### v1

Abstract
This paper extends the Isotonic Mechanism from the single-owner to multi-owner settings, in an effort to make it applicable to peer review where a paper often has multiple authors. Our approach starts by partitioning all submissions of a machine learning conference into disjoint blocks, each of which shares a common set of co-authors. We then employ the Isotonic Mechanism to elicit a ranking of the submissions from each author and to produce adjusted review scores that align with both the reported ranking and the original review scores. The generalized mechanism uses a weighted average of the adjusted scores on each block. We show that, under certain conditions, truth-telling by all authors is a Nash equilibrium for any valid partition of the overlapping ownership sets. However, we demonstrate that while the mechanism’s performance in terms of estimation accuracy depends on the partition structure, optimizing this structure is computationally intractable in general. We develop a nearly linear-time greedy algorithm that provably finds a performant partition with appealing robust approximation guarantees. Extensive experiments on both synthetic data and real-world conference review data demonstrate the effectiveness of this generalized Isotonic Mechanism.
1 Introduction
In recent years, major machine learning conferences such as NeurIPS and ICML have faced a concerning decline in the quality of peer review, which poses a significant challenge to the global machine learning community [16, 5, 24]. In particular, the NeurIPS 2014 experiment showed that $49.5\%$ of the papers accepted by one committee would be rejected by another [17, 8, 16]. This inconsistency probability was $50.6\%$ for NeurIPS 2021 [6]. This worrying trend is, in part, due to the surge in submission volumes [20]. For example, NeurIPS 2022 witnessed a record-breaking 10,411 submissions. In contrast, as demonstrated in [15, 19, 21], the growth rate of qualified reviewers lags significantly behind this surge in submissions.
Figure 1: An example of an author-paper ownership set shown as a bipartite graph. An edge between an individual and a paper means that this individual is the author of the paper.
To mitigate this issue, there has been a progressive trend to propose various strategies aimed at enhancing the peer review process in machine learning (see, e.g., [20]). These strategies include improved reviewer assignments, the incorporation of additional questions, and many others [9, 14, 25, 13, 12]. Among these, an emerging approach employs mechanism design techniques to solicit private information from authors, thereby enabling a more accurate estimation of review scores [22, 23]. This approach, termed the “Isotonic Mechanism”, solicits authors of multiple submissions to rank their papers based on their perceptions. The adjusted review scores maintain consistency with the author-provided ranking while remaining as close as possible to the original review scores.
To formally describe this method, imagine that an author, say, Alice, submits $n$ papers to a conference. Right after submission, Alice under the Isotonic Mechanism provides a ranking $\pi$ in the form of a permutation of $1,2,\ldots,n$ that sorts her papers in descending order of quality. Letting $y_{1},\ldots,y_{n}$ denote the (average) review scores of the $n$ papers, the Isotonic Mechanism yields adjusted review scores that are the solution to the following convex optimization program:
$\mathop{min}_{\bm{r}\in\mathbb{R}^{n}}$ $\sum_{i=1}^{n}(y_{i}-r_{i})^{2}$
s.t. $r_{\pi(1)}\geq r_{\pi(2)}\geq\dots\geq r_{\pi(n)}.$
Under certain utility assumptions, [22] shows that Alice is incentivized to provide the true ranking under the Isotonic Mechanism, which in turn improves the estimation accuracy of the “ground-truth” review scores. The improvement is substantial when the number of submissions is large and the review scores contain a high level of noise. Moreover, [23] proves that the ranking of the submissions is the most fine-grained information that can be truthfully elicited from the author.
The Isotonic Mechanism was recently experimented at ICML 2023111See https://icml.cc/Conferences/2023/CallForPapers.. However, this method in its current form cannot fully adapt to practical settings wherein papers have multiple authors. Specifically, it is not clear how the rankings of Alice’s co-authors would affect the utility of Alice. A more challenging situation is where different papers partially overlap in authorship, as illustrated in Figure 1. In such a case, Alice’s utility might be affected by other authors’ rankings that include papers Alice did not contribute, if one applies the Isotonic Mechanism to every author’s submissions and subsequently averages the adjusted review scores for each paper. This is an unintended effect that we aim to eliminate in an ideal mechanism.
In this paper, we introduce a generalized Isotonic Mechanism to address the challenge of overlapping authorship in large machine learning conferences. Our approach begins by partitioning the author-paper ownership sets into blocks of papers, ensuring that each block shares a common set of authors. We then apply the Isotonic Mechanism to each block, soliciting paper rankings from each common author, and outputting a weighted average of all adjusted review scores from all calls of the Isotonic Mechanism for this block. Under certain conditions, this generalized mechanism is marked by the following key characteristic:
Main Result 1.
Truthfully reporting rankings by all authors forms a Nash equilibrium under the generalized Isotonic Mechanism.
Along with other results, the truthfulness result is formally presented in Section 3. This result suggests that the optimal strategy for an author to maximize her utility is to report the ground-truth ranking, given that all the other authors do likewise. Its proof involves a novel technique concerning majorization. Additionally, we demonstrate that this Nash equilibrium is the most favored, as any other potential Nash equilibrium would result in lower utility for every author.
As is conceivable, the statistical performance of this generalized Isotonic Mechanism depends on the partition. For example, a partition for Figure 1 can be simply $\{p_{1},p_{2},p_{3},p_{4},p_{5}\}$, which share a single common author $\{a_{3}\}$, or can be $\{p_{1},p_{2},p_{3}\},\{p_{4}\}$, and $\{p_{5},p_{6}\}$, which share $\{a_{1},a_{2},a_{3}\},\{a_{3}\}$, and $\{a_{3},a_{4}\}$, respectively. To find high-quality author-paper ownership partitions, we define a natural class of criteria to assess the quality of a partition. Within this class of criteria, we obtain our second main result concerning the generalized Isotonic Mechanism:
Main Result 2.
A simple linear-time greedy algorithm can obtain a partition of the author-paper ownership set for the generalized Isotonic Mechanism that is simultaneously near-optimal for every criterion in the class.
This result is formally stated in Section 4. In the same section, we show that it is NP-hard to find the exact optimal partition even for a certain simple criterion. In contrast, the greedy algorithm is computationally efficient while achieving constant-ratio approximation simultaneously for most natural criteria.
We validate this mechanism through a series of experiments on both ICLR review data and synthetic data. Our experiments exhibit that this generalized Isotonic Mechanism enhances estimation performance relative to the original review scores, a benefit observed consistently across various experimental settings.
2 Notation and Preliminaries
This paper extends the problem setup in [23] to a general model of overlapping ownership. There are $m$ owners and $n$ items. The $j$-th owner owns a set of items $\mathcal{I}^{j}\subseteq[n]=\{1,2,\dots,n\}$ and let $|\mathcal{I}^{j}|=n_{j}$. The ownership relation can be visualized as a bipartite graph (e.g., see Figure 1 and their formal connections in Section 6.2). For each item $i=1,\ldots,n$, let $R_{i}$ be its ground-truth quality and
$y_{i}=R_{i}+z_{i}$
be a noisy rating of its quality by reviewers. As a friendly reminder of our notation, $i$ and $\mathcal{I}$ are used to denote item index and item set, respectively. Denote by $\bm{z}=(z_{1},z_{2},\dots,z_{n}),\bm{R}=(R_{1},R_{2},\dots,R_{n})$ and $\bm{y}=(y_{1},y_{2},\dots,y_{n})$, respectively, the vector of noise, ground-truth item scores and review scores. Let $\mathrm{U}^{j}:\mathbb{R}^{n_{j}}\to\mathbb{R}$ be the $j$-th owner’s utility function on the set of items she owns. We adopt the same assumptions (listed below) as [23] about owner knowledge, noise distribution and utility.
Assumption 2.1 (Informed Owners [23]).
For each $j$, the $j$-th owner has sufficient knowledge of the ground-truth item quality $\bm{R}$ to determine the true ranking of her own items $\mathcal{I}^{j}$.
Assumption 2.2 (Exchangeable Noise Distribution [23]).
The review noise vector $\bm{z}=(z_{1},\dots,z_{n})$ follows an exchangeable distribution in the sense that $(z_{1},\dots,z_{n})$ has the same probability distribution in $\mathbb{R}^{n}$ as $\rho\circ\bm{z}:=(z_{\rho(1)},\dots,z_{\rho(n)})$ for any permutation $\rho$ of $1,2,\dots,n$.
Assumption 2.3 (Convex Utility [23]).
For each $j$, the $j$-th owner’s utility function takes the form of $\mathrm{U}^{j}(\widehat{\bm{R}})=\sum_{i\in\mathcal{I}^{j}}{U}^{j}(\widehat{R}_{i})$, where ${U}^{j}:\mathbb{R}\to\mathbb{R}$ is a non-decreasing convex function.
For simplicity of presentation, we focus on the most standard knowledge partition specified by the $n!$ permutations of the ordinal order of the item qualities. For Assumption 2.3, the utility function can be slightly relaxed to a non-separable utility function $\mathrm{U}^{j}:\mathbb{R}^{n_{j}}\to\mathbb{R}$ that satisfies Schur-convexity, namely,
$(x_{i}-x_{i^{\prime}})\left(\frac{\partial\mathrm{U}^{j}(\bm{x})}{\partial x_{i}}-\frac{\partial\mathrm{U}^{j}(\bm{x})}{\partial x_{i^{\prime}}}\right)\geq 0,$
for all $i,i^{\prime}\in[n_{j}]$. Notably, Assumption 2.3 does not require ${U}^{j}={U}^{j^{\prime}},\forall j\neq j^{\prime}$, so each owner may have different utility on the same scoring of their items. We refer the readers to [23] for the detailed discussions and justification for these assumptions.
At a high level, our goal is to design a scoring mechanism $\mathcal{M}$ to estimate item qualities $\bm{R}$. The mechanism is expected to mitigate noises in review scores $\bm{y}$ by incentivizing as many owners to truthfully share as much information on their item qualities as possible. The key to this special kind of mechanism is the design problem of a statistical estimator $\widehat{\mathrm{R}}_{\mathcal{M}}$ about the item qualities $\bm{R}$. To improve on the observed review score $\bm{y}$, the estimator will additionally account for the ranking information elicited from the owners $\{\pi^{j}\}_{j=1}^{m}$. Motivated by the practical applications of conference peer review, we require the mechanism to proceed in the following stages:
  1. 1.
    Before Submission: Each owner is notified of the scoring mechanism and obtains some knowledge on the qualities of the items that she owns.
  2. 2.
    During Submission: Each owner $j$ reports a ranking $\pi^{j}$ of her items $\mathcal{I}^{j}$ [^2] If an owner does not report a ranking, a uniformly random ranking will be used. As will be shown later, this design ensures rational owners participate in the mechanism. It is also without loss of generality to assume that owners report the full ranking of their items, and we defer the discussion to Remark 1.2.
  3. 3.
    After Submission: Once the review score $\bm{y}$ is realized, the mechanism accordingly estimates the item qualities as $\widehat{\bm{R}}=\widehat{\mathrm{R}}_{\mathcal{M}}(\{\pi^{j}\}_{j=1}^{m};\bm{y})$. Each owner $j$ then derives utility $\mathrm{U}^{j}\left([\widehat{R}_{i}]_{i\in\mathcal{I}^{j}}\right)$.
We use the standard Bayesian game framework [10] to model the owners’ strategic decision-making under the uncertainty of realized review scores $\bm{y}$. We say a profile of owners’ report $\{\pi^{j}\}_{j=1}^{m}$ forms a Bayes-Nash Equilibrium (NE) under mechanism $\mathcal{M}$  [^3] It suffices to work with this special set of pure-strategy NEs in most part of this paper and we defer the additional notations for the more general definition of mixed-strategy NEs to Appendix 1. if for any owner $j\in[m]$, given others’ report $\pi^{-j}=\{\pi^{j^{\prime}}\}_{j^{\prime}\neq j}$, $j$’s expected utility by reporting $\pi^{j}$ is no worse than reporting any other possible ranking $\widetilde{\pi}^{j}$ or, mathematically,
[…]
We say a mechanism $\mathcal{M}$ is truthful if every owner $j$ reporting the true ranking $\pi^{j}$ of her items forms a Bayes-Nash equilibrium under scoring mechanism $\mathcal{M}$.
3 Truthfulness under Completely Overlapping Ownership
The Isotonic Mechanism [23] in case of a single author uses any reported ranking $\pi$ from the owner and the realized review scores $\bm{y}$ for his items to solve for the item quality estimates,
$\widehat{\mathrm{R}}(\pi;\bm{y})=\mathop{argmin}_{\bm{r}\in\mathbb{R}^{n}}$ $\|\bm{y}-\bm{r}\|^{2}$
s.t. $r_{\pi(1)}\geq r_{\pi(2)}\geq\dots\geq r_{\pi(n)},$
[…]
Now returning to our more general setup with the overlapping ownership, to aggregate information from multiple owners, the most natural design is perhaps to use the (weighted) averaged scores from the quality estimates based on each owner’s reported ranking. This leads to our extension of the Isotonic Mechanism [23] for the multi-owners setup, as presented in Mechanism 1. In this mechanism, one can simply set the weight parameter $\alpha^{j}=1/m$ to evenly weigh on each owner’s reported ranking, or set any $\alpha^{j}\in[0,1]$ for each owner $j$ (e.g., due to their expert levels, track records) as a convex combination of scores calibrated by each owner’s reported ranking — we leave such kind of implementation choices to the practitioner’s discretion.
Input : Review scores $\bm{y}\in\mathbb{R}^{n}$, Weights $\{\alpha^{j}\}_{j=1}^{m}.$
for  $j\in[m]=\{1,2,\dots,m\}$  do
       Elicit ranking $\pi^{j}$ from owner $j$.
       Solve for ranking-calibrated scores $\widehat{\bm{R}}^{j}\leftarrow\widehat{\mathrm{R}}(\pi^{j};\bm{y}).$
Set $\widehat{\bm{R}}\leftarrow\sum_{j=1}^{m}\alpha^{j}\widehat{\bm{R}}^{j}$.
return $\widehat{\bm{R}}$.
Mechanism 1 Isotonic Mechanism under Completely Overlapping Ownership
In this section, we consider the setting where every submission has the same set of authors, referred to as the special case of completely overlapping ownership. As it turns out, truth-telling forms a Nash equilibrium in this setting under Mechanism 1444This result can be related to the ex-post incentive compatibility of Mechanism 1, a common goal of truthful mechanism design. Moreover, this equilibrium is payoff dominant [11], i.e., one that is Pareto superior to all other Nash equilibria in the game. Put simply, all owners would prefer this equilibrium, because it simultaneously gives every owner the highest equilibrium utility among all possible equilibrium outcomes. This nice property makes it more plausible to expect agents’ truthful behaviors, despite the potential existence of multiple equilibria in the game; more discussions about this mechanism’s truthfulness from the perspective of behavioral theory can be found in Remark 1.1.
Theorem 1.
Under Assumptions 2.1, 2.2, and 2.3, if the ownership is completely overlapping,
  1. 1.
    It forms a Bayes-Nash equilibrium for each owner to truthfully report the ranking in Mechanism 1.
  2. 2.
    This equilibrium of truthful report is payoff dominant.
Remark 1.1 (Additional Discussions on the Strategic Game Induced by the Isotonic Mechanism).
The strategic game induced by Mechanism 1 features a certain structure that is known as the common interest game in the literature [11, 4]. We briefly highlight this connection here and discuss evidence from both empirical human preferences and behavioral theory about why truthful behavior should be expected in such games. In its simplest form, the payoff matrix of a common interest game typically takes one of the three structures illustrated in Table 1 (only orders of the payoff values matter). Namely, there is an action profile (the upper left cell of each game) that represents the common interest of both players; this corresponds to the truthful action profile $(\pi^{\star},\pi^{\star})$ under the Isotonic Mechanism which forms a payoff-dominant Nash equilibrium, according to Theorem 1. Meanwhile, the three variants of common interest games capture the different possible orders of payoff values in the lower right cell and those in the off-diagonal cells. Any of these payoff variants may be realized under the Isotonic Mechanism, since our proof using the Jensen’s inequality only offers upper bounds on the utilities of these non-truthful action profiles (see e.g., Equation (4)).
$\pi^{\star}$ $\pi^{\circ}$
$\pi^{\star}$ $10,11$ $4,5$
$\pi^{\circ}$ $4,5$ $3,4$
$\pi^{\star}$ $\pi^{\circ}$
$\pi^{\star}$ $10,11$ $1,1$
$\pi^{\circ}$ $1,1$ $3,4$
$\pi^{\star}$ $\pi^{\circ}$
$\pi^{\star}$ $10,11$ $1,5$
$\pi^{\circ}$ $5,1$ $3,4$
Table 1: An illustration of three typical common interest game payoff matrices.
It turns out that in experimental games and also in real life, people tend to always choose the upper left cell — the payoff dominant outcomes — in all the three possible game structures above [4]. This behavior is expected under the first payoff matrix since $(\pi^{\star},\pi^{\star})$ is the unique Nash equilibrium there. It is somewhat reasonable under the second payoff matrix, as argued by Harsanyi et al. [11] for its “self-reinforcing property” — everyone thinks the other has no reason to prefer the payoff-dominated equilibrium outcomes. However, realizing the payoff dominant outcome becomes less clear under the third payoff matrix, because $(\pi^{\circ},\pi^{\circ})$ forms a “risk-dominant” Nash equilibrium — i.e., compared to $\pi^{\star}$, $\pi^{\circ}$ is the less risky action under uncertainty of the opponent’s choice [11] (specifically, under $(\pi^{\circ},\pi^{\circ})$ equilibrium, row player gets $5$ instead of $3$ if the opponent deviates from the equilibrium action $\pi^{\circ}$ to $\pi^{\star}$). Interestingly, standard game theory cannot account for this phenomenon of coordination on the payoff dominant outcome under the second and third payoff matrices, i.e., the Hi-Lo Paradox. One major theory that explains this phenomenon is the team reasoning by Bacharach [4] as a non-game-theoretic rationale, i.e., “What do we want? And what should I do to play my part in achieving this?”. If there is common knowledge that both players adopt the team-reasoning mode of choosing their strategies, then both would choose the payoff dominant equilibrium. These also serve as behavioural evidences of agents’ truthfulness in the Isotonic Mechanism.
3.1 Non-truthfulness Beyond Completely Overlapping Ownership.
Despite the stability of truthful NE shown in Theorem 1, the ownership relation is not complete in many application scenarios, e.g., different papers are written by different sets of authors. A natural extension of Mechanism 1 for the case of incompletely overlapping ownership is to average the scores according to the ownership, $\widehat{R}_{i}\leftarrow{\sum_{j=1}^{m}e^{j}_{i}\widehat{R}^{j}_{i}}/{\sum_{j^{\prime}=1}^{m}e^{j^{\prime}}_{i}}$, where the binary ownership indicator $e^{j}_{i}=1$ if and only if author $j$ owns item $i$. The question is whether truthful reporting still forms a Nash equilibrium. Unfortunately, the answer turns out to be “No”. Specifically, the owners may gain from the misreporting strategy and one possible manipulation is to use the reputation of good items to promote a bad item — it may only hurt the good item a little, but help the bad item a lot. We demonstrate such kind of strategic manipulations via a concrete example below.
Example 1 (Non-truthfulness of the Isotonic Mechanism under Partially Overlapping Ownership).
Consider a case of $m$ owners and $n=3$ items. The items have ground-truth qualities $R_{1}=9,R_{2}=8,R_{3}=4$ — one weak and two strong. The ownership is not completely overlapping. The first $m-1$ owners work together on the two strong items $1,2$, while the $m$-th owner works on the item $2,3$. Let the $j$-th owners’ utility be $\mathrm{U}^{j}(\widehat{\bm{R}})=\sum_{i\in\mathcal{I}^{j}}\mathop{max}\{R_{i}-5,0\}$. For simplicity, suppose the reviews are noiseless so that, if all owners report their ranking truthfully, the calibration is perfect, i.e., $\widehat{\bm{R}}=\textbf{y}=\bm{R}$. See Table 2 for an illustration.
True Scores
$R_{1}=9$
$R_{2}=8$
$R_{3}=4$
Item Owner $1$ $2$ $m-1$ $m$
1 1 1 1 0
2 1 1 1 1
3 0 0 0 1
Table 2: An illustration of item scores and the ownership matrix $E=(e_{i}^{j})_{m\times n}$ in Example 1
However, the $m$-th owner does not have the incentive to report the ranking of items $1,2$ truthfully, given that the first $m-1$ owners report the true ranking. That is, under the flipped ranking that $\widetilde{R}_{1}^{m}\geq\widetilde{R}_{2}^{m}$, Mechanism 1 would have $\widetilde{R}_{1}^{m}=\widetilde{R}_{2}^{m}=6$ and averaged calibrated scores $\widetilde{R}_{1}=9,\widetilde{R}_{2}=\widetilde{R}_{2}^{m}+\sum_{j=1}^{m-1}R_{2}^{j}=8-\frac{2}{m},\widetilde{R}_{3}=\widetilde{R}_{3}^{m}=6$. We can see that the utility of $m$-th owner under truthful ranking is, $13<3-\frac{2}{m}+$ 1, strictly worse than that under the non-truthful ranking for any $m\geq 3$.
At a high level, such strategic behavior is due to the owner’s uneven influence on the average scores of its different items. One may speculate that we can potentially resolve this kind of situation through a careful reweighing process of each owner’s influence. The answer also turns out to be “No”. Specifically, in the example above, $m$-th owner, as the sole owner of item 1, would always fully determine the score of item 1; yet, this owner can never fully determine the score of item 2 unless we choose to ignore other owners’ opinion or assign some weight to the raw review scores, which more or less defeats the purpose of this community effort of calibrating review scores. Hence, we need to seek a different approach to help us truthfully aggregate the authors’ information in practice.
4 Restoring Truthfulness via Partitioning
To ensure the truthfulness of the Isotonic Mechanism beyond complete ownership, this section studies a partition-based approach. Our main idea is to partition the ownership sets into multiple blocks, ensuring complete overlap within each block by some owners. We then apply the Isotonic Mechanism separately to each block, eliciting truthful rankings only from the owners who completely own that block. For example, this includes authors who wrote all the papers within that block (refer to the illustrations in Figure 2 and Table 3). Finally, we use the elicited rankings from these blocks to estimate the item quality. This generalized approach is described in the following Mechanism 2.
Input : Review score $\bm{y}\in\mathbb{R}^{n}$.
Input : Any partition ${\mathcal{S}}=\{{\mathcal{S}}_{1},\ldots,{\mathcal{S}}_{K}\}$ of all items (i.e., $\bigcup_{k=1}^{K}{\mathcal{S}}_{k}=[n]$ and ${\mathcal{S}}_{k}\cap{\mathcal{S}}_{k^{\prime}}=\emptyset$).
for every ${\mathcal{S}}_{k}\in{\mathcal{S}}$ do
       Find all owners with complete ownership of ${\mathcal{S}}_{k}$, i.e., ${\mathcal{T}}_{k}\leftarrow\{j\in[m]:{\mathcal{S}}_{k}\subseteq\mathcal{I}^{j}\}$.
       if $|{\mathcal{T}}_{k}|=0$ or $|{\mathcal{S}}_{k}|=1$  then
            Do not elicit any ranking information.
      else
             Apply Mechanism 1 to item set ${\mathcal{S}}_{k}$ with owner set ${\mathcal{T}}_{k}$ to estimate the quality of items in ${\mathcal{S}}_{k}$, denoted by $\widehat{\bm{R}}[{\mathcal{S}}_{k}]$.
return $\widehat{\bm{R}}=\{\widehat{\bm{R}}[{\mathcal{S}}_{k}]\}_{k=1}^{K}$.
Mechanism 2 Partition-based Isotonic Mechanism for Partially Overlapping Ownership
We remark that in Mechanism 2, each owner $j\in{\mathcal{T}}_{k}$ has the complete ownership of the items in ${\mathcal{S}}_{k}$ by construction. As a corollary of Theorem 1, it is easy to show that Mechanism 2 is truthful under any partition ${\mathcal{S}}=\{{\mathcal{S}}_{1},\ldots,{\mathcal{S}}_{K}\}$ (which is part of the input to Mechanism 2), because we only elicit rankings of blocks from those who completely own the block. Thus, all owners $j\in{\mathcal{T}}_{k}$ will report truthful ranking over items in ${\mathcal{S}}_{k}\subseteq\mathcal{I}^{j}$. Owing to the independence of quality estimates for items between different partition blocks, the overall mechanism is truthful; we formalize it in the following corollary.
Corollary 1.
Under Mechanism 2 with any item set partition, it forms a Bayes-Nash equilibrium for each owner to truthfully report the ranking of items within each block specified by the partition.
Remark 1.2 (Practical Implementation of Mechanism 2).
While Mechanism 2 only elicits ranking of items within each block ${\mathcal{S}}_{k}$, in practical implementations, the mechanism designer may simply ask the owners to rank all their items but commit to only use part of each owner’s ranking as specified by the mechanism. Under such a committed mechanism, any owner will be indifferent about the order of any two items across any two different partition sets ${\mathcal{S}}_{k}$ and ${\mathcal{S}}_{k^{\prime}}$ since their order will never be used by the mechanism. Thus, the owner can be assumed to break ties in favor of the designer and reveal their full ranking truthfully. [^5] There is also a more game-theoretic way to justify such tie-breaking by adding a negligible amount of randomness for picking an arbitrarily different partition, which consequently creates a negligible amount of incentive for any owner to report other rankings truthfully. Such equilibrium refinement via randomization has appeared in mechanism design literature [18]. Notably, however, the designer’s commitment to not using any order information beyond what the mechanism specified (despite such information being elicited) is important since otherwise, truthfulness will not hold. In reality, the mechanism designer as a trusted authority (e.g., the organizers of a large ML conference) has such commitment power.
[Uncaptioned image] Figure 2: A partition of partially overlapping ownership Item Owner $1$ $2$ $3$ $4$ $1$ 1 1 1 0 $2$ 1 1 1 0 $3$ 1 1 1 0 $4$ 0 0 1 0 $5$ 0 0 1 1 $6$ 0 0 1 1 Table 3: A visualization via table
While any partition will lead to a truthful Mechanism 2, what partition we choose will significantly affect how much information we can elicit from owners to refine our quality estimation. Intuitively, the smaller the size of a block is, the less ranking information we can elicit. [^6] For instance, the ranking of a size-$10$ item block has $\frac{10\times 9}{2}=45$ pairs of comparison, whereas the rankings of two size-$5$ item blocks have $202\times\frac{5\times 4}{2}=$ 20 pairs of comparison. Thus, the core of our partition-based approach is to design an efficient algorithm to compute high-quality partition structure, which we now elaborate on.
4.1 Partition Optimization and its Hardness
As illustrated above, the choice of the item partition plays an important role in both the performance and robustness of Mechanism 2. Below, we formalize two different metrics for partitions, which determine the performance and robustness of the resulting mechanism, respectively. We will study their implications for the efficient computation of a high-quality partition, as well as their intrinsic tradeoff.
  • Informativeness (as optimization objective). In regard to the performance, we aim to identify partitions that are most informative for the review score calibration. As mentioned previously, generally larger-sized blocks are preferred — in the extreme case, if a block contains only one item, the mechanism cannot elicit any ranking information on this single item in the block. However, to get large blocks, it may sometimes be unavoidable to have a small amount of small or even trivial blocks. This creates a complex tradeoff about partition choices. To systemically evaluate the coverage quality of a partition ${\mathcal{S}}$, we introduce an informativeness objective function
    $\operatorname{obj}({\mathcal{S}})=\sum_{k=1}^{K}w(|{\mathcal{S}}_{k}|),$
    determined by any weight function $w$ on the size of each partition block. For example, two choices of weight function $w(\cdot)$ are of particular interest. One is $w(x)=x^{2}$ as a simple quadratic function of block size, which relates to the number of pairwise comparisons within a block:
[…]
    Another natural choice is $w(x)=\mathop{max}\{x-1,0\}$, which is called size-based defined as follows:
[…]
    where the second equation is because any block in the partition has at least $1$ item (i.e., $|{\mathcal{S}}_{k}|\geq 1$). Since $n$ is a constant, maximizing the size-based objective is equivalent to minimizing the size of the partition $K$ (thus the name “size-based”).
    Generally, it is reasonable to expect that function $w(\cdot)$ should satisfy increasing monotonicity and convexity — i.e., the larger the size is, the faster the objective increases, and the better the Isotonic Mechanism performs. Another natural requirement is $w(0)=0$ in order to prevent empty blocks in the optimal partitions. Both functions above satisfy these properties, and so do any $w(x)=x^{p}$ for $p\geq 1$. However, it appears to be an intractable task to precisely describe the relationship between the structure of partition and the statistical guarantee of the mechanism. Thus, in Section 4.2, we seek to overcome this challenge by designing algorithms that can simultaneously perform well for many natural weight functions satisfying these properties.
  • Robustness (as optimization constraint). If the informativeness metric above is viewed as the objective we want the partition to maximize, then the robustness metric can be viewed as a constraint we wish the partition to satisfy. In regard to robustness, aggregating information from multiple owners helps to mitigate each individual owner’s noisy perception of their items. Thus, the minimal number of owners for each partition block matters. Formally, we say a partition ${\mathcal{S}}=\{{\mathcal{S}}_{1},\ldots,{\mathcal{S}}_{K}\}$ is $L$-strong if there are at least $L$ owners who share all items in the same partition block, i.e., $\left|{\mathcal{T}}_{k}\right|\geq L,\forall{\mathcal{S}}_{k}\in{\mathcal{S}}$ with $|{\mathcal{S}}_{k}|>1$. The larger $L$ is, the less prone the Isotonic Mechanism is to each individual owner’s randomness in her perception. Notably, $L\geq 1$ is the bare minimum requirement for Mechanism 2 to work. It is intuitive that robustness and informativeness are a tradeoff in general: a larger $L$ necessarily leads to less informative partitions, since a stricter constraint necessarily leads to a worse objective value.
Fortunately, generating a feasible $L$-strong partition reduces to generating a feasible $1$-strong partition for a different ownership instance, as shown in the following proposition. This reduction helps us to convert any $L$-strongness constraint to a certain $1$-strongness constraint during our design of the optimization algorithm.
Proposition 1 (A Reduction from $L$-Strong to $1$-Strong Partition).
For any ownership instance $\mathcal{O}$, we can construct in $O\left(m^{L}n\right)$ time a different ownership instance $\mathcal{O}^{\prime}$ such that any $L$-strong partition ${\mathcal{S}}$ in $\mathcal{O}$ is a 1-strong partition in $\mathcal{O}^{\prime}$.
Proposition 1 shows that to optimize the informativeness objective, it is without loss of generality to design algorithms to maximize the objective under $1$-strongness, for a transformed ownership set instance. A simple example to illustrate this reduction: any $2$-strong partition in the ownership instance $\mathcal{O}=\{\{1,2\},\{1,2,3\},\{2,3\}\}$ is a $1$-strong partition in the ownership instance $\mathcal{O}^{\prime}=\{\{1,2\},\{2\},\{2,3\}\}$. Essentially, the instance $\mathcal{O}^{\prime}$ is constructed to have an owner with items shared by every $L$ owner in the original instance $\mathcal{O}$; we leave the details to the proof in Appendix. This is notably a polynomial reduction for any constant $L$ — e.g., the number of co-authors on each paper is typically small. However, this simplification does not fundamentally simplify the problem due to the intrinsic hardness of maximizing the informativeness objective, as shown in the following proposition.
Proposition 2 (Hardness of Partition Optimization).
It is NP-hard to find the optimal item partition ${\mathcal{S}}=\{{\mathcal{S}}_{1},\ldots,{\mathcal{S}}_{K}\}$ that maximizes the size-based objective in Equation (2) subject to $1$-strongness.
4.2 Fast Greedy Partition with Robust Approximation Guarantees
Despite the computational hardness for the optimal partition, we show that a natural greedy algorithm (as also used in [23]) can efficiently find partitions with a provably good approximation guarantee. In words, the greedy algorithm just iteratively selects the largest residual paper set owned by some owner. This algorithm can be implemented in time almost linear in the input size $\sum_{j\in[m]}{|\mathcal{I}^{j}|}$, up to a $\log(m)$ factor, and thus is essentially the fastest algorithm one can design.
The highlight of this greedy algorithm is that, though the algorithm itself is fully agnostic to any partition objectives, its output is a robustly high-quality partition for every natural weight function, i.e., monotone and convex weight functions, as formalized in the following Theorem 2. In particular, consider the following hypothesis class $\mathcal{W}$ of weight functions
[…]
Algorithm 3 enjoys the following strong approximation guarantees for every $w$ within $\mathcal{W}$.
Theorem 2 (Robust Approximation of Greedy).
For any ownership instance $\{\mathcal{I}^{j}\}_{j\in[m]}$ with $N=\sum_{j\in[m]}{|\mathcal{I}^{j}|}$ total number of author-paper pairs, Algorithm 3 runs in $O(N\log m)$ time and outputs a 1-strong partition that is simultaneously a $c(w)$-approximation of the optimal objective for every $w\in\mathcal{W}$, where
$c(w)=\inf\Big{\{}\frac{w(x)}{w^{\prime}_{-}(x)x}\Big{|}w^{\prime}_{-}(x)>0,x\geq 2\Big{\}}.$
Above, $w^{\prime}_{-}(a)=\lim_{x\to a^{-}}\frac{w(x)-w(a)}{x-a}$ denotes the left derivative of $w$ at $a$.
Input : Ownership sets $\{\mathcal{I}^{j}\}_{j\in[m]}$.
Initialize the partition as ${\mathcal{S}}=\{\}$ and set of selected items $\overline{\mathcal{I}}=\varnothing$.
while  ¯[n]$\overline{\mathcal{I}}\subset[n]$ do
       Determine the largest residual paper set ${\mathcal{S}}^{*}=\mathcal{I}^{j^{*}}\setminus\overline{\mathcal{I}}$ where $j^{*}=\mathop{argmax}_{j\in[m]}|\mathcal{I}^{j}\setminus\overline{\mathcal{I}}|$.
       Update ¯¯$\overline{\mathcal{I}}\leftarrow\overline{\mathcal{I}}\cup{\mathcal{S}}^{*}$ and ${\mathcal{S}}\leftarrow{\mathcal{S}}\cup\{{\mathcal{S}}^{*}\}$.
return ${\mathcal{S}}$.
Algorithm 3 A Greedy Algorithm for $1$-Strong Partition [23]
Remark 2.1.
Theorem 2 readily implies that the partition found by the greedy algorithm simultaneously approximates at least $1/2$ of both the optimal comparison-based objective and size-based objective. More generally, for $\alpha$-th degree polynomial functions $w(x)=x^{\alpha}$, the greedy algorithm simultaneously guarantees $(1/\alpha)$-fraction of the optimal objective since $w(x)/[w^{\prime}(a)x]=1/\alpha$ for every $x$. Such objective-agnostic property of Algorithm 3 is especially desirable in our application scenarios where the concrete form of the partition objective is difficult (if not impossible) to know.
Moreover, the approximation ratio in Theorem 2 for the greedy algorithm is provably tight under every monomial objective $f({\mathcal{S}})=\sum_{k=1}^{K}|{\mathcal{S}}_{k}|^{\alpha}$, $\alpha>1$. Lower bound instances can be found in Section 6.2.1.
Remark 2.2.
While Algorithm 3 only finds a partition satisfying $1$-strongness, the algorithm as well as its constant approximation guarantees can be readily extended to find the optimal $L$-strong partition in $O\left(m^{L}n\right)$ time, using the reduction in Proposition 1.
5 Experiments
One major challenge in evaluating the proposed scoring mechanism is that their performance measures rely on the underlying true item qualities. Such information is unattainable in most applications — indeed, if we already know the true qualities, peer review would not be needed anymore. Fortunately, we are able to collect both the ownership relations and the review scores from ICLR 2021-2023 [1, 2, 3] that are made publicly available from openreview.net. Based on the recorded review scores, we simulate the true item qualities $\bm{r}=\textbf{y}-\bm{z}$, where $\bm{z}\sim\mathcal{N}(0,\sigma^{2})$ is a zero-mean Gaussian random variable with standard deviation $\sigma\in\{1,2,3,4\}$. We use the proposed Isotonic Mechanism to determine the calibrated score $\widehat{\bm{R}}$ from review scores y and the elicited ranking information. For comparison, we also consider the alternative calibration approaches: baseline, which directly uses the review scores; random, which runs the Isotonic Mechanism under a randomly generated partition. We measure the performance of the proposed mechanism by the mean squared error, $\frac{1}{n}\|\bm{R}-\widehat{\bm{R}}\|^{2}$, and the precision@kmetric, $\frac{|\widehat{\mathcal{P}}_{@k}\cap{\mathcal{P}}_{@k}|}{|{\mathcal{P}}_{@k}|}$, where ${\mathcal{P}}_{@k}$ (resp., $\widehat{\mathcal{P}}_{@k}$) is the set of items within $k$-th percentile with regard to the true quality $\bm{R}$ (resp., calibrated score $\widehat{\bm{R}}$). We choose the percentile $k\in\{1.5,5,30\}$ to roughly match with the criteria of “oral”, “spotlight” and “poster” presentations at ICLR, respectively.
#authors #papers $\operatorname{obj}({\mathcal{S}}_{\text{greedy}})$ $\operatorname{obj}^{\prime}({\mathcal{S}}_{\text{greedy}})$ $\operatorname{obj}({\mathcal{S}}_{\text{random}})$ $\operatorname{obj}^{\prime}({\mathcal{S}}_{\text{random}})$
ICLR 2021 8875 2964 5242 1365 2515 1132
ICLR 2022 10583 3328 6123 1574 2966 1324
ICLR 2023 15372 4881 9764 2489 4600 2045
Table 4: Statistics of ICLR 2021-2023, where ${\mathcal{S}}_{\text{greedy}},{\mathcal{S}}_{\text{random}}$ are greedy and random partitions generated in the ownership graph of each year’s conference, $\operatorname{obj},\operatorname{obj}^{\prime}$ are respectively the comparison-based and size-based objective, defined in Equations (1) and (2).
In Table 4, we summarize the basic statistics of ICLR 2021-2023 that we collected as well as the partitions determined from their ownership sets. There is a clear growing trend in conference size in terms of submission and author number. Moreover, from both partition objective functions, the greedy partition is a clearly better choice than the randomly generated partition, though we are unable to efficiently find the exact optimal partition and empirically examine the greedy approximation ratio.
(a) ICLR 2021
(b) ICLR 2022
(c) ICLR 2023
Figure 3: The loss (normalized as the percentage change, $\frac{\text{model}-\text{baseline}}{\text{baseline}}$) of Isotonic Mechanism under varying level of review noise $\sigma$ on ICLR.
(a) ICLR 2021 Oral
(b) ICLR 2022 Oral
(c) ICLR 2023 Oral
(d) ICLR 2021 Spotlight
(e) ICLR 2022 Spotlight
(f) ICLR 2023 Spotlight
(g) ICLR 2021 Poster
(h) ICLR 2022 Poster
(i) ICLR 2023 Poster
Figure 4: The percentile precision of Isotonic Mechanism under varying level review noise $\sigma$ on ICLR.
We compare the empirical performance of the Isotonic Mechanism with different baselines at different review noise levels. In Figures 3 and 4, we can see that the proposed Isotonic Mechanism can mitigate a substantial amount of noise from the review score and improve the precision metric over the baseline approach. The percentage loss change in Figure 3 suggests our proposed method is able to mitigate about $10-20\%$ amount of review noise in the past three years of ICLR. In addition, the performance of the Isotonic Mechanism steadily improves as the conference size grows — an encouraging sign for the applicability of the Isotonic Mechanism. Notably, the improvement on precision@30% suggests that $5\%\sim 15\%$ of the mistakenly rejected papers can be rectified to acceptance through the proposed score calibration method. Meanwhile, it is also intriguing to understand the degrading performance on precision@1.5%. This pattern is consistent in the past three years of ICLR. The potential reason may be twofold. On the one hand, the best papers may require more community critics rather than the authors’ self-evaluation, making the Isotonic Mechanism less applicable. On the other hand, it may attribute to the limitation of our data generation process, as we simulate the paper quality purely based on its review score, which is not necessarily a good indicator of a top paper being selected as “oral”.
Figure 5: The loss of Isotonic Mechanism under different levels of owners’ perception noise with varying number of owners per partition block and different review noise $\sigma=$ 1 (left), 2 (middle), 4 (right) on a $7$-level binary tree (top) and $11$-level ternary tree (bottom).
Moreover, we seek to understand the structural influence of partition on the informativeness and robustness of the Isotonic Mechanism. An immediate observation is that the Isotonic Mechanism under the greedy partition outperforms that under the random partition, which matches our hypothesis in the theoretical analysis in Section 4. We also examine the empirical tradeoff of robustness and informativeness through a simulation of hierarchical ownership, given that the authors have a certain level of perception noises on their item qualities. We construct a $11$-level binary tree with $2^{11}$ items and $12^{11}-$ 1 owners, as well as a $7$-level ternary tree with $3^{7}$ items and $(3^{7}-1)/2$ owners. For the intuition, the partition at one extreme is to have only a single block of all items owned by the root node (owner), and there is only one owner we can elicit from. At the other extreme, the partition is to have one block for the set of items owned by each leaf node, and we can elicit from the owners at this leaf node and all its ancestor nodes. More generally, there exist $L$-strong partitions with $L$ from 1 to 11 in the binary tree, or from 1 to 7 in the ternary tree. In Figure 5, we can observe a clear trend of tradeoff: in the case of low noise (blue curve), the large block size is preferred to have more pairwise information, while in the case of high noise (red curve), the small block size is preferred to have more owners per block (larger $L$). In cases of medium noise (green or orange curve), the optimal block size is a careful balance of the block size and $L$-strongness. The trend showcases the importance of eliciting from multiple owners, e.g., to enforce the $L$-strong partition constraint, though the parameters shall be tailored carefully for the noise level in real-world applications.
6 Proofs
In this section, we prove Theorems 1 and 2 as well as some technical lemmas.
6.1 Proof of Theorem 1
Definition 1 (Mixed Strategy Nash equilibrium).
In a mixed strategy Nash equilibrium, each owner $j$ adopts a mixed strategy $x^{j}$ as a distribution supported on all possible rankings. We say a profile of owners’ report $\{x^{j}\}_{j=1}^{m}$ forms a mixed strategy Bayes-Nash Equilibrium (NE) under mechanism $\mathcal{M}$ if for any owner $j\in[m]$, given others’ randomized report profile $x^{-j}=\{x^{j^{\prime}}\}_{j^{\prime}\neq j}$, the expected utility under $x^{j}$ is no worse than any other possible randomized report $\widetilde{x}^{j}$, i.e.,
$\mathop{\mathbf{E}}_{\bm{y}}\left[\mathop{\mathbf{E}}_{\pi^{j}\sim x^{j}}\mathop{\mathbf{E}}_{\pi^{-j}\sim x^{-j}}\mathrm{U}^{j}\left(\widehat{\mathrm{R}}_{\mathcal{M}}(x^{j},x^{-j};\bm{y})\right)\right]\geq\mathop{\mathbf{E}}_{\bm{y}}\left[\mathop{\mathbf{E}}_{\widetilde{\pi}^{j}\sim\widetilde{x}^{j}}\mathop{\mathbf{E}}_{\pi^{-j}\sim x^{-j}}\mathrm{U}^{j}\left(\widehat{\mathrm{R}}_{\mathcal{M}}(\widetilde{x}^{j},x^{-j};\bm{y})\right)\right].$
Definition 2.
We say $\bm{a}$ majorizes $\bm{b}$, denoted $\bm{a}\succeq\bm{b}$, if $\sum_{i=1}^{k}a_{(i)}\geq\sum_{i=1}^{k}b_{(i)},\forall k<n$ and $\sum_{i=1}^{n}a_{(i)}=\sum_{i=1}^{n}b_{(i)}$, where $a_{(1)}\geq\cdots\geq a_{(n)}$ and $b_{(1)}\geq\cdots\geq b_{(n)}$ are sorted in descending order from $\bm{a}$ and $\bm{b}$, respectively.
Lemma 1 (Hardy–Littlewood–Pólya inequality).
For any vector $\bm{a},\bm{b}\in\mathbb{R}^{n}$, the inequality
$\sum_{i=1}^{n}h(a_{i})\geq\sum_{i=1}^{n}h(b_{i})$
holds for all convex functions $h$ if and only if $\bm{a}\succeq\bm{b}$.
Proof of Theorem 1.
We first show that truthful reporting forms a Bayes-Nash equilibrium. Pick any owner and let its utility function be $\mathrm{U}(\widehat{\bm{R}})=\sum_{i=1}^{n}{U}(\widehat{R}_{i})$. Suppose that the other $m-1$ owners are already truth-telling and without loss of generality we can index them from $1$ to $m-1$. With slight abuse of notation, we let $\widehat{\bm{R}}^{j}(\pi)$ be the random variable for the item quality estimates using owner $j$’s reported ranking $\pi$ w.r.t. The randomness of review noise $\bm{z}$, i.e, $\mathop{\mathbf{Pr}}[\widehat{\bm{R}}^{j}(\pi)=\widehat{\mathrm{R}}^{j}(\pi;\bm{R}+\bm{z})]=\mathop{\mathbf{Pr}}[\bm{z}]$. Suppose that the first $m-1$ estimates use the true ranking and the true ranking $\pi^{\star}$ is the identity permutation for simplicity. It is clear that the item quality estimates with all $m-1$ owners’ truthfully reported rankings are all identical,
$\widehat{\bm{R}}^{1}(\pi^{\star})=\widehat{\bm{R}}^{2}(\pi^{\star})=\cdots=\widehat{\bm{R}}^{m-1}(\pi^{\star}):=\widehat{\bm{R}}(\pi^{\star}).$
Suppose the $m$-th owner reports $\pi$. Pick any $\{\alpha^{j}\}_{j=1}^{m}$ such that $\sum_{j=1}^{m}\alpha^{j}=1$. Then, her expected utility is
$\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}(\alpha^{m}\widehat{\bm{R}}^{m}(\pi)+\sum_{j=1}^{m-1}\alpha^{j}\widehat{\bm{R}}^{j}(\pi^{\star}))\right]=\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}(\alpha^{m}\widehat{\bm{R}}^{m}(\pi)+(1-\alpha^{m})\widehat{\bm{R}}(\pi^{\star}))\right].$
Since the noise terms are exchangeable in distribution, the expectation can be replaced by averaging the term above with $n!$ permuted noise terms. Let $\widehat{\bm{R}}(\pi;\bm{z}):=\widehat{\bm{R}}(\pi;\bm{R}+\bm{z})$ and $\rho\circ\bm{z}$ be the noise vector $\bm{z}$ after permutation $\rho$. By the linearity of expectation, we have
$\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}(\alpha^{m}\widehat{\bm{R}}^{m}(\pi)+(1-\alpha^{m})\widehat{\bm{R}}(\pi^{\star}))\right]=\mathop{\mathbf{E}}_{\bm{z}}\left[\sum_{\rho}\mathrm{U}\left(\alpha^{m}\widehat{\mathrm{R}}^{m}(\pi;\rho\circ\bm{z})+(1-\alpha^{m})\widehat{\mathrm{R}}(\pi^{\star};\rho\circ\bm{z})\right)\right].$
Note that $\widehat{\mathrm{R}}^{m}(\pi;\rho\circ\bm{z})=\widehat{\mathrm{R}}(\pi;\rho\circ\bm{z})$ for any $\pi$. Then, it suffices to show that, for any vector $\bm{z}\in\mathbb{R}^{n}$,
$\sum_{\rho}\mathrm{U}\left(\alpha^{m}\widehat{\mathrm{R}}(\pi;\rho\circ\bm{z})+(1-\alpha^{m})\widehat{\mathrm{R}}(\pi^{\star};\rho\circ\bm{z})\right)\leq\sum_{\rho}\mathrm{U}\left(\alpha^{m}\widehat{\mathrm{R}}(\pi^{\star};\rho\circ\bm{z})+(1-\alpha^{m})\widehat{\mathrm{R}}(\pi^{\star};\rho\circ\bm{z})\right).$ (3)
Let $\bm{a}^{+}$ be the projection of $\bm{a}$ onto the standard isotonic cone $\{\bm{a}|a_{1}\geq a_{2}\geq\dots\geq a_{n}\}$. Recall that we assume without loss of generality that $R_{1}\geq R_{2}\dots\geq R_{n}$. Following from the coupling argument in the proof of Theorem 2 in [23], we have $\mathrm{R}(\pi^{\star};\bm{z})=(\bm{R}+\bm{z})^{+}$ and $\mathrm{R}(\pi;\bm{z})=\pi^{-1}\circ(\pi\circ\bm{R}+\pi\circ\bm{z})^{+},\forall\bm{z}\in\mathbb{R}^{n}$. So we can rewrite the inequality (3) as,
$\sum_{\rho}\mathrm{U}\left((1-\alpha^{m})(\bm{R}+\rho\circ\bm{z})^{+}+\alpha^{m}\pi^{-1}\circ(\pi\circ\bm{R}+\pi\circ\rho\circ\bm{z})^{+}\right)\leq\sum_{\rho}\mathrm{U}\left((\bm{R}+\rho\circ\bm{z})^{+}\right).$
We can prove this inequality with the following steps:
$\sum_{\rho}\mathrm{U}\left((1-\alpha^{m})(\bm{R}+\rho\circ\bm{z})^{+}+\alpha^{m}\pi^{-1}\circ(\pi\circ\bm{R}+\pi\circ\rho\circ\bm{z})^{+}\right)$
$\leq$ $\sum_{\rho}\mathrm{U}\left((1-\alpha^{m})(\bm{R}+\rho\circ\bm{z})^{+}+\alpha^{m}(\pi\circ\bm{R}+\pi\circ\rho\circ\bm{z})^{+}\right)$
$\leq$ $\sum_{\rho}\mathrm{U}\left((1-\alpha^{m})(\bm{R}+\rho\circ\bm{z})^{+}+\alpha^{m}(\bm{R}+\pi\circ\rho\circ\bm{z})^{+}\right)$
$=$ $\sum_{i=1}^{n}\sum_{\rho}U\left((1-\alpha^{m})(\bm{R}+\rho\circ\bm{z})^{+}_{i}+\alpha^{m}(\bm{R}+\pi\circ\rho\circ\bm{z})^{+}_{i}\right)$
$\leq$ $\sum_{i=1}^{n}\sum_{\rho}U\left((1-\alpha^{m})(\bm{R}+\rho\circ\bm{z})^{+}_{i}+\alpha^{m}(\bm{R}+\rho\circ\bm{z})^{+}_{i}\right)$
$=$ $\sum_{\rho}\mathrm{U}\left((\bm{R}+\rho\circ\bm{z})^{+}\right)$
The first inequality is due to Lemmas 1 and 2, as $\pi^{-1}$ is some permutation on $\bm{b}^{+}=(\pi\circ\bm{R}+\pi\circ\rho\circ\bm{z})^{+}$. The second inequality follows from Lemmas 1, 3 as well as Lemma 6.5 of [23], which shows that $(\bm{R}+\pi\circ\rho\circ\bm{z})^{+}\succeq(\pi\circ\bm{R}+\pi\circ\rho\circ\bm{z})^{+},\forall\rho$. The third inequality follows from Lemma 1 and 2 where for each distinct permutation $\rho_{j}$, $(1-\alpha^{m})(\bm{R}+\rho_{j}\circ\bm{z})^{+}_{i},\alpha^{m}(\bm{R}+\rho_{j}\circ\bm{z})^{+}_{i}$ are the $j$-th entry of $\bm{a}^{+},\bm{b}^{+}\in\mathbb{R}^{n!}$ and $\pi$ is some permutation on $\bm{b}^{+}$. The equalities are by the definition of the utility function and its linearity.
It now remains to prove that truthful NE gives every owner the highest equilibrium utility among all possible equilibria of the game. Again, we pick arbitrary owner $j$. Under truthful NE, the owner’s utility is $\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}^{j}(\widehat{\bm{R}}(\pi^{\star}))\right]$, where $\widehat{\bm{R}}(\pi^{\star})$ is the item quality estimate under the truthful ranking. We compare it to any (possibly mixed) Nash equilibrium (see Definition 1). Consider any pure strategy profile in this mixed Nash equilibrium, $\bm{\pi}=(\pi^{1},\pi^{2},\dots,\pi^{m})\sim\bm{x}=(x^{1},x^{2},\dots,x^{m})$. This sampled profile of reported ranking corresponds to the weighted average of quality estimates, $\sum_{j=1}^{m}\alpha^{j}\widehat{\bm{R}}(\pi^{j})$. We can see that the owner’s expected utility under this pure strategy profile is no larger than that under the truthful strategy profile,
$\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}^{j}\left(\sum_{j=1}^{m}\alpha^{j}\widehat{\bm{R}}(\pi^{j})\right)\right]$ $\leq\sum_{j=1}^{m}\alpha^{j}\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}^{j}(\widehat{\bm{R}}(\pi^{j}))\right]$ (4)
$\leq\mathop{max}_{j\in[m]}\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}(\widehat{\bm{R}}(\pi^{j}))\right]$
$\leq\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}^{j}(\widehat{\bm{R}}(\pi^{\star}))\right],$
where the first inequality is due to Jensen’s inequality and linearity of expectation, the second inequality is by the property of arithmetic mean, last inequality is due to Theorem 6.1. Hence, the owner’s expected utility under this Nash equilibrium, taking an expectation over the distribution of pure strategy profile, $\mathop{\mathbf{E}}_{\bm{\pi}\sim\bm{x}}\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}^{j}(\sum_{j=1}^{m}\alpha^{j}\widehat{\bm{R}}^{j})\right]\leq\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}^{j}(\widehat{\bm{R}}(\pi^{\star}))\right]$, is no larger than the expected utility under the truthful NE. ∎
Remark 2.3.
In fact, one may have noticed that the proof of payoff dominance above does not rely on whether the non-truthful strategy profile forms an equilibrium or not. Hence, a slightly stronger result holds, that is, the truthful NE gives every owner the highest utility among all possible strategy profiles of the game. However, we would like to also point out that this additional observation does not imply that the truthful report is necessarily a dominant strategy for any owner, i.e., the best response to any other strategy profile of the other owners. Nor is the truthful NE necessarily unique in this game. That said, given its payoff-dominant property, there is other strong evidence, especially from empirical human preference and behavioral theory, to expect the truthful equilibrium outcome, as pointed out in Remark 1.1.
Theorem 6.1 (Restatement of Theorem 2 in [23]).
Under Assumption 2.1, 2.2, 2.3, it is optimal for a single owner to truthfully report the ranking of its items $\pi^{\star}$,
$\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}(\widehat{\bm{R}}(\pi^{\star}))\right]=\mathop{max}_{\pi}\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}(\widehat{\bm{R}}(\pi))\right].$
Lemma 2.
For any vector $\bm{a},\bm{b}\in\mathbb{R}^{n}$, $\bm{a}^{+}+\bm{b}^{+}\succeq\bm{a}^{+}+\rho\circ\bm{b}^{+}$ for any permutation $\rho$ on $\bm{b}^{+}$.
Proof of Lemma 2.
We prove it by contradiction. Suppose $\rho$ maximizes the sum. Let $\bm{b}^{\prime}=\rho\circ\bm{b}^{+}$. We have for some $i<j$, $b^{\prime}_{i}<b^{\prime}_{j}$. In this case, consider another vector $\widetilde{\bm{b}}=\widetilde{\rho}\circ\bm{b}^{+}$ with $\widetilde{b}_{j}=b^{\prime}_{i},\widetilde{b}_{i}=b^{\prime}_{j}$ and $\widetilde{b}_{\ell}=b^{\prime}_{\ell},\forall\ell\neq i,j$. Given that $a_{i}^{+}>a_{j}^{+}$, we have $a_{i}^{+}+b^{\prime}_{j}>a_{j}^{+}+b^{\prime}_{i}$. Hence, $\bm{a}^{+}+\widetilde{\rho}\circ\bm{b}^{+}\succeq\bm{a}^{+}+\rho\circ\bm{b}^{+}$, which reaches the contradiction. ∎
Lemma 3.
Suppose $\bm{a},\bm{a}^{\prime},\bm{b}$ are in the same descending order and $\bm{a}\succeq\bm{a}^{\prime}$, then $\bm{a}+\bm{b}\succeq\bm{a}^{\prime}+\bm{b}.$
Proof of Lemma 3.
We first verify that equality holds. Since $\sum_{i=1}^{n}a_{i}=\sum_{i=1}^{n}a_{i}^{\prime}$,
$\sum_{i=1}^{n}(a_{i}+b_{i})=\sum_{i=1}^{n}(a^{\prime}_{i}+b_{i}).$
For the inequality at any $k<n$, we have $\sum_{i=1}^{k}a_{i}\geq\sum_{i=1}^{k}a^{\prime}_{i}$. This readily implies that
$\sum_{i=1}^{k}(a_{i}+b_{i})\geq\sum_{i=1}^{k}(a^{\prime}_{i}+b_{i}).$
Therefore, by definition, we have $\bm{a}+\bm{b}\succeq\bm{a}^{\prime}+\bm{b}$. ∎
6.2 Proof of Theorem 2
A Convenient Bipartite Graph View of the Ownership Relation. To ease our discussion in the formal proofs regarding the optimal partition, we start by representing the overlapping ownership model as a bipartite graph $\mathcal{G}=(\mathcal{P},\mathcal{A},\mathcal{E})$. The vertex sets $\mathcal{P},\mathcal{A},$ with $|\mathcal{P}|=n,|\mathcal{A}|=m$ correspond to the disjoint set of $n$ items (papers) and $m$ owners (authors), respectively. The edge set $\mathcal{E}\subseteq\mathcal{P}\times\mathcal{A}$ corresponds to the ownership relation: the owner $j$ owns the item $i$, if and only if there is an edge $(p_{i},a_{j})\in\mathcal{E}$ between the vertices $p_{i}\in\mathcal{P}$ and $a_{j}\in\mathcal{A}$. The number of edges $|E|=\sum_{j\in[m]}|\mathcal{I}^{j}|$ is precisely the $N$ in the theorem statement, which captures the order of the problem’s input size. Let $E=(e_{i}^{j})_{m\times n}$ be the $0-1$ bi-adjacency matrix of the ownership in which $e_{i}^{j}=1$ for each edge $(p_{i},a_{j})\in\mathcal{E}$ — as is illustrated in Table 2 and 3. Hence, the set of items owned by $j$ is exactly the neighbor set of vertex $a_{j}$, i.e., $\mathcal{I}^{j}=\{p_{i}\in\mathcal{P}|e_{i}^{j}=1\}$.
We now formalize the partition-based scheme in Mechanism 2 using the bipartite graph characterization. Specifically, we consider the partition ${\mathcal{S}}=\{{\mathcal{S}}_{1},\ldots,{\mathcal{S}}_{K}\}$ on the vertex set $\mathcal{P}$ of the bipartite graph. By definition of a partition, we have $\bigcup_{k=1}^{K}{\mathcal{S}}_{k}=\mathcal{P}$ and ${\mathcal{S}}_{k^{\prime}}\cap{\mathcal{S}}_{k}=\varnothing,\forall k\neq k^{\prime}\in[K]$. For each partition block ${\mathcal{S}}_{k}$, we denote ${\mathcal{T}}_{k}=\{j\in\mathcal{A}|{\mathcal{S}}_{k}\subseteq\mathcal{I}^{j}\}$ as the set of owners who owns all items in ${\mathcal{S}}_{k}$. For the blocks with $|{\mathcal{S}}_{k}|\leq 1$ or $|{\mathcal{T}}_{k}|\leq 0$, the mechanism simply uses the raw review score. Otherwise, by construction, the sub-bipartite-graph of $\mathcal{G}$ induced by each $({\mathcal{S}}_{k},{\mathcal{T}}_{k})$ and their incidence edges $\{(p_{i},a_{j})\in\mathcal{E}|p_{i}\in{\mathcal{S}}_{k},a_{j}\in{\mathcal{T}}_{k}\}$ forms a complete bipartite graph, i.e., a complete overlapping ownership. Therefore, we can also visualize the partition of the partially overlapping ownership using the bi-adjacency matrix illustrated in Table 3.
We now utilize the notations above to formally prove Theorem 2.
Proof of Theorem 2.
Pick any ownership instance represented as a bipartite graph $\mathcal{G}$ and any weight function $w\in\mathcal{W}$. Let $\operatorname{OPT}(\mathcal{G}):=\operatorname{obj}({\mathcal{S}}^{*})$, where ${\mathcal{S}}^{*}=\{{\mathcal{S}}_{k}^{*}\}_{k=1}^{K}$ is the optimal $1$-strong partition of $\mathcal{G}$. Consider the greedy algorithm that keeps picking the largest residual item set owned by some author, until all items are assigned to some block, as described in Algorithm 3. Let ${\mathcal{S}}$ denote the partition obtained by Algorithm 3. By construction, ${\mathcal{S}}$ is 1-strong — i.e., there is at least one owner who owns all items of each item subset in ${\mathcal{S}}$. We let $\operatorname{ALG}(\mathcal{G}):=\operatorname{obj}({\mathcal{S}})$ be the objective value of the partition obtained by Algorithm 3. The remainder of the proof will argue that the output ${\mathcal{S}}$ simultaneously satisfies,
$\frac{\operatorname{ALG}(\mathcal{G})}{\operatorname{OPT}(\mathcal{G})}=\frac{\operatorname{obj}({\mathcal{S}})}{\operatorname{obj}({\mathcal{S}}^{*})}\geq\inf\big{\{}\frac{w(x)}{w^{\prime}_{-}(x)x}\big{|}w^{\prime}_{-}(x)>0,x\geq 2\big{\}},$
for every $w\in\mathcal{W}$.
Let $\mathcal{G}\setminus j$ denote the induced subgraph of $\mathcal{G}$ by removing the vertex $j$ as well as all its adjacent item nodes and corresponding edges from $\mathcal{G}$. Formally, in this induced subgraph $\mathcal{G}\setminus j=(\mathcal{P}^{\prime},\mathcal{A}^{\prime},\mathcal{E}^{\prime})$, we have $\mathcal{P}^{\prime}=\mathcal{P}\setminus\mathcal{I}^{j},\mathcal{A}^{\prime}=\mathcal{A}\setminus\{j\}$ and =$\mathcal{E}^{\prime}=\{(i,j)|(i,j)\in\mathcal{E}\text{ and }i\in\mathcal{P}^{\prime},j\in\mathcal{A}^{\prime}\}$.
We now prove the approximation ratio of the greedy algorithm by induction on the owner set’s size $m=|\mathcal{A}|$. For the base case, for any bipartite graph $\mathcal{G}=(\mathcal{P},\mathcal{A},\mathcal{E})$ with $|\mathcal{A}|=1$, it is clear that the greedy approach finds the exact optimal partition. For the inductive case, we show that if the approximation guarantee holds for any bipartite graph $\mathcal{G}=(\mathcal{P},\mathcal{A},\mathcal{E})$ with $|\mathcal{A}|\leq k$, then it must hold for any bipartite graph $\mathcal{G}$ with $|\mathcal{A}|=k+1$ as well.
Let $j$ be the very first owner picked by Algorithm 3 on the bipartite graph instance $\mathcal{G}$ with $|\mathcal{A}|=k+1$. Thus, the first block (i.e., item set) picked by the algorithm must be $\mathcal{I}^{j}$. Let ${\mathcal{S}}^{*}$ be the optimal partition of $\mathcal{G}$ under weight $w$. Note that the algorithm output ${\mathcal{S}}$ is independent of $w$, but ${\mathcal{S}}^{*}$ typically depends on $w$. For each block ${\mathcal{S}}_{k}^{*}\in{\mathcal{S}}^{*}$, let $p_{k}^{*}:=\left|{\mathcal{S}}_{k}^{*}\right|$ be the number of items in this block and $a_{k}:=\left|{\mathcal{S}}^{*}_{k}\cap\mathcal{I}^{j}\right|$ be the number of items covered by the greedy algorithm’s first choice $\mathcal{I}^{j}$. Hence, $\left|\mathcal{I}^{j}\right|=\sum_{k=1}^{K}a_{k}$. We can then bound the approximation ratio of the greedy algorithm by decomposing the objective value as follows,
$\frac{\operatorname{ALG}(\mathcal{G})}{\operatorname{OPT}(\mathcal{G})}=$ $\frac{w(\mathcal{I}^{j})+\operatorname{ALG}(\mathcal{G}\setminus j)}{\sum_{k=1}^{K}w(p^{*}_{k})-\sum_{k=1}^{K}w(p^{*}_{k}-a_{k})+\sum_{k=1}^{K}w(p^{*}_{k}-a_{k})}$
$\geq$ $\frac{w(\mathcal{I}^{j})+\operatorname{ALG}(\mathcal{G}\setminus j)}{\sum_{k=1}^{K}w(p^{*}_{k})-\sum_{k=1}^{K}w(p^{*}_{k}-a_{k})+\operatorname{OPT}(\mathcal{G}\setminus j)},$
where the inequality uses the suboptimality of the partition $\{{\mathcal{S}}^{*}_{k}\setminus\mathcal{I}^{j}\}_{k=1}^{K}$ for the graph $\mathcal{G}\setminus j$.
We will bound the approximation ratio by considering the two parts $\frac{w\left(\sum_{k=1}^{K}a_{k}\right)}{\sum_{k=1}^{K}w(p^{*}_{k})-\sum_{k=1}^{K}w(p^{*}_{k}-a_{k})}$ and $\frac{\operatorname{ALG}(\mathcal{G}\setminus j)}{\operatorname{OPT}(\mathcal{G}\setminus j)}$ separately, and then apply the mediant inequality to derive the lower bound. The second part can be easily bounded according to the induction hypothesis. However, the first part could be unbounded in general (this may happen when $p^{*}_{k}$’s are all at most 1). Fortunately, the analysis for such special case can be carried out separately. Specifically, if $\mathop{max}_{k\in[K]}p^{*}_{k}\leq 1$ in $\mathcal{G}$, by convexity of $w$ function, this implies that any 1-strong partition of the items are at least as good as ${\mathcal{S}}^{*}$ and thus the greedy algorithm is optimal.
Next, we consider the more typical situation with $p:=\mathop{max}_{k\in[K]}p^{*}_{k}\geq 2$. Here, invoking the mediant inequality and a technical Lemma 4 below with $p\geq 2$, we can lower bound the approximation ratio as follows,
$\frac{w(\mathcal{I}^{j})+\operatorname{ALG}(\mathcal{G}\setminus j)}{\sum_{k=1}^{K}w(p^{*}_{k})-\sum_{k=1}^{K}w(p^{*}_{k}-a_{k})+\operatorname{OPT}(\mathcal{G}\setminus j)}$
$\geq$ $\mathop{min}\bigg{\{}\frac{w\left(\sum_{k=1}^{K}a_{k}\right)}{\sum_{k=1}^{K}w(p^{*}_{k})-\sum_{k=1}^{K}w(p^{*}_{k}-a_{k})},\ \frac{\operatorname{ALG}(\mathcal{G}\setminus j)}{\operatorname{OPT}(\mathcal{G}\setminus j)}\bigg{\}}$
$\geq$ $\inf\left\{\frac{w(x)}{w^{\prime}_{-}(x)x}\big{|}w^{\prime}_{-}(x)>0,x\geq 2\right\}.$
Finally, for running time analysis, Algorithm 3 can be efficiently implemented in almost linear time using appropriately chosen data structures. Specifically, we employ the red-black tree data structure which can retrieve the maximum element $j^{*}=\arg\mathop{max}_{j\in[m]}|\mathcal{I}^{j}\setminus\overline{\mathcal{I}}|$, i.e., the maximal residual paper sets among all authors, using $O(\log m)$ time. We will then update the paper sets of all the authors who have at least one paper in $\mathcal{I}^{j^{*}}\setminus\overline{\mathcal{I}}$. The number of operations is at most the number of edges connected to ${\mathcal{S}}^{j^{*}}$ in the bipartite graph. We then update the red-black tree structure using the new sizes of the residual paper sets of the involved papers. Note that the update of each author’s paper set size takes $O(\log m)$ time by first deleting it in the red-black tree and then re-inserting it. The above recursion takes at most $m$ rounds since each round at least one author’s residual paper set will become empty and there are only $m$ authors. In total, the total running time is thus $O(|\mathcal{E}|\log m)$.
Lemma 4.
Consider any $\{p_{k}^{*}\}_{k\in[K]}$ and $\{a_{k}\}_{k\in[K]}$ such that $p^{*}_{k}\geq a_{k}\geq 0$ for any $k\in[K]$ and $\sum_{k=1}^{K}a_{k}\geq p:=\mathop{max}_{k\in[K]}p^{*}_{k}$. For any $w\in\mathcal{W}$ with $w(p)>0$, we have
$\frac{w\left(\sum_{k=1}^{K}a_{k}\right)}{\sum_{k=1}^{K}w(p^{*}_{k})-\sum_{k=1}^{K}w(p^{*}_{k}-a_{k})}\geq\inf_{x\geq p}\frac{w(x)}{w^{\prime}_{-}(x)x}$
Proof of Lemma 4.
Any $w\in\mathcal{W}$ is convex, so its left derivative $w^{\prime}_{-}$ is non-deceasing such that $w(p^{*}_{k})-w(p^{*}_{k}-a_{k})\leq w^{\prime}_{-}(p_{k}^{*})a_{k}$. In addition, $w^{\prime}_{-}(p^{*}_{k})\geq 0$ since $w(0)=0,w(p^{*}_{k})\geq 0$. Let $\sum_{k=1}^{K}a_{k}=a$, we have
$\frac{w\left(\sum_{k=1}^{K}a_{k}\right)}{\sum_{k=1}^{K}w(p^{*}_{k})-\sum_{k=1}^{K}w(p^{*}_{k}-a_{k})}\geq\frac{w(a)}{\sum_{k=1}^{K}w^{\prime}_{-}(p^{*}_{k})a_{k}}.$
Since $a\geq\mathop{max}_{k\in[K]}p^{*}_{k}=p$ and thus $w^{\prime}_{-}(p_{k}^{*})\leq w^{\prime}_{-}(a)$ for any $k$ by convexity of $w$. Since $w(p)>0$ and $w(0)=0$, we must have $0<w^{\prime}_{-}(p)\leq w^{\prime}_{-}(x)$ for any $x\geq p$. Therefore,
$\frac{w(a)}{\sum_{k=1}^{K}w^{\prime}_{-}(p^{*}_{k})a_{k}}\geq\frac{w(a)}{\sum_{k=1}^{K}w^{\prime}_{-}(a)a_{k}}=\frac{w(a)}{w^{\prime}_{-}(a)a}\geq\inf_{x\geq p}\frac{w(x)}{w^{\prime}_{-}(x)x}.$
6.2.1 Tightness of Greedy’s Approximation Ratio in the Monomial Class
One implication of Theorem 2 is that for $\alpha$-th degree polynomial function $w=|{\mathcal{S}}_{k}|^{\alpha}$, the greedy algorithm has $\frac{1}{\alpha}$-approximation. Next, we show that this approximation ratio is provably tight for every weight in the monomial class.
Consider an instance with $n=MN$ items and $m=M+L$ owners for some positive integer $L,M,N$ such that $M^{L-1}\mid N$ and $N(1-1/M)^{L}\geq L$. The ownership relations are as follows (see also Figure 6 for the visualization of the instance):
  • There are $M$ owners $1,\cdots,M$, each of whom owns a disjoint set of $N$ items. Thus, collectively, they own all the $MN$ items. Let the items owned by these owners as $\mathcal{I}^{1},\dots,\mathcal{I}^{M}$ respectively.
  • We denote the item owned by the remaining $L$ owners as $\mathcal{I}^{\ell+M}$ for $\ell\in[1,\dots,L]$. Let $\left|\mathcal{I}^{\ell+M}\cap\mathcal{I}^{\ell}\right|=(1-1/M)^{\ell-1}N/M+1$ and $\left|\mathcal{I}^{\ell+M}\cap\mathcal{I}^{\ell^{\prime}}\right|=(1-1/M)^{\ell-1}N/M,\forall\ell^{\prime}\neq\ell\in[M]$. Hence, $|\mathcal{I}^{\ell+M}|=(1-1/M)^{\ell-1}N+1$.
Observe that $\mathcal{I}^{M+1}$ is the largest item set and $\mathcal{I}^{M+2}$ becomes the largest item set after $\mathcal{I}^{M+1}$ is removed. Therefore, we can see that the greedy algorithm proceeds by picking the $\ell$-th owner of the last $L$ owners from $\ell\in[1,\dots,L]$ and lastly the first $M$ owners, i.e., when each of whom has remaining item less than $N-\sum_{\ell=1}^{L}(1-1/M)^{\ell-1}N/M=(1-1/M)^{L}N$. As such, the greedy partition achieves the objective of at least $\operatorname{ALG}(\mathcal{G})\leq\sum_{\ell=1}^{L}[(1-1/M)^{\ell-1}N+1]^{\alpha}+M(1-1/M)^{\alpha L}N^{\alpha}.$
On the other hand, consider a partition strategy that only picks the first $M$ owners, which achieves the objective $\operatorname{OPT}(\mathcal{G})=MN^{\alpha}$. Simply using it as a lower bound for the optimal objective, we can bound the approximation ratio as,
$\frac{\operatorname{ALG}(\mathcal{G})}{\operatorname{OPT}(\mathcal{G})}$ $\leq\frac{\sum_{\ell=1}^{L}[(1-1/M)^{\ell-1}N+1]^{\alpha}+M(1-1/M)^{\alpha L}N^{\alpha}}{MN^{\alpha}}$
$=\frac{\sum_{\ell=1}^{L}[(1-1/M)^{\ell-1}N+1]^{\alpha}}{MN^{\alpha}}+(1-1/M)^{\alpha L}.$
For any constant $M>1$, we have as $N,L\to\infty$,
$\lim_{N,L\to\infty}\frac{\sum_{\ell=1}^{L}[(1-1/M)^{\ell-1}N+1]^{\alpha}}{MN^{\alpha}}+(1-1/M)^{\alpha L}$ $=\lim_{N,L\to\infty}\sum_{\ell=1}^{L}(1-1/M)^{(\ell-1)\alpha}/M$
$=\frac{1/M}{1-(1-1/M)^{\alpha}}.$
Finally, as $M\to\infty$, the ratio matches with the worst case approximation guarantee $1/\alpha$,
$\lim_{M\to\infty}\frac{1/M}{1-(1-1/M)^{\alpha}}=\lim_{M\to\infty}\frac{1}{\alpha-1/M}=1/\alpha.$
every picture/.style=line width=0.75pt
[x=0.75pt,y=0.75pt,yscale=-1,xscale=1]
[…]
Figure 6: Illustration of the worst case instance for the greedy algorithm. $l=1,\cdots,L$.
Meanwhile, despite the hardness result for the size-based objective in Proposition 2, we would like to point out that the efficient algorithm for optimal partition under the $\alpha$-th degree polynomial objective may exist. In particular, consider the following instance with $m=n+1$, where the optimal partition cannot realize the set-cover problem for the hardness reduction. The owner $i$ owns the items $\{2i-1,2i\}$,$\forall i\in[n]$ and owner $n+1$ owns the items $\{1,3,\cdots,2n-1\}$. In this case, the optimal partition is not the minimum set cover, since
$\operatorname{obj}\left(\left\{\{1,3,\cdots,2n-1\},\{2\},\cdots\{2n\}\right\}\right)=n^{\alpha}+n>\operatorname{obj}\left(\left\{\{1,2\},\cdots,\{2n-1,2n\}\right\}\right)=n\times 2^{\alpha}.$
Hence, it remains an open question on the hardness of finding the optimal partition under the $\alpha$-th degree polynomial objective.
7 Discussion
In this work, we have established a generalized Isotonic Mechanism designed to enhance peer review in machine learning conferences, specifically addressing the practical context where multiple authors contribute to various papers. We demonstrate the truthfulness of our mechanism and design a simple polynomial-time algorithm to partition the ownership sets, which achieves a constant ratio approximation of the optimal partition that is computationally hard to find. This facilitates the practical implementation of our mechanism.
Nonetheless, our mechanism is not without limitations, which present opportunities for future investigations. A natural extension lies in adapting the mechanism to accommodate review scores drawn from exponential family distributions [26]. Additionally, the introduction of randomness into authors’ private information presents an intriguing research direction. This modeling perspective could more accurately reflect the reality that authors may assign different values to their co-authored submissions. Lastly, recognizing the potential heterogeneity of the utility function poses a significant, yet crucial challenge. In this context, an author’s value of a paper could depend on factors such as authorship order, the length of the author’s list, and her relative contributions to the paper [7].
Acknowledgments
This work was supported in part by NSF through CCF-2132506 and CCF1934876, ARO through W911NF-23-1-0030, Analytics at Wharton, and Wharton AI and Analytics for Business.
Appendix A Appendix
Proof of Proposition 1.
We prove that for any ownership instance $\mathcal{O}=\{\mathcal{I}^{j}\}_{j\in[m]}$, we can construct a different ownership instance $\mathcal{O}^{\prime}=\{\overline{\mathcal{I}}^{j}\}_{j\in[m]}$ such that any $L$-strong partition ${\mathcal{S}}$ in $\mathcal{O}$ is a 1-strong partition in $\mathcal{O}^{\prime}$.
Pick any bipartite graph $\mathcal{G}=(\mathcal{A},\mathcal{P},\mathcal{E})$ from the ownership instance $\mathcal{O}=\{\mathcal{I}^{j}\}_{j\in[m]}$. Suppose we were to find $L$-strong partition ${\mathcal{S}}$ in $\mathcal{G}$. We can construct a bipartite graph $\mathcal{G}^{\prime}=(\mathcal{A}^{\prime},\mathcal{P}^{\prime},\mathcal{E}^{\prime})$, where the item set remains the same $\mathcal{P}^{\prime}=\mathcal{P}$, the owner set $\mathcal{A}^{\prime}$ corresponds to all the $L$ element subsets of $\mathcal{A}$, $\mathcal{A}^{\prime}=\{\ell|u^{\ell}\subseteq\mathcal{A},|u^{\ell}|=L\}$ and the edge set captures the commons items of every $L$ owners, =$\mathcal{E}^{\prime}=\{(i,\ell)|i\in\bigcap_{j\in u^{\ell}}\mathcal{I}^{j}\}$. We show that for any $L$-strong partition ${\mathcal{S}}=\{{\mathcal{S}}_{1},{\mathcal{S}}_{2},\dots,{\mathcal{S}}_{K}\}$ of $\mathcal{G}$, there is an equivalent $1$-strong greedy partition ${\mathcal{S}}^{\prime}=\{{\mathcal{S}}^{\prime}_{1},{\mathcal{S}}^{\prime}_{2},\dots,{\mathcal{S}}^{\prime}_{K}\}$ of $\mathcal{G}^{\prime}$ such that ${\mathcal{S}}_{k}={\mathcal{S}}^{\prime}_{k},\forall k\in K$.
The proof is a straightforward verification of this bipartite graph construction. Pick any ${\mathcal{S}}_{k}\in{\mathcal{S}}$ with $|{\mathcal{T}}_{k}|\geq L$ in $\mathcal{G}$, there exists an ${\mathcal{S}}^{\prime}_{k}$ with $|{\mathcal{T}}^{\prime}_{k}|=1$ in $\mathcal{G}^{\prime}$. That is, we have ${\mathcal{T}}^{\prime}_{k}=\{\ell\}$ for some $u^{\ell}\subseteq{\mathcal{T}}_{k}$, then ${\mathcal{S}}^{\prime}_{k}={\mathcal{S}}_{k}\subseteq\bigcap_{j\in u^{\ell}}\mathcal{I}^{j}$ by construction. Therefore, ${\mathcal{S}}_{k}={\mathcal{S}}^{\prime}_{k},\forall k\in[K]$. ∎
Proof of Proposition 2.
We consider the weight function $w(x)=\mathop{max}\{x,1\}-1$. For any partition with $K$ blocks, the objective function is $\operatorname{obj}({\mathcal{S}})=n-K$. We show that finding the optimal partition under such an objective function is at least as hard as finding the minimum set. We prove via a standard hardness reduction argument: for any set cover instance, we can construct a bipartite graph instance in which the partition that maximizes the objective function is $\operatorname{obj}({\mathcal{S}})=n-K$ can be turned into a minimum set cover.
Pick arbitrary set cover instance with $m$ subsets of $n$ item, $\mathcal{V}_{1},\mathcal{V}_{2},\dots,\mathcal{V}_{m}$. We construct a bipartite graph $\mathcal{G}=(\mathcal{A},\mathcal{P},\mathcal{E})$ where $\mathcal{A}=[m],\mathcal{P}=[n]$ and =$\mathcal{E}=\{(p_{i},a_{j})|p_{i}\in\mathcal{V}_{j},j\in[m]\}$. There exists an optimal partition ${\mathcal{S}}^{*}$ with $\operatorname{obj}({\mathcal{S}}^{*})=n-K^{*}$, if and only if the minimum set cover has size $K^{*}$. For the “if” direction, we construct a partition ${\mathcal{S}}^{*}$ with $\operatorname{obj}({\mathcal{S}}^{*})=n-K^{*}$ from the set cover of size $K^{*}$. For each subset selected by the set cover, we construct a partition block with items in the subset minus the items already in the partition. This partition must be valid and have $K^{*}$ blocks. For the “only if” direction, we construct the set cover of size $K^{*}$ from the optimal partition ${\mathcal{S}}^{*}$. For every block ${\mathcal{S}}_{k}^{*}$ of the optimal partition ${\mathcal{S}}^{*}$, we pick any vertex $a_{j}\in{\mathcal{T}}_{k}^{*}$ from the corresponding owner set. Based on the $*K*$ distinct vertices, we select the subsets with the same indices, and it forms a set cover of all $n$ items by construction. ∎