### v3

# An Isotonic Mechanism for Overlapping Ownership

Abstract
Motivated by the problem of improving peer review at large scientific conferences, this paper studies how to elicit self-evaluations to improve review scores in a natural many-to-many owner-item (e.g., author-paper) situation with overlapping ownership. We design a simple, efficient and truthful mechanism to elicit self-evaluations from item owners that can be used to calibrate their noisy review scores in the existing evaluation process (e.g., papers’ review scores from peers).
Our approach starts by partitioning the owner-item relation structure into disjoint blocks, each sharing a common set of co-owners. We then elicit the ranking of items from each owner and employ isotonic regression to produce adjusted item scores, aligning with both the reported rankings and raw item review scores. We prove that truth-telling by all owners is a payoff dominant Nash equilibrium for any valid partition of the overlapping ownership sets under natural conditions. Moreover, the truthfulness depends on eliciting rankings independently within each block, making block partition optimization crucial for improving statistical efficiency. Despite being computationally intractable in general, we develop a nearly linear-time greedy algorithm that provably finds a performant block partition with appealing robust approximation guarantees. Extensive experiments on both synthetic data and real-world conference review data demonstrate the effectiveness of our mechanism in a pressing real-world problem.
1 Introduction
In recent years, major AI and machine learning conferences such as NeurIPS, ICML, and ICLR have faced a concerning decline in the quality of peer review, posing a significant challenge to the global machine learning community [21, 6, 39, 24]. In particular, the NeurIPS 2014 experiment showed that $49.5\%$ of the papers accepted by one committee would be rejected by another [22, 11, 21]. This inconsistency probability was $50.6\%$ for NeurIPS 2021 [9]. This troubling trend can be largely attributed to a structural imbalance between the surge in submission volumes and lagged growth of the number of qualified reviewers [32]. As demonstrated by Figure 1, many of these conferences have handled around 10,000 full paper submissions recently, reflecting an almost ten-fold growth over the past decade, whereas a large portion of the reviewers are inexperienced researchers (e.g., junior graduate students) [33, 34, 31]. Similar trends are seen across various research communities [27, 20, 14, 8], highlighting a systemic issue in the today’s peer review systems. To mitigate this issue, there have been progressive research efforts to improve peer review processes. Some approaches employ machine learning and optimization techniques to improve reviewer assignments, reduce bias, and automate review procedures, while others adopt an economic approach modeling the incentives of participants to encourage high-quality reviews (see, e.g., a survey by [32]).
Figure 1: The number of paper submissions in major machine learning and related conferences from 2014 to 2024.
Motivated by pressing concerns as mentioned above, this paper proposes a mechanism that blends together statistical and economic design, and can effectively incentivize authors to share evaluations of their own papers to calibrate peer review scores. Conceptually, the motivation is natural: if the reviewer pool cannot supply sufficient information for all the submissions, conference organizers could resort to other sources — in our case, the authors themselves. However, while self-criticism is a merit in science, self-evaluation has been largely overlooked due to the obvious conflicts of interest. That is, authors may be reluctant to provide honest assessments, if revealing a mediocre rating of their own work puts them at disadvantage. Consequently, estimations from manipulated data may be hardly useful. This presents a fundamental design challenge:
How can statistical estimation processes effectively elicit data from strategic actors?
In particular, is it possible to balance between statistical efficiency and incentive compatibility?
Examining the above question within a clean and stylized model of overlapping ownership, this paper embarks on a path to bring the mechanism design perspective to statistical estimation and proposes an approach, termed the Isotonic Mechanism, that demonstrates a definitive solution to this design challenge. In particular, while most of the learning algorithm designs have focused on improving the statistical efficiency, this work takes one step forward to further trace back the source of data. Specifically, through a careful balance of statistical efficiency and incentive compatibility, we design a mechanism that enables the elicitation and use of information traditionally deemed inaccessible. We view the demonstration of this feasibility in a high-stakes application as an important conceptual contribution, which we hope to spark a broader research agenda at the interplay of statistical and incentive design. Furthermore, given the prevalence of many-to-many owner-item relationships with overlapping ownership structures in real-world settings, we anticipate that our framework can be similarly applied to refine evaluation and award selection across a variety of domains — including internal team projects within large organizations, collaboratively produced artworks, and properties co-listed by multiple real estate agents.
To elicit an owner’s evaluation about her owned items, a key challenge to begin with, even putting incentive issues aside, is to determine what kind of data could be elicited from owners with reasonable accuracy. One approach might be to ask for highly fine-grained data, such as the owners’ own evaluation values (called item scores henceforth) for their items. However, even if owners are completely honest, they may lack the precise knowledge needed to provide reliable scores. Alternatively, one could request more general information (e.g., asking owners to identify their favorite item), but this may not offer sufficient value for improving review scores. The challenge lies in finding an effective middle ground—eliciting data that is both accurate and useful. To this end, we propose focusing on the ranking of an owner’s items (e.g., an author’s papers). Such relative information tends to be less noisy than absolute measures and has been successfully applied in various fields, such as learning from human preferences [44, 3].
Next comes our core scientific question: is it possible to truthfully elicit ranking data from owners to improve the statistical efficiency of any pre-given noisy review scores (e.g., papers’ original peer review scores)? Our starting point is an earlier work [36] which showcased the possibility in a much simplified situation with only a single owner, say, Alice. For this stylized case, Su [36] considered a natural statistical model for review score estimation and proved that truthful ranking elicitation is possible under natural owner utility assumptions. To formally describe this method, suppose Alice is the owner of $n$ items (e.g., a single author for $n$ papers). Letting $y_{1},\ldots,y_{n}$ denote any pre-given noisy review scores of the $n$ items (e.g., average reviewer scores for Alice’s single-authored $n$ papers). The mechanism elicits from Alice a ranking $\pi$ in the form of a permutation of $1,2,\cdots,n$ that sorts her items in descending order of quality, and then use this elicited ranking to compute adjusted review scores that are the solution to the following convex optimization program:
$\mathop{min}_{\bm{r}\in\mathbb{R}^{n}}$ $\sum_{i=1}^{n}(y_{i}-r_{i})^{2}$
s.t. $r_{\pi(1)}\geq r_{\pi(2)}\geq\dots\geq r_{\pi(n)}.$
This optimization yields the well-known isotonic regression [5], and its solution ensures consistency with the owner-provided ranking while remaining as close as possible to the original review scores in the least squares sense. Under natural statistical modeling of the pre-given review scores, this approach provably improves items’ score estimation. Moreover, it also guarantees truthful elicitation of Alice’s ranking, under standard utility and prior knowledge assumptions. Notably, the latter is a non-trivial property. Specifically, it crucially hinges on geometric properties of isotonic regressions, and would not hold for other estimation methods (e.g., switching from $l_{2}$-norm above to $l_{1}$-norm). This illustrates the importance of designing appropriate statistical estimation methods that can align with incentives.
Figure 2: An example of an owner-item ownership set shown as a bipartite graph. An edge between an owner and an item indicates that this individual owns the item.
This work seeks to address the more general situation where many individuals owners many items, with overlapping ownership (e.g., overlapping paper authorship). This strict generalization exhibits multiple new challenges. First, since Alice’s co-owners will also submit rankings for (some of) Alice’s items, Alice would have to account for the effect of her co-owners’ data on the utility of Alice. This intricacy is due to the fundamental difference between single-agent and multi-agent decision making setup, and it leads to our different solution concepts, generalizing from Bayesian optimal decision to Bayes Nash equilibrium. Second, an even more challenging issue is that different items can partially overlap in ownership, as illustrated in Figure 2. This is an especially common situation in today’s popular machine learning conferences. In such cases, Alice’s utility may be affected by other co-owners’ rankings on items Alice does not own. Naive adaptation of the previous Isotonic Mechanism designed for single-owner situations can lead to serious incentive issues. For instance, one natural approach is to apply the Isotonic Mechanism for every owner’s items and output final item scores by averaging their adjusted review scores. Unfortunately, it is easy to find undesirable examples with untruthful behaviors under this mechanism (see Section 3.2) because some co-owners would misreport their items’ ranking information in order to improve their own utilities. Therefore, a more principled “truth serum” is needed to effectively incorporate self-evaluations from a complex network of overlapping ownership into the reviewing mechanism.
To address the aforementioned challenges, this paper develops an approach that advances the earlier design using novel techniques from both algorithm design and mechanism design. Specifically, the mechanism first resorts to an algorithmic preprocessing procedure that partitions the owner-item ownership structure into blocks of items such that each block shares a common set of owners. It then uses the ranking information elicited from the common owners in each block to calibrate the pre-given review scores of their items. Under certain conditions, this proposed mechanism is marked by the following key characteristic:
Main Result 1.
Truthfully reporting rankings by all owners forms a Nash equilibrium under the proposed mechanism.
Along with other results, the main result of truthfulness property is formally presented in Section 3.111We make a remark on practical implementation of our mechanism. To guarantee truthfulness, it suffices for the designer to commit to only use the partial ranking information within the partitioned blocks. Thus, to ease the practical implementation, the designer can elicit every owner’s full ranking of all her items but commit to use only part of the ranking information for the mechanism. In paper review applications, the mechanism designer is typically the organizing committee of large conferences. They do have such power of commitment, since their decisions are usually audited by steering committees as well. This result significantly extends the previous truthfulness guarantee of the earlier work [36] to cases with overlapping ownership and suggests that the optimal strategy for each owner to maximize her utility is to report the ground-truth ranking, given that all the other owners do likewise. In particular, our proof of this new result involves a novel technique concerning majorization in the ordering of item scores with respect to both the pre-given reviews’ noise and other owners’ reports. Moreover, we show the mechanism encourages truthfulness behaviors in several ways. From a game-theoretical perspective, this Nash equilibrium of truthfulness is the most favored, as any other potential Nash equilibrium results in no better utility for any of the owners. From empirical studies, the mechanism induces a common-interest game and the human subjects are found to always prefer the truthful behaviors for the payoff-dominant outcome.
A potential criticism to the above partition-based approach is that we have to give up eliciting ranking information across blocks. One may naturally wonder whether the Isotonic Mechanism can be carefully tailored to elicit additional comparison information among items across partitioned blocks as such information can help address the efficiency of isotonic regression — so long as it can be truthfully elicited. Unfortunately, our second main result gives a negative answer.
Main Result 2.
Within a significantly more general class of isotonic-regression-based mechanisms (Mechanism 3), any truthful calibration mechanism has to be partition-based (for some choice of item partitions), thus cannot use any ordering information about items across the partitioned blocks.
This result is formally presented in Section 4.1. It illustrates the fundamental trade-off between the statistical efficiency and incentive requirements in designing statistical estimators using data elicited from strategic sources. Hence the only room left for us to optimize the statistical efficiency of our designed mechanism, while retaining its truthfulness, is to optimize the choice of the item partition; this leads to our third main result. As is conceivable, the choice of partition affects the statistical performance of our mechanism. For example, a partition for the instance in Figure 2 can be simply $\{p_{1},p_{2},p_{3},p_{4},p_{5}\}$, which share a single common owner $\{a_{3}\}$, or can be $\{p_{1},p_{2},p_{3}\},\{p_{4}\}$, and $\{p_{5},p_{6}\}$, which share $\{a_{1},a_{2},a_{3}\},\{a_{3}\}$, and $\{a_{3},a_{4}\}$, respectively. To find high-quality owner-item ownership partitions, we define a natural class of criteria to assess the quality of a partition. Within this class of criteria, we obtain the following result on this proposed mechanism:
Main Result 3.
A simple nearly linear-time greedy algorithm can obtain a partition of the owner-item ownership structure for the proposed mechanism that is near-optimal simultaneously for every criterion in the class.
This result is formally stated in Section 4.3. We also show that it is NP-hard to find the exactly optimal partition even under some simple criterion. In contrast, our greedy algorithm is computationally efficient while achieving constant-ratio approximation simultaneously for all the criteria in our considered class. Such simultaneous approximation is a surprisingly nice property of our problem and is a rare phenomenon generally.  [^2] We are aware of only one other situation in approximated majorization for minimizing symmetric convex objectives [13], where there is a simultaneous approximation guarantee but their optimization problem is very different from ours and their ratio is logarithmic in general. This powerful result has recently been used by Banerjee et al. [4] for the welfare guarantee in information design.
Finally, to validate the empirical performance of our mechanism, we applied it to the conference peer review application and conducted a series of experiments on both synthetic data and ICLR review data. Our experiments exhibit that the proposed mechanism enhances estimation performance relative to the original review scores, a benefit observed consistently across various experimental settings.
2 Problem Formulation
This paper considers the review score calibration problem under a network of overlapping ownership. [^3] Besides the motivating example of academic conference review, similar problems with overlapping ownership also exist in reviews for company projects, songs, movies, etc. Hence, our technical sections study an abstract model with generic terms “owner/item”. There are $m$ owners and $n$ items. The $j$-th owner owns a set of items $\mathcal{I}^{j}\subseteq[n]=\{1,2,\dots,n\}$ and let $|\mathcal{I}^{j}|=n_{j}$ [^4] The ownership relation forms a bipartite graph (see e.g., Figure 2), and we formalize their connections in our proofs in Appendix E. For each item $i\in[n]$, let $R_{i}$ be its ground-truth score and
$y_{i}=R_{i}+z_{i}$
be the raw review score given by its reviewers with noise $z_{i}$. As a friendly reminder of our notation, $i$ and $\mathcal{I}$ are used to denote item index and item set, respectively, whereas $j$ is used to index the owner. Denote by $\bm{z}=(z_{1},z_{2},\dots,z_{n}),\bm{R}=(R_{1},R_{2},\dots,R_{n})$ and $\bm{y}=(y_{1},y_{2},\dots,y_{n})$, respectively, the vector of noise, the ground-truth scores, and the raw review scores.
In applications such as conference paper reviews, the ground-truth scores $\bm{R}$ can be interpreted as the mean evaluation scores of the papers perceived across all experts in the community. This item evaluation score is mathematically well defined and always exists, [^5] We remark that paper evaluation score should be carefully distinguished from “true merit” of the paper, which is a vague concept and whose existence might be arguable. Unlike merits, a paper’s true evaluation score is well-defined, and is what peer review procedure as well as statistical methods like ours is trying to estimate. but it is very difficult to be accurately observed during the review phase for various reasons including insufficient resources to have enough reviews and insufficient time to truly observe the items’ impact. This is essentially the motivation of peer review, during which this ground-truth score will be estimated by a few experts sampled from the community. The averaged review scores of these experts form our modeling of the raw review scores $\bm{y}$ — a noisy estimation of ground-truth scores $\bm{R}$. Therefore, the high-level objective of a calibration mechanism $\mathcal{M}$ is to output the adjusted review scores $\widehat{\bm{R}}$ that provide a more accurate estimation of the ground truth scores $\bm{R}$. An advantage of such a calibration mechanism for estimating the community’s expected evaluation of papers is that it is naturally compatible with existing peer review systems.
To tackle the review score calibration problem with the limited supply of qualified reviewers, we resort to the design of owner-assisted calibration mechanisms that elicit and utilize self-evaluation data from the item owners. We abstract this mechanism design problem into the construction of an estimator $\widehat{\mathrm{R}}_{\mathcal{M}}$ for $\bm{R}$ that combines the raw scores $\bm{y}$ with the information elicited from owners $\{\pi^{j}\}_{j=1}^{m}$ to improve the estimates of $\bm{R}$. In this paper, we focus on eliciting each owner’s ranking of their items, where $\pi^{j}$ is a permutation of $1,2,\ldots,n_{j}$ that specifies an ordering of the $j$-th owner’s items in $\mathcal{I}^{j}$. The benefits of this design choice and the possible relaxations are discussed in Section 6. The owner-assisted calibration mechanism then proceeds in the following stages.
  1. 1.
    Before Review: Each owner $j$ observes some information of her items $\mathcal{I}^{j}$, is notified of the calibration mechanism and then asked to report a ranking $\pi^{j}$ of her items $\mathcal{I}^{j}$ [^6] If an owner does not report a ranking, a uniformly random ranking will be used. As will be shown later, this design ensures that rational owners participate in the mechanism. It is also without loss of generality to assume that owners report the full ranking of their items, and we defer the discussion to Remark 2.
  2. 2.
    After Review: Once the raw review score $\bm{y}$ is realized, the mechanism accordingly estimates and outputs the adjusted review scores as $\widehat{\bm{R}}=\widehat{\mathrm{R}}_{\mathcal{M}}(\{\pi^{j}\}_{j=1}^{m};\bm{y})$.
The key challenge of this design problem is to balance the tension between the statistical efficiency and incentive compatibility. To formalize this problem, we adopt the mechanism design approach in modeling the strategic incentives of item owners on the calibration results. That is, we assume each owner has a utility function on their items’ adjusted review scores. Let owner $j$’s utility function be $\mathrm{U}^{j}:\mathbb{R}^{n_{j}}\to\mathbb{R}$ such that each owner $j$ derives utility $\mathrm{U}^{j}\left([\widehat{R}_{i}]_{i\in\mathcal{I}^{j}}\right)$ from adjusted review scores $\widehat{\bm{R}}$, output by the calibration mechanism. We discuss how this utility function based on review scores can be translated to the one based on the final outcomes in Section 6. We employ the Bayesian game framework [15] to study the owners’ strategic decision-making at the information elicitation stage during submission, where the realization of the raw review scores $\bm{y}$ is uncertain to the owners (e.g., due to the unknown reviewer assignments and noise). We say a profile of owners’ report $\{\pi^{j}\}_{j=1}^{m}$ forms a Bayes-Nash Equilibrium (BNE)  [^7] It suffices to consider pure-strategy NEs in most part of this paper and we defer the additional notations for the more general definition of mixed-strategy NEs to Appendix 1. under mechanism $\mathcal{M}$ if for any owner $j\in[m]$, given others’ report $\pi^{-j}=\{\pi^{j^{\prime}}\}_{j^{\prime}\neq j}$, $j$’s expected utility by reporting $\pi^{j}$ is no worse than reporting any other possible ranking $\widetilde{\pi}^{j}$ or, mathematically,
$\quad\mathop{\mathbf{E}}_{\bm{y}}\left[\mathrm{U}^{j}\left(\widehat{\mathrm{R}}_{\mathcal{M}}(\pi^{j},\pi^{-j};\bm{y})\right)\right]\geq\mathop{\mathbf{E}}_{\bm{y}}\left[\mathrm{U}^{j}\left(\widehat{\mathrm{R}}_{\mathcal{M}}(\widetilde{\pi}^{j},\pi^{-j};\bm{y})\right)\right],\quad\forall j\in[m].$ (Equilibrium Condition)
We say a mechanism $\mathcal{M}$ is truthful if every owner $j$ reporting the true ranking $\pi^{j}$ of her items forms a Bayes-Nash equilibrium under $\mathcal{M}$. Through a standard revelation principle argument [29], it is without loss of generality to restrict the design space to truthful mechanisms. This equilibrium condition is sufficient to guarantee the truthfulness against unilateral deviations by each individual owner. We discuss potential extensions to address group manipulation in Section 6.
To demonstrate the theoretical guarantees of our mechanism, we make the following assumptions that are natural in peer review applications. In Section 5, we also conduct empirical studies that demonstrates the robust performance of our mechanism for settings even beyond these assumptions.
Assumption 2.1 (Informed Owners).
For each $j$, the $j$-th owner knows the relative ranking of her own items in $\mathcal{I}^{j}$, according to their ground-truth scores.
Assumption 2.2 (Exchangeable Noise Distribution).
The review noise vector $\bm{z}=(z_{1},\dots,z_{n})$ follows an exchangeable distribution in the sense that $(z_{1},\dots,z_{n})$ has the same probability distribution in $\mathbb{R}^{n}$ as in $\rho\circ\bm{z}:=(z_{\rho(1)},\dots,z_{\rho(n)})$ for any permutation $\rho$ of $1,2,\dots,n$.
Assumption 2.3 (Convex Utility).
For each $j$, the $j$-th owner’s utility function takes the form of $\mathrm{U}^{j}(\widehat{\bm{R}})=\sum_{i\in\mathcal{I}^{j}}{U}^{j}(\widehat{R}_{i})$, where ${U}^{j}:\mathbb{R}\to\mathbb{R}$ is a non-decreasing convex function.
Assumption 2.1 is standard in mechanism design and reflects each owner’s private knowledge about their items without requiring any knowledge about other items. [^8] For instance, in basic models of auction design, bidders are assumed to perfectly know their value about the item [29]. Our assumption of owners knowing the ranking information is also inspired by many machine learning applications that learn parameters from humans’ ranking/comparison data which are posited to be more accurate than their knowledge of absolute values [44, 3]. Assumption 2.2 imposes symmetry on review noises, and Assumption 2.3 captures a utility structure that naturally arises in “high-risk-high-reward” tournament settings (Section 6 discusses more about the root of this assumption, its adoption in similar economic problems, and empirical evidences from ICLR data).
3 An Isotonic Mechanism for Completely Overlapping Ownership
We begin with a rank-calibrated score estimator $\widehat{\mathrm{R}}(\pi;\bm{y})$ that employs isotonic regression on raw review scores $\bm{y}$ and any reported ranking $\pi$ from the owner, via the following convex program,
$\widehat{\mathrm{R}}(\pi;\bm{y})=\mathop{argmin}_{\bm{r}\in\mathbb{R}^{n}}$ $\|\bm{y}-\bm{r}\|^{2}$
s.t. $r_{\pi(1)}\geq r_{\pi(2)}\geq\dots\geq r_{\pi(n)},$
[…]
In the general setup with overlapping ownership, we would like to design a mechanism that aggregates information from as many owners as possible, while preserving the desirable properties of truthfulness and statistical efficiency. The most natural design is perhaps to use the (weighted) averaged scores from the estimates based on each owner’s reported ranking. The mechanism takes the input of a problem instance, specified by review scores $\bm{y}\in\mathbb{R}^{n}$ and owner credentials $\{\alpha^{j}\}_{j=1}^{m}\in[0,1]^{m}$, and outputs the adjusted review score $\widehat{\bm{R}}$ using rankings elicited from all owners. Notably, the owner credential is a set of weights $\{\alpha^{j}\}_{j=1}^{m}$ that pre-specifies the different levels of influence of the owners; we view them as part of the problem instance that reflects each owner’s expert level, reputation scores and track records in the given instance. In practice, one could simply set $\alpha^{j}=1$ to evenly weigh on each owner’s reported ranking, or set personalized $\alpha^{j}$ for each owner $j$ to account for different owner’s expert level — we include it as a part of the input so that it is a choice to the practitioner’s discretion. We formally describe these procedures in Mechanism 1. The rest of this section is to analyze the situations when it does or does not work.
Input: Review scores $\bm{y}\in\mathbb{R}^{n}$, owner credentials $\{\alpha^{j}\}_{j=1}^{m}$.
for every $j\in[m]=\{1,2,\dots,m\}$  do
       Elicit ranking $\pi^{j}$ from owner $j$.
       Solve for ranking-calibrated scores $\widehat{\bm{R}}^{j}\leftarrow\widehat{\mathrm{R}}(\pi^{j};\bm{y}).$
Set $\widehat{\bm{R}}\leftarrow\sum_{j=1}^{m}\alpha^{j}\widehat{\bm{R}}^{j}/\sum_{j=1}^{m}\alpha^{j}$.
return $\widehat{\bm{R}}$.
Mechanism 1 Isotonic Mechanism under Completely Overlapping Ownership
3.1 Truthfulness under Completely Overlapping Ownership
We start by considering a useful special case in which every submission is owned by every owner, referred to the completely overlapping ownership situation. As it turns out, truth-telling forms a Nash equilibrium in this setting under Mechanism 1999This result can be related to the ex-post incentive compatibility of Mechanism 1, a common goal of truthful mechanism design. Moreover, this equilibrium is payoff dominant [16], i.e., one that is Pareto superior to all other Nash equilibria in the game. Put simply, all owners would prefer this equilibrium, because it simultaneously gives every owner the highest equilibrium utility among all possible equilibrium outcomes. This nice property makes it more plausible to expect agents’ truthful behaviors, despite the potential existence of multiple equilibria in the game; more discussions about the truthfulness of this mechanism from the perspective of behavioral theory can be found in Remark 1.
Theorem 1.
Under Assumptions 2.1, 2.2, and 2.3, if the ownership is completely overlapping,
  1. 1.
    It forms a Bayes-Nash equilibrium for each owner to truthfully report the ranking in Mechanism 1.
  2. 2.
    This equilibrium of truthful report is payoff dominant.
We defer the proof of Theorem 1 to Appendix A and conclude this subsection with additional evidence on the truthfulness of the Isotonic Mechanism according to the behavioral game theory.
Remark 1 (Truthfulness of Isotonic Mechanism from a behavioral angle).
Mechanism 1 induces the strategic game with special structures known as the common interest games [16, 2]. We briefly highlight the connections here and discuss evidence from both empirical human preferences and behavioral theory on why truthful behavior should be expected in such games. In its simplest form, the payoff matrix of a common interest game typically takes one of the three structures illustrated in Table 1 (only orders of the payoff values matter). Namely, there is an action profile (the upper left cell of each game) that represents the common interest of both players; this corresponds to the truthful action profile $(\pi^{\star},\pi^{\star})$ under the Isotonic Mechanism which forms a payoff-dominant Nash equilibrium, according to Theorem 1. Meanwhile, the three variants of common interest games capture the different possible orders of payoff values in the lower right cell and those in the off-diagonal cells. Any variant of these payoff matrices may be realized under the Isotonic Mechanism, since our proof using Jensen’s inequality only offers upper bounds on the utilities of these non-truthful action profiles (see e.g., Equation (5)).
$\pi^{\star}$ $\pi^{\circ}$
$\pi^{\star}$ $10,11$ $4,5$
$\pi^{\circ}$ $4,5$ $3,4$
$\pi^{\star}$ $\pi^{\circ}$
$\pi^{\star}$ $10,11$ $1,1$
$\pi^{\circ}$ $1,1$ $3,4$
$\pi^{\star}$ $\pi^{\circ}$
$\pi^{\star}$ $10,11$ $1,5$
$\pi^{\circ}$ $5,1$ $3,4$
Table 1: An illustration of three typical common interest game payoff matrices.
It turns out that in experimental games and also in real life, people tend to always choose the upper left cell — the payoff dominant outcomes — in all the three possible game structures above [2]. This behavior is expected under the first payoff matrix since $(\pi^{\star},\pi^{\star})$ is the unique Nash equilibrium there. It is somewhat reasonable under the second payoff matrix, as argued by Harsanyi et al. [16] for its “self-reinforcing property” — everyone thinks the other has no reason to prefer the payoff-dominated equilibrium outcomes. However, realizing the payoff dominant outcome becomes less clear under the third payoff matrix, because $(\pi^{\circ},\pi^{\circ})$ forms a “risk-dominant” Nash equilibrium — i.e., compared to $\pi^{\star}$, $\pi^{\circ}$ is the less risky action under uncertainty of the opponent’s choice [16] (specifically, under $(\pi^{\circ},\pi^{\circ})$ equilibrium, row player gets $5$ instead of $3$ if the opponent deviates from the equilibrium action $\pi^{\circ}$ to $\pi^{\star}$). Interestingly, standard game theory cannot account for this phenomenon of coordination on the payoff dominant outcome under the second and third payoff matrices, i.e., the Hi-Lo Paradox. One major theory that explains this phenomenon is the team reasoning by Bacharach [2] as a non-game-theoretic rationale, i.e., “What do we want? And what should I do to play my part in achieving this?”. If there is common knowledge that both players adopt the team-reasoning mode of choosing their strategies, then both would choose the payoff dominant equilibrium. These also serve as behavioural evidences of agents’ truthfulness in the Isotonic Mechanism.
3.2 Non-truthfulness Beyond Completely Overlapping Ownership
Despite the nice properties of the Nash equilibrium shown in Theorem 1, its assumption of completely overlapping ownership is usually not satisfied in many application scenarios — different papers are often written by different sets of authors. To extend Mechanism 1 to the case of incompletely overlapping ownership, one natural choice is to average the scores according to the ownership, $\widehat{R}_{i}\leftarrow{\sum_{j=1}^{m}e^{j}_{i}\widehat{R}^{j}_{i}}/{\sum_{j^{\prime}=1}^{m}e^{j^{\prime}}_{i}}$, where the binary ownership indicator $e^{j}_{i}=1$ if and only if owner $j$ owns item $i$. The question is whether truthful reporting still forms a Nash equilibrium. Unfortunately, the answer turns out to be “No”. Specifically, the owners may gain from the misreporting strategy and one possible manipulation is to use the reputation of good items to promote a bad item — it may only hurt the good item a little, but help the bad item a lot. We demonstrate such kind of strategic manipulations via a concrete example below.
Example 1 (Non-truthfulness under Partially Overlapping Ownership).
Consider a case of $m$ owners and $n=3$ items. The items have ground-truth scores $R_{1}=9,R_{2}=8,R_{3}=4$ — one weak and two strong. The ownership is not completely overlapping. As illustrated in Table 2, the first $m-1$ owners work together on the two strong items $1,2$, while the $m$-th owner works on the item $2,3$. For each $j$, let the $j$-th owner’s utility be $\mathrm{U}^{j}(\widehat{\bm{R}})=\sum_{i\in\mathcal{I}^{j}}\mathop{max}\{R_{i}-5,0\}$.
Item Owner $1$ $2$cdots $m-1$ $m$
1 1 1 $\cdots$ 1 0
2 1 1 $\cdots$ 1 1
3 0 0 $\cdots$ 0 1
True Scores
$R_{1}=9$
$R_{2}=8$
$R_{3}=4$
Table 2: An illustration of the ownership matrix $E=(e_{i}^{j})_{m\times n}$ and ground-truth scores in Example 1.
For simplicity, suppose the reviews are noiseless so that, if all owners report their ranking truthfully, the calibration is perfect, i.e., $\widehat{\bm{R}}=\textbf{y}=\bm{R}$. However, we can observe that the $m$-th owner does not have the incentive to report the ranking of items $1,2$ truthfully, given that the first $m-1$ owners report the true ranking. That is, under the flipped ranking that $\widetilde{R}_{2}^{m}\leq\widetilde{R}_{3}^{m}$, Mechanism 1 would have $\widetilde{R}_{2}^{m}=\widetilde{R}_{3}^{m}=6$ and the adjusted scores $\widetilde{R}_{1}=9,\widetilde{R}_{2}=\frac{1}{m}(\widetilde{R}_{2}^{m}+\sum_{j=1}^{m-1}R_{2}^{j})=8-\frac{2}{m},\widetilde{R}_{3}=\widetilde{R}_{3}^{m}=6$. We can see that the utility of $m$-th owner under truthful ranking is, $13<3-\frac{2}{m}+$ 1, strictly worse than that under the non-truthful ranking for any $m\geq 3$.
At a high level, such strategic behavior is due to the owner’s uneven influence on the average scores of its different items — we will formalize this intuition in Theorem 2. One may speculate that we can potentially resolve this kind of situation through a careful reweighing process of each owner’s influence. The answer also turns out to be “No”. Specifically, in the example above, $m$-th owner, as the sole owner of item 3, would always fully determine the score of item 3; yet, this owner can never fully determine the score of item 2 unless we choose to ignore other owners’ opinion or assign some weight to the raw review scores, which more or less defeats the purpose of this community effort of calibrating review scores. Hence, we need to seek a different approach to help us truthfully aggregate the owners’ information in practice.
4 Restoring Truthfulness via Partitioning
To ensure the truthfulness beyond complete ownership, this section studies a partition-based approach. Our main idea is to partition the ownership sets into multiple blocks, ensuring complete overlap within each block by some owners. We then individually apply Mechanism 1 to each block, eliciting truthful rankings only from the owners who completely own that block, as illustrated in Figure 3. Finally, we use the elicited rankings from these blocks to estimate the ground-truth scores. We formally described this procedure in the following Mechanism 2.
Figure 3: A partition of partially overlapping ownership.
Parameters : Item set partition $\mathfrak{S}=\{{\mathcal{S}}_{1},\ldots,{\mathcal{S}}_{K}\}$ based on ownership relations $\{\mathcal{I}^{j}\}_{j=1}^{m}$.
Input: Review score $\bm{y}\in\mathbb{R}^{n}$, owner credentials $\{\alpha^{j}\}_{j=1}^{m}$.
for every ${\mathcal{S}}_{k}\in\mathfrak{S}$ do
       Find all owners with complete ownership of ${\mathcal{S}}_{k}$, i.e., ${\mathcal{T}}_{k}\leftarrow\{j\in[m]:{\mathcal{S}}_{k}\subseteq\mathcal{I}^{j}\}$.
       if $|{\mathcal{T}}_{k}|=0$ or $|{\mathcal{S}}_{k}|=1$  then
            Do not elicit any ranking information.
      else
             Apply Mechanism 1 to item set ${\mathcal{S}}_{k}$ with owner set ${\mathcal{T}}_{k}$ and weights $\{\alpha^{j}\}_{j\in{\mathcal{T}}_{k}}$ to estimate the ground-truth score of items in ${\mathcal{S}}_{k}$, denoted by $\widehat{\bm{R}}[{\mathcal{S}}_{k}]$.
return $\widehat{\bm{R}}=\{\widehat{\bm{R}}[{\mathcal{S}}_{k}]\}_{k=1}^{K}$.
Mechanism 2 Partition-based Isotonic Mechanism for Partially Overlapping Ownership
Compared to Mechanism 1, Mechanism 2 additionally determines a set of parameters which include a partition of item set $\mathfrak{S}=\{{\mathcal{S}}_{1},\ldots,{\mathcal{S}}_{K}\}$ that satisfies $\bigcup_{k=1}^{K}{\mathcal{S}}_{k}=[n],{\mathcal{S}}_{k}\cap{\mathcal{S}}_{k^{\prime}}=\varnothing$. We ask the parameters to only depend on the ownership relations for practical considerations: since the mechanism may be truthful only under some proper choice of parameters, the owners should be informed of the parameters during submission, when the ownership relations are formed but the review scores are not realized. One can observe that in Mechanism 2, each owner $j\in{\mathcal{T}}_{k}$ has the complete ownership of the items in ${\mathcal{S}}_{k}$ by construction. As a corollary of Theorem 1, it is easy to show that Mechanism 2 is truthful under any partition $\mathfrak{S}=\{{\mathcal{S}}_{1},\ldots,{\mathcal{S}}_{K}\}$ and any weights $\{\{\beta^{j}_{k}\}_{k=1}^{K}\}_{j=1}^{m}$, because we only elicit rankings of items within each block from those who completely own the block. Thus, all owners $j\in{\mathcal{T}}_{k}$ will report truthful ranking over items in ${\mathcal{S}}_{k}\subseteq\mathcal{I}^{j}$. Owing to the independence of score estimates for items between different partition blocks, the overall mechanism is truthful; we formalize it in the following corollary.
Corollary 1.
Mechanism 2 is truthful for any input instance and parameter choice in the following sense: it forms a Bayes-Nash equilibrium for each owner to truthfully report the ranking of their items within each block specified by the partition.
Moreover, Mechanism 2 can be implemented in practice to promote both simplicity and fairness for participants. Below, we highlight several key aspects and advocate for field experiments to better evaluate and optimize for its practical applicability.
Remark 2 (Practical Implementation and Fairness Considerations).
While Mechanism 2 only elicits the ranking of items within each block ${\mathcal{S}}_{k}$, the mechanism designer may implement the mechanism by simply asking owners to rank all their items with the commitment to only use part of the ranking of each owner as specified by the mechanism. Under such a committed mechanism, any owner will be indifferent about the order of any two items in any two different partition sets ${\mathcal{S}}_{k}$ and ${\mathcal{S}}_{k^{\prime}}$ since their order will never be used by the mechanism. Thus, the owner can be assumed to break ties in favor of the designer and reveal their full ranking truthfully. [^10] There is also a formal justification for such tie-breaking by adding a negligible amount of randomness for picking an arbitrarily different partition, which consequently creates a negligible amount of incentive for any owner to report other rankings truthfully. Such equilibrium refinement via randomization has appeared in the literature on mechanism design [29]. Notably, however, the designer’s commitment to not using any order information beyond what the mechanism specified (despite such information being elicited) is important, since otherwise, truthfulness will not hold. In reality, the mechanism designer as a trusted authority (e.g., the organizers of a large ML conference) usually has such commitment power.
When applying our mechanism to conference paper peer review, a potential limitation one might argue is that if an author only has one or two papers, then the mechanism may not elicit ranking information from this author, which may be “unfair” to this author’s papers. We remark that this is not a major concern because most papers in today’s ML conferences have multiple authors, so long as some author of a paper offered ranking information, that paper’s review score will be calibrated. For instance, in ICLR 2023, approximately 76.5% of papers had at least one “senior” author with multiple submissions and all these papers’ review scores could be calibrated based on these “senior” authors’ reported ranking. In addition, there are other ways to further mitigate potential unfairness. For example, while partitioning may lead to some owners’ information being unused across blocks, we could randomize the choice of partitions in the mechanism to ensure that each owner’s reported ranking is incorporated in the learning process with non-zero probability. That said, while our mechanism could mitigate the issue of fairness to some extent, we acknowledge that it does not come with a guarantee of perfect fairness which in general is an extremely deep concept, both socially and philosophically, and is also multifaceted. We expect follow-up research to more thoroughly address this fairness issue. For instance, one possible direction is to identify Pareto-optimal designs that reduce overall review noise while maintaining a reasonable degree of fairness.
4.1 The Necessity of Partition-based Isotonic Mechanisms
{tblr}
colspec = ccccc, cell22 = red!30, cell32 = red!30, cell43 = yellow!30, cell44 = yellow!30, Item Owner & $1$ $2$ $3$
$1$ 1 0 1
$2$ 1 1 0
$3$ 0 1 1
  {tblr}colspec = ccccc, cell22 = red!30, cell32 = red!30, cell33 = yellow!30, cell43 = yellow!30, cell24 = cyan!20, cell44 = cyan!20, Item Owner & $1$ $2$ $3$
$1$ 1 0 1
$2$ 1 1 0
$3$ 0 1 1
Table 3: An illustration of partition-based mechanism (left) and a more general mechanism (right). In the partition-based mechanism, each color denotes a partitioned item set block. In the more general mechanism, different owners can have different item set partitions, as marked by different colors.
At this point, one may wonder whether restricting ranking elicitation to partitioned item sets is necessary, as it inevitably forces the mechanism to ignore some of the owners’ revealed information across the partition blocks. For example, consider an instance with $m=n=3,\mathcal{I}^{1}=\{1,2\},\mathcal{I}^{2}=\{2,3\},\mathcal{I}^{3}=\{1,3\}$ — i.e., each owner owns two items, and each item has two co-owners (see the illustration in Table 3). We can see that, with any partition of the item set, Mechanism 2 can elicit ranking information from at most one owner. However, a more aggressive design could be eliciting the ranking information from every owner $j$ about her items in $\mathcal{I}^{j}$ and then applying a generalized version of Mechanism 1 by using the elicited (partial) ranking of $j$’s items. This is certainly a statistically more efficient design, but the key question is whether a mechanism of such kind can still be guaranteed to be truthful. Our study next shows that the answer is unfortunately “No”. We show that, within a much broader class of isotonic-regression-based mechanisms, the partition-based mechanism as prescribed in Mechanism 2 is essentially the only candidate that can guarantee truthful owner behaviors — that is, one may have to give up eliciting the comparison information between the partitioned blocks in order to trade for incentive properties. This reveals an intrinsic tradeoff between statistical efficiency and incentive guarantee within the general class of Isotonic Mechanisms. We leave it as an intriguing open direction to explore alternative estimation methods other than isotonic regression that can balance statistical efficiency and incentive compatibility — no such method is known so far even for the single-owner case.
We start by generalizing Mechanism 2 to a broader class of Isotonic Mechanisms, as described in Mechanism 3. In this more generalized Isotonic Mechanism, the calibrated score of any item $i\in[n]$ is now allowed to depend on the information elicited from all its owners, denoted by $\mathcal{J}^{i}\subseteq[m]$. Concretely, the input to Mechanism 2 and Mechanism 3 is the same, but the parameter choices in Mechanism 3 are strictly more general: (1) the item partition $\mathfrak{S}^{j}=\{{\mathcal{S}}^{j}_{1},{\mathcal{S}}^{j}_{2},\dots\mathcal{S}^{j}_{K^{j}}\}$, satisfying $\bigcup_{k=1}^{K^{j}}{\mathcal{S}}_{k}^{j}=\mathcal{I}^{j},{\mathcal{S}}_{k}^{j}\cap{\mathcal{S}}_{k^{\prime}}^{j}=\varnothing$, is now allowed to be different across owners; (2) additionally, a weight vector $\bm{\beta}^{j}=[\beta_{1}^{j},\beta_{2}^{j},\dots,\beta_{n}^{j}]\in\mathbb{R}_{\geq 0}^{n}$ is introduced to allow owner $j$’s more fine-grained influence on her items with itemized weights. This is much more powerful than $j$’s influence in Mechanism 2 that is restricted to be the same among items. The parameters are also determined from the input problem instance, specifically the ownership relation $\{\mathcal{I}^{j}\}_{j=1}^{m}$. The adjusted score for each item is then similarly determined by a weighted linear combination of the rank-calibrated scores from each owner. Notably, parameter $\beta_{i}^{j}$ for any $i\not\in\mathcal{I}^{j}$ is never used by Mechanism 3 hence can be arbitrary. We nevertheless defined $\bm{\beta}^{j}$ as a vector in $\mathbb{R}_{\geq 0}^{n}$ mainly for notational convenience.
Parameters : Personalized item partition $\mathfrak{S}^{j}=\{{\mathcal{S}}_{1}^{j},\ldots,{\mathcal{S}}_{K^{j}}^{j}\}$ and itemized weights $\bm{\beta}^{j}$ for each owner $j\in[m]$ based on ownership relations $\{\mathcal{I}^{j}\}_{j=1}^{m}$.
Input: Review scores $\bm{y}\in\mathbb{R}^{n}$, owner credentials $\{\alpha^{j}\}_{j=1}^{m}$.
for every $j\in[m]$ do
       Apply Mechanism 1 to owner $j$ with partition $\mathfrak{S}^{j}$ to estimate the ground-truth score, denoted by $\widehat{R}^{j}_{i}$, for each item $i$ in $\mathcal{I}^{j}$.
return $\big{\{}\widehat{R}_{i}=\sum_{j\in\mathcal{J}^{i}}\alpha^{j}\beta^{j}_{i}\widehat{R}^{j}_{i}/\sum_{{j^{\prime}}\in\mathcal{J}^{i}}\alpha^{j^{\prime}}\beta^{j^{\prime}}_{i}\big{\}}_{i=1}^{n}$.
Mechanism 3 Isotonic Mechanism with Personalized Partition
It is easy to see that the class of Mechanism 3 strictly contains the class of Mechanism 2. On the one hand, it can be immediately verified that Mechanism 2 with any item set partition $\mathfrak{S}=\{{\mathcal{S}}_{1},\ldots,{\mathcal{S}}_{K}\}$ can be formulated as Mechanism 3 with some parameter $\big{\{}\mathfrak{S}^{j},\bm{\beta}^{j}\big{\}}_{j=1}^{m}$ (see the proof of Theorem 2 for the details) On the other hand, Mechanism 3 does have strictly more expressivity than Mechanism 2. For example, the non-partition-based mechanism illustrated above in Table 3 can be captured by the following elicitation rule: $\mathfrak{S}^{1}=\{\{1,2\}\},\bm{\beta}^{1}=[1,1,0],\mathfrak{S}^{2}=\{\{2,3\}\},\bm{\beta}^{2}=[0,1,1]$ and so on. One may now wonder whether this strictly more general Mechanism 3 is also strictly more powerful. Our next result shows that the answer is “NO” if the mechanism is truthful.
Theorem 2 (The Necessity of Global Partition).
Under Assumptions 2.1, 2.2, and 2.3, for any $\mathcal{M}$ in the format of Mechanism 3 that is truthful for every input, there exists a $\mathcal{M}^{\prime}$ in the format of Mechanism 2 that elicits no less ranking information. Formally, if the order of any two items $i,i^{\prime}$ is truthfully elicited by $\mathcal{M}$, then their order is also truthfully elicited by $\mathcal{M}^{\prime}$.
The formal proof is somewhat involved hence deferred to Appendix D. At a high level, our proof is structured with two main steps. First, we characterize the necessary conditions under which Mechanism 3 is truthful, which is based on the following key Lemma 1. Second, we show that for any Mechanism 3 under the necessary condition, a Mechanism 2 can be constructed to elicit at least as much ranking information.
Lemma 1.
Under Assumptions 2.1, 2.2, and 2.3, if a Mechanism 3 with parameters $\big{\{}\mathfrak{S}^{j},\bm{\beta}^{j}\big{\}}_{j=1}^{m}$ is truthful, then the following two conditions must both hold:
  1. (I)
    Each owner has balanced influence on the items within each partition blocks for any input, i.e.,
    for $\text{for any }j\in[m],{\mathcal{S}}\in\mathfrak{S}^{j},i,i^{\prime}\in{\mathcal{S}},\qquad\text{we have }\,\omega^{j}_{i}=\omega^{j}_{i^{\prime}},$
    where $\omega_{i}^{j}=\alpha^{j}\beta^{j}_{i}/\sum_{j^{\prime}\in\mathcal{J}^{i}}\alpha^{j^{\prime}}\beta^{j^{\prime}}_{i}$ denotes owner $j$’s relative influence on the score of item $i$.
  2. (II)
    The parameters $\big{\{}\mathfrak{S}^{j},\bm{\beta}^{j}\big{\}}_{j=1}^{m}$ has a valid partition structure in the following sense,
    for $\text{for any }j,j^{\prime}\in[m],{\mathcal{S}}\in\mathfrak{S}^{j},{\mathcal{S}}^{\prime}\in\mathfrak{S}^{j^{\prime}},\text{ if }{\mathcal{S}}\cap{\mathcal{S}}^{\prime}\neq\varnothing,\text{ then }\beta^{j}_{i}=\beta^{j}_{i^{\prime}},\text{ for all }i,i^{\prime}\in{\mathcal{S}}\cup{\mathcal{S}}^{\prime}.$ (1)
Moreover, these two conditions are equivalent to each other.
4.2 Partition Optimization and Its Hardness
As illustrated above, any partition will make Mechanism 2 truthful and that partition-based Isotonic Mechanisms are the only truthful mechanisms within a much broader class of Isotonic Mechanisms. These together point to the last key piece of our design — identifying the item partition for the optimal statistical efficiency. In Remark 3, we make two key observations to formalize the relationship between partition structure and statistical efficiency.
Remark 3 (On Partition Optimization for Proximal Statistical Efficiency).
It is intuitive to see that both the number of owners and items of a partition block would influence the statistical performance of the review score calibration. We formalize these observations into two monotonicity laws about partition optimization. First, the coarser the partition is, the better the estimation accuracy. The proof is based on the observation that projecting onto a strictly smaller feasible region that contains the ground-truth leads to a smaller mean squared error in the estimated scores [37, Proposition 3.4]. Second, the more owners each partition block has, the better the estimation accuracy. This is straightforward when owners have independent perception noise about their item scores. That is, let $\widehat{\bm{R}}^{j}\sim\bm{R}+\bm{z}$ be the score estimate based on owner $j$’s reported ranking, where $\bm{R}$ is the true score vector and $\bm{z}$ consists of i.i.d. noise with bias $b$ and variance $\sigma^{2}$. Then, for any item $i$, its score estimate aggregated across $n$ owners, $\widehat{R}_{i}=\frac{1}{n}\sum_{i=1}^{n}\widehat{R}_{i}$, has mean squared error $b^{2}+\frac{\sigma^{2}}{n}$, which clearly decreases as the number of owners $n$ increases.
However, these monotonicity laws alone are not sufficient to solve the partition optimization problem. The set of feasible partitions is only partially ordered, and the rate at which noise is reduced remains difficult to quantify. As a result, precisely characterizing the relationship between partition structure and statistical efficiency remains an open question, if not an intractable one. Proving the optimality of statistical estimators is notoriously challenging, as exemplified by Stein’s paradox, where even the seemingly natural sample mean turns out to be suboptimal under certain conditions.
Following from Remark 3, we resort to the design of proxy objectives for partition optimization under two general principles: (1) Informativeness — larger blocks with more ranking information; [^12] For instance, the ranking of a size-$10$ item block has $\frac{10\times 9}{2}=45$ pairs of comparison, whereas the rankings of two size-$5$ item blocks have $202\times\frac{5\times 4}{2}=$ 20 pairs of comparison. (2) Robustness — each block has more owners. However, there is an inherent tension between the two principles, as larger blocks usually have fewer joint owners since it is less usual for multiple owners to jointly own many items. Thus the pursuit of larger blocks will prevent us from eliciting rankings from multiple owners. To address this tradeoff, we formalize two distinct metrics that quantify the informativeness and robustness of a partition and study the computational complexity of partition optimization for proximal statistical efficiency.
Informativeness (as optimization objective). In regard to the performance, we aim to identify partitions that are most informative for the review score calibration. On one hand, larger-sized blocks generally contain more ranking information, whereas in the extreme case if a block contains only one item, the mechanism cannot elicit any ranking information within the block. On the other hand, the selection of a large block may sometimes render other blocks small, leading to complex tradeoff about partition choices. To systemically evaluate the coverage quality of a partition $\mathfrak{S}$, we introduce an informativeness objective function,
$\operatorname{obj}(\mathfrak{S})=\sum_{k=1}^{K}w(|{\mathcal{S}}_{k}|),$
determined by some wellness function $w$ on the size of each partition block. For examples, consider the following two choices of the wellness function $w(\cdot)$. One is $w(x)=x^{2}$ as a simple quadratic function of block size, which relates to the number of pairwise comparisons within a block:
[…]
Another choice is $w(x)=\mathop{max}\{x-1,0\}$, which is called size-focused objective, defined below:
[…]
where the second equation is because any block in the partition has at least $1$ item (i.e., $|{\mathcal{S}}_{k}|\geq 1$). Since $n$ is a constant, maximizing the size-focused objective is equivalent to minimizing the size of the partition $K$ (thus the name “size-focused”).
Generally, it is reasonable to expect that function $w(\cdot)$ should be monotonically increasing and convex — to see convexity, if a block of length $l$ is broken into two blocks with length $l_{1},l_{2}$, then the original longer block must be preferred, i.e., $w(l)\geq w(l_{1})+w(l_{2})$ for any $l=l_{1}+l_{2}$, which is precisely the definition of convexity. Another natural requirement is $w(0)=0$ in order to prevent empty blocks to contribute any value to the partition wellness. Both functions above satisfy these properties, and so does any monomial $w(x)=x^{p}$ for $p\geq 1$. Beyond these properties, however, it appears impossible to exactly know the format of $w(\cdot)$. To overcome this challenge, we resort to robust algorithms in subsequent subsection 4.3 and look to design algorithms that can simultaneously perform well for every wellness function satisfying these properties.
Robustness (as optimization constraint). If the informativeness metric above is regarded as the objective we want the partition to maximize, then the robustness metric can be viewed as a constraint we wish the partition to satisfy. In regard to robustness, aggregating information from multiple owners can help mitigate each individual owner’s noisy perception of their items in situations where Assumption 2.1 does not hold. Thus, the minimal number of owners for each partition block matters. Formally, we say a partition $\mathfrak{S}=\{{\mathcal{S}}_{1},\ldots,{\mathcal{S}}_{K}\}$ is $L$-strong if there are at least $L$ owners who share all items in the same partition block, i.e., $\left|{\mathcal{T}}_{k}\right|\geq L,\forall{\mathcal{S}}_{k}\in\mathfrak{S}$ with $|{\mathcal{S}}_{k}|>1$. The larger $L$ is, the less prone Isotonic Mechanism is to the noise in each owner’s perceived ranking. Notably, $L\geq 1$ is the bare minimum requirement for Mechanism 2 to work. Next we show that generating a feasible $L$-strong partition reduces to generating a feasible $1$-strong partition for a different ownership instance. This reduction helps us convert any $L$-strongness constraint to a certain $1$-strongness constraint during our design of the partition optimization algorithm.
Proposition 1 (Reduction to $1$-Strong Partition).
For any ownership instance $\mathcal{O}=\{\mathcal{I}^{j}\}_{j\in[m]}$, we can construct in $O\left(m^{L}n\right)$ time a different ownership instance $\mathcal{O}^{\prime}=\{\overline{\mathcal{I}}^{j}\}_{j\in[m]}$ with the same owner and item set such that any partition ${\mathcal{S}}$ is $L$-strong in $\mathcal{O}$ if and only if ${\mathcal{S}}$ is $1$-strong in $\mathcal{O}^{\prime}$.
Proposition 1 shows that to optimize the informativeness objective, it is without loss of generality to design algorithms to maximize the objective under $1$-strongness, for a transformed ownership set instance. We defer its formal proof to Appendix B but provide a simple example below to illustrate the main idea of the reduction: any $2$-strong partition in the ownership instance $\mathcal{O}=\{\{1,2\},\{1,2,3\},\{2,3\}\}$ is a $1$-strong partition in the ownership instance $\mathcal{O}^{\prime}=\{\{1,2\},\{2\},\{2,3\}\}$, and vice versa. Essentially, the instance $\mathcal{O}^{\prime}$ is constructed by using one owner to represent each subset of $L$ owners in the original instance $\mathcal{O}$, whose common items are owned by this owner. This is a polynomial-time reduction for small constant $L$, which is often the case in academic conference review due to the small number of co-authors on a paper. Notably, this reduction does not fundamentally simplify the problem due to the intrinsic hardness of maximizing the informativeness objective, as shown in the following proposition; see Appendix C for its formal proof.
Proposition 2 (Hardness of Partition Optimization).
It is NP-hard to find the optimal partition $\mathfrak{S}=\{{\mathcal{S}}_{1},\ldots,{\mathcal{S}}_{K}\}$ that maximizes the size-focused objective in Eq. (3) subject to $1$-strongness.
4.3 Fast Greedy Partition with Robust Approximation Guarantees
Input: Ownership sets $\{\mathcal{I}^{j}\}_{j=1}^{m}$.
Initialize the partition as $\mathfrak{S}=\{\}$ and set of selected items $\overline{\mathcal{I}}=\varnothing$.
while  ¯$\overline{\mathcal{I}}\subset[n]$ do
       Determine the largest residual item set ${\mathcal{S}}^{*}=\mathcal{I}^{j^{*}}\setminus\overline{\mathcal{I}}$ where $j^{*}=\mathop{argmax}_{j\in[m]}|\mathcal{I}^{j}\setminus\overline{\mathcal{I}}|$.
       Update ¯¯$\overline{\mathcal{I}}\leftarrow\overline{\mathcal{I}}\cup{\mathcal{S}}^{*}$ and $\mathfrak{S}\leftarrow\mathfrak{S}\cup\{{\mathcal{S}}^{*}\}$.
return $\mathfrak{S}$.
Algorithm 4 A Greedy Algorithm for $1$-Strong Partition
Despite the computational hardness for the optimal partition, we show that a natural greedy algorithm presented in Algorithm 4 can efficiently find $1$-strong partitions with a provably good approximation guarantee. In words, this greedy algorithm just iteratively selects the largest residual item set owned by some owner. This algorithm can be implemented in time almost linear in the input size $\sum_{j\in[m]}{|\mathcal{I}^{j}|}$, up to a $\log(m)$ factor, thus is essentially the fastest algorithm one could hope to design.
The highlight of this greedy algorithm is that, though the algorithm itself is fully agnostic to any partition objectives, its output is a robustly high-quality partition for every natural wellness function, i.e., monotone and convex wellness functions, as formalized in the following Theorem 3. In particular, consider the following hypothesis class $\mathcal{W}$ of wellness functions
[…]
Algorithm 4 enjoys the following strong approximation guarantees for every $w$ within $\mathcal{W}$, whose proof is deferred to Appendix E. In Remark 4, we give the exact approximation ratio for some optimization objectives of special interests and show its tightness. It is also worth mentioning that while Algorithm 4 only finds a partition satisfying $1$-strongness, the algorithm as well as its constant approximation guarantees can be readily extended to find the optimal $L$-strong partition in $O\left(m^{L}n\right)$ time, using the reduction in Proposition 1.
Theorem 3 (Robust Approximation of Greedy).
For any ownership instance $\{\mathcal{I}^{j}\}_{j\in[m]}$ with $N=\sum_{j\in[m]}{|\mathcal{I}^{j}|}$ total number of owner-item pairs, Algorithm 4 runs in $O(N\log m)$ time and outputs a 1-strong partition that is simultaneously a $c(w)$-approximation of the optimal objective for every $w\in\mathcal{W}$, where
$c(w)=\inf\Big{\{}\frac{w(x)}{w^{\prime}_{-}(x)x}\Big{|}w^{\prime}_{-}(x)>0,x\geq 2\Big{\}}.$
Above, $w^{\prime}_{-}(a)=\lim_{x\to a^{-}}\frac{w(x)-w(a)}{x-a}$ denotes the left derivative of $w$ at $a$.
Remark 4.
Theorem 3 readily implies that the partition found by the greedy algorithm simultaneously approximates at least $1/2$ of both the optimal comparison-focused objective and size-focused objective. More generally, for $\alpha$-th degree polynomial functions $w(x)=x^{\alpha}$, the greedy algorithm simultaneously guarantees $(1/\alpha)$-fraction of the optimal objective since $w(x)/[w^{\prime}(a)x]=1/\alpha$ for every $x$. Such objective-agnostic property of Algorithm 4 is especially desirable in our application scenarios where the concrete form of the partition objective is difficult (if not impossible) to know. Moreover, the approximation ratio in Theorem 3 for the greedy algorithm is provably tight under every monomial objective $f(\mathfrak{S})=\sum_{k=1}^{K}|{\mathcal{S}}_{k}|^{\alpha}$, $\alpha>1$, and the construction of lower bound instances can be found in Appendix E.1.
5 Empirical Evaluation in the Context of ML Conference Review
Figure 4: An illustration of applying our proposed Isotonic Mechanism to the conference reviewing procedure.
In this section, we study the empirical performance of our proposed mechanism by applying it to our primary motivating domain of academic conference peer review. An illustrative procedure of applying our mechanism to conference peer review can be found in Figure 4. One major challenge of evaluation is that its performance measures rely on the underlying ground-truth scores, and such information is unattainable in most applications — indeed, if we already know the ground-truth scores, peer review would not be needed anymore. Fortunately, from both ICLR and ICML, we are able to collect some parts of their data and synthesize some of the unobservable parts. Below, we describe our experiment setups and results.
5.1 Experiment Setups: Datasets, Baselines and Metrics
We start by elaborating on the preprocessing procedures and characteristics of the two real-world datasets we experimented on.
  1. 1.
    ICLR 2021-2023: We collect ownership relations and review scores from ICLR 2021-2023 [17, 18, 19] that are made publicly available from OpenReview.net. Based on the recorded review scores, we simulate the ground-truth scores $\bm{R}=\textbf{y}-\bm{z}$, where $\bm{z}\sim\mathcal{N}(0,\sigma^{2})$ is a zero-mean Gaussian random variable with standard deviation $\sigma\in\{1,2,3,4\}$.
  2. 2.
    ICML 2023: We collect ownership relations, review scores as well as the authors’ reported ranking on their papers from ICML 2023, based on survey from OpenRank.cc. We only keep the authors (and their papers) who have participated in the survey and provided a ranking of their submissions. For each paper, we split its review scores by their reviewers’ confidence level. We set the review score with the least confidence level as the paper’s raw review score $\bm{y}$, and the average score of remaining reviews as the paper’s ground-truth score $\bm{R}$.
Notably, we adopted two different approaches to generate the ground-truth scores in the two datasets above. While neither approach can perfectly mimic the ground-truth score, they help us understand the effectiveness of the Isotonic Mechanism from different perspectives. For the ICLR dataset, we generate the ground-truth scores based on the i.i.d. review score noises assumption. In this way, we are able to empirically test the Isotonic Mechanism under the score generation model of this paper. For the ICML dataset, we use the high confidence review scores (most likely from domain experts) as the surrogate to approximate the ground-truth scores. While this score generation process does not necessarily match our model, it allows us to directly observe whether the Isotonic Mechanism can adjust review scores with high noise towards the consensus of experts, which tends to better align with the ground-truth score.
We use the Mechanism 2 with uniform owner credential ($\alpha^{j}=1,\forall j$) to determine the adjusted score $\widehat{\bm{R}}$ from review scores y and the elicited ranking information. For comparison, we consider the following alternative calibration approaches, evaluated under the mean squared error (MSE), $\frac{1}{n}\|\bm{R}-\widehat{\bm{R}}\|^{2}$: baseline, which directly uses the raw review scores; random, which uses a randomly generated partition for the Isotonic Mechanism.
Table 4 summarizes the basic statistics of ICLR 2021-2023 that we collected as well as the partitions determined from their ownership sets. There is a clear growing trend in conference size in terms of submission and author number. Moreover, from both partition objective functions, the greedy partition is a clearly better choice than the randomly generated partition, though we are unable to empirically examine the approximation ratio of Greedy (Algorithm 4), since it is intractable to find the exactly optimal partition for such large instances given its NP-hardness.
#authors #papers $\operatorname{obj}({\mathcal{S}}_{\text{greedy}})$ $\operatorname{obj}({\mathcal{S}}_{\text{random}})$ $\operatorname{obj}^{\prime}({\mathcal{S}}_{\text{greedy}})$ $\operatorname{obj}^{\prime}({\mathcal{S}}_{\text{random}})$
ICLR 2021 8875 2964 5242 2515 1365 1132
ICLR 2022 10583 3328 6123 2966 1574 1324
ICLR 2023 15372 4881 9764 4600 2489 2045
Table 4: Statistics of ICLR 2021-2023, where ${\mathcal{S}}_{\text{greedy}},{\mathcal{S}}_{\text{random}}$ are greedy and random partitions generated in the ownership graph of each year’s conference, $\operatorname{obj},\operatorname{obj}^{\prime}$ are respectively the comparison-based and size-based objective, defined in Equations (2) and (3).
5.2 Experiment Results
Given the two rich sets of real world peer review data, we empirically study the following three questions about the performance of our proposed mechanism.
Q1: How does the mechanism’s calibration quality change under different levels of noise? Here we use the ICLR dataset where we can control the noise level in our simulation of ground-truth score. Figure 5 compares the empirical performance of the Isotonic Mechanisms with different baselines at different review noise levels. We can see that the proposed Isotonic Mechanism can mitigate a substantial amount of noise from the review score and improve the precision metric over the baseline approach. The percentage MSE loss in Figure 5 suggests our proposed method is able to mitigate about $10-20\%$ amount of review noise in the past three years of ICLR. In addition, the performance of the Isotonic Mechanisms steadily improves as the conference size grows — an encouraging sign for the applicability of the Isotonic Mechanism. This performance is also reflected in determining accepted paper (i.e., top 30%) as illustrated in the lower panel of Figure 5. It suggests that $5\%\sim 15\%$ of the mistakenly rejected papers can be rectified to acceptance through the proposed score calibration method. [^13] In Appendix F, we plot how MSE improvement influences the accuracy of more stringent paper selections, e.g., spotlight (top 5%) and oral (top 1.5%) presentations. The performance has similar trend of improvement, hence is omitted here to avoid repetition.
(a) MSE on ICLR 2021
(b) MSE on ICLR 2022
(c) MSE on ICLR 2023
(d) Accept Accuracy
(e) Accept Accuracy
(f) Accept Accuracy
Figure 5: Plots of mechanism performances under varied noise level $\sigma$. Upper: The mean square error (MSE) loss of score estimation (normalized as the percentage change, $\frac{\text{model}-\text{baseline}}{\text{baseline}}$); Lower: The accuracy of accepted paper, described by $\frac{|\widehat{\mathcal{P}}_{@30}\cap{\mathcal{P}}_{@30}|}{|{\mathcal{P}}_{@30}|}$, where ${\mathcal{P}}_{@30}$ is the set of truly top 30% papers (ICLR has acceptance rate  30%).
Q2: The informativeness and robustness tradeoff of the Isotonic Mechanism when there is authors’ perception noise. Our Isotonic Mechanism offers flexibility for practitioners to decide the number of authors from whom they would like to elicit ranking information (i.e., the $L$-strongness of the mechanism). Eliciting from a single author can lead to larger blocks in the partition hence more efficiency, but suffers more risk of miscalibration when the authors have noise perception of the ranking. Our second experiment tests such informativeness vs robustness tradeoff. However, the $L$ parameter can hardly go beyond 1 in the authorship graph of real conference, as the papers that share more one common authors are very rare. This requires us to generate problem instances with carefully structured ownership relations in order for a diverse set of partitions to emerge. We thus use pure simulation data here.
Specifically, in our simulation, the authors perceive score as $R_{i}+\zeta^{j}_{i}$ and control the variance of authors’ perception noise variable $\zeta^{j}_{i}$ in a range from $\{0.1,0.5,1,2\}$ — in a smaller scale than the review noise, as we expect that authors tend to have less noise perception of their papers than the reviewers (roughly conforming to our modeling assumption). Meanwhile, to vary the partition structure, we construct a class of instances whose connections can be characterized by a $7$-level ternary tree with $3^{7}$ items and $(3^{7}-1)/2$ authors. Intuitively, the partition at one extreme is to have only a single block of all items owned by the root node (author), and there is only one author we can elicit. At the other extreme, the partition is to have one block for the set of items owned by each leaf node, and we can elicit from the authors at this leaf node and all its ancestor nodes. Hence, there exist $L$-strong partitions with $L$ from 1 to 7. In Figure 6, we can observe a clear trend of tradeoff: in the case of low noise (blue curve), the large block size is preferred to have more pairwise information, while in the case of high noise (red curve), the smaller (though not necessarily the smallest) block size is preferred to have more authors per block (larger $L$). In cases of medium noise (green or orange curve), the optimal block size is a careful balance of the block size and $L$-strongness. The trend showcases the importance of eliciting from multiple authors, e.g., to enforce the $L$-strong partition constraint, though the parameters should be tailored carefully for the noise level in real-world applications.
Figure 6: The loss (MSE) of Isotonic Mechanisms under different levels of authors’ perception noise with varying number of authors per partition block and different review noise $\sigma=$ 1 (left), 2 (middle), 4 (right).
Q3: Evaluations on the real ICML 2023 data. Our last set of experiments focuses on the ICML dataset which is richer than the ICLR data since this data contains reported ranking from the real-world survey, hence allowing us to conduct almost all real experiments, except that the authors’ report in the survey would not affect the final outcomes of their paper. The performance of Isotonic Mechanisms is presented in Table 5. From these results, we are able to state with high confidence that Isotonic Mechanisms under both partition schemes are capable of reducing the noise in the review process.
Baseline Greedy Random
MSE $1.4267$ $1.3012$ $1.3210$
F-Statistics - $15.828$ $9.911$
P-Value - $7.032\times 10^{-5}$ $1.652\times 10^{-3}$
Table 5: The loss of Isotonic Mechanisms under different partition scheme in the ICML dataset, along with the results of F-test and P-value on whether the loss indeed decreases.
6 Conclusion and Additional Discussions
In this work, we have established the Isotonic Mechanism primarily designed to assist peer review in machine learning conferences, specifically addressing the challenges where multiple authors contribute to various papers. We demonstrate the truthfulness of our mechanism and design a simple polynomial-time algorithm to partition the ownership sets, which achieves a constant ratio approximation of the optimal partition that is computationally hard to find. This facilitates the practical implementation of our mechanism. However, our mechanism is not without limitations, offering avenues for future exploration. We begin with the applicability of the technical assumptions made in our theoretical analysis and point out the potential extensions and relaxations to investigate in future work.
Owner’s Private Information. Assumption 2.1 is a standard private knowledge assumption, commonly adopted in the mechanism design literature (e.g., in auction design it is often assumed that bidders know their values or private signals). Here we assume owners have accurate knowledge of the information we elicit, i.e., relative ranking of their own items. This assumption is not to ask owners to have precise knowledge on ground-truth scores of their items — only agreement on the (much coarser) ranking information suffices. We make a remark on the intrinsic tradeoff when eliciting information from owners. On one hand, eliciting very fine-grained score information can be risky since owners may not have accurate information. On the other hand, eliciting overly coarse information may not be useful any more. As a compromise, we believe the elicitation of ranking information is a reasonable “middle ground” to begin with, though it is an interesting future direction to further explore other possibilities. Additionally, in Section 5 where we evaluate our algorithm on situations with authors’ noisy perception of the true paper ranking, we still observe improvements on score estimation by the Isotonic Mechanisms, indicating that this approach remains statistically effective when owners can supply additional (even somewhat noisy) information not yet included in review scores.
That said, as demonstrated in the following Remark 5, our current mechanism cannot guarantee a truthful Nash equilibrium if some co-owners would “truthfully” report a misleading ranking according to his noisy knowledge of his papers’ relative ranking. The mechanism also ceases to be truthful if an owner can, to some degree, predict the realization of review noise, e.g., due to ex-ante insider information on the reviewers. Therefore, it remains a fundamental design challenge for future work to ensure the authors’ incentive compatibility when they hold outside information or conflicting information with each other. Nevertheless, we believe our current model and its truthfulness result remain valuable, as they encourage coauthors to discuss and refine their perceptions for more accurate rankings. We also conjecture that truthful behavior may be the most robust strategy when authors are uncertain about others’ response model, and we leave a formal analysis to future work.
Remark 5 (On the Strategyproofness of the Isotonic Mechanism).
[…]
$\mathbb{E}_{z}\bigl{[}\,\mathrm{U}(\pi_{2}=(1,3,2))\,\bigr{]}=5.33\quad>\quad 5.25=\mathbb{E}_{z}\bigl{[}\,\mathrm{U}(\pi_{2}=(1,2,3))\,\bigr{]}.$
This example demonstrates that, even in a two-owner, three-item scenario, if one owner’s report is non-truthful (e.g., corrupted by noise), it could be suboptimal for the other owner to report the ranking truthfully.
Noise Structures in Review Scores. The exchangeable noise structure in Assumption 2.2 is also a natural choice and has been widely adopted in mechanism design literature. The typical rationales behind this assumption are as follows: an owner determines the ranking of his paper ex-ante, i.e., before any submission and reviewer assignment have been made, a moment that he does not know how his papers will be reviewed. Thus, the best this owner could assume is the symmetry of review noises for all his papers, which is precisely described by our exchangeable noise assumption. Similar justification on such exchangeable noise assumption is also given in a recent work by [26], though for a completely different mechanism design problem. We point out that our Assumption 2.2 is a strict relaxation of the i.i.d. noise assumption, which is perhaps more often adopted in the peer review literature (See e.g., [1, 25, 38]). Notice that the noise satisfying Assumption 2.2 does not even need to have a zero mean. In addition, if the review score of each paper is averaged over several reviews and the number of reviewers assigned to each paper is treated as an i.i.d. random variable, then the exchangeable noises allow the papers to have different variances.
Finally, an advantage of the exchangeable noise assumption is that it is non-parametric, without the need of modeling the distribution family. If one is willing to assume specific format of the noise distributions, then the exchangeable noise assumption may be further relaxed (e.g., see recent work by Yan et al. [43] using exponential family to model review score distributions). However, without such structural distribution assumption, we conjecture it will be quite difficult (if not impossible) to relax this exchangeable noise assumption in our setup.
Owners’ Utility Structures in Peer Review. The convex utility in Assumption 2.3 may not appear as common as the two assumptions above, hence worth some more discussions here. To begin, it is important to distinguish two traits that affect utilities, the quality and quantity. While utility theory often assumes concavity in the quantity of products due to the law of diminishing return, it is not uncommon to assume utility’s convexity in quality, including the seminal work of Mussa and Rosen [28] on pricing quality-differentiated goods and its numerous follow-up works. [^14] More concretely, for tractability reasons, these papers often normalize utility to be linear (i.e., weakly convex) in quality by re-parameterization, but shift the convexity into cost functions. In our setting, utility convexity in ratings reflects the nature that items receiving higher rating tend to generate significantly more reward. We believe such convex utility is particularly suitable in the context of evaluating research (e.g., conference review), fundamentally due to the “high-risk-high-reward” nature of research, recognized by various funding agencies [42, 7, 12]. A spiritually similar context is in the tournament design [23, 30], where participants are incentivized to exert more effort or take on more risk to compete for the best as the potential rewards are significantly larger at the top; participants’ utilities are commonly assumed to be convex in their ratings (akin to the scores of papers in our setting). In fact, this assumption of convexity in scarcity is naturally consistent with the classic assumption of concavity in quantity. This is because, if a product is hardly accessible (e.g., papers with high scores), or an award is conferred to only a few, we tend to derive more utility — a phenomenon famously known as the “Veblen effects” in behavioral economics and social psychology [40, 41]. In line with the “Veblen effects”, most publication venues have indeed been implementing various metrics to promote (more scarce) high-quality research papers. For instance, with the massive number of accepted papers at AI/ML venues today, these conferences start to distinguish accepted papers with labels such as “poster”, “spotlight” and “oral”; oral papers are promoted with extra resources such as longer presentations and more official media highlights, which naturally lead to much higher utility than an accepted poster paper.
Figure 7: The first two plots based on the ICLR 2022 dataset illustrate the probability of an paper getting accepted as a “poster”, “spotlight”, “oral” w.r.t. its averaged review score. The dashed line denotes the estimated probability from raw data, the smooth line denotes the probability predicted by logistic regression. The last plot illustrates the expected utility curve based on a paper’s chance of receiving different acceptance labels.
[…]
Remark 6 (A case of the “convexifying” owner utilities in peer review).
The strong convexity of utility could result from not only the high-risk-high-reward nature of research but also the artful design of peer review systems. Existing measures include the selection of best papers and selective labels of “oral“ and “spotlight”. We observe that a potential redesign of the review rating metrics can even further convexify owners’ utility. Specifically, the method is to set less level choices for excellent papers (e.g., merging traditional “7:very good” and “8:excellent” to just a single score of “8:excellent”) while use more fine-grained rating for regular papers. Assuming the exchangeable review noises, such kind of designs will make the differences between regular papers wider while the differences between excellent papers narrower. An explicit goal of this mechanism is to better distinguish papers of different qualities at the acceptance borderline. Meanwhile, despite having narrower score differences, the papers of high ratings typically require much less deliberation (e.g., most of them are accepted with probability $>95\%$ at ICLR 2022). More importantly, this mechanism can implicitly encourage the convexity of acceptance curve by, intuitively, “stretching” the middle borderline region of the curve whereas “condensing” its high-score region. Such redesign has already taken place at ICLR 2024, where reviewers can only choose ratings from a list of non-linear scores (“1”,“3”,“5”,“6”,“8”,“10”) — we look forward to its impact on the ongoing review process.
To relax the current utility assumption, one direction is to study the utility functions that are monotone in the adjusted scores with the design of sequential review, despite the downside of prolonged peer review procedures [45]. Meanwhile, an author’s value of a paper could depend on factors such as authorship order, the length of the author’s list, and his relative contributions to the paper [10]. This modeling perspective could more accurately reflect the reality that authors may assign different values to their co-authored submissions. Recognizing the potential heterogeneity of the utility function poses a significant, yet crucial challenge for future work. Another direction is to understand the potential group manipulations or collusion under our mechanism. This is generally a challenging problem in mechanism design, as it requires incentive analysis beyond the unilateral deviation from Nash equilibrium — our Remark 5 shows that truthful behaviors do not form a dominant strategy Nash Equilibrium under the current setup. For future work, one could investigate how defense techniques used in paper assignment [32, Section 4.2] can be adopted to detect and mitigate group manipulations from paper authors.
Lastly, it is important to realize that theory might only take us so far in this field, due to the complex nature of human subjects. We view our work as a theoretical foundation for the initiatives aimed at eliciting self-evaluations and managing incentives in scientific reviews. While empirical results from the OpenRank survey at ICML 2023 are promising [35], further effort is required to implement this method in practice, where various challenges may arise. As an initial step, the Isotonic Mechanism can function as an independent layer atop existing peer review processes, with its calibrated scores serving as reference points to assist expert human decision-making (e.g., helping area chairs to detect anomalies and request additional reviewers). We look forward to more real-world experiments to better understand the potential limitations of the Isotonic Mechanism in practice and drive iterative improvements towards solving the owner-assisted calibration problem.
Acknowledgments
This work was supported in part by NSF through CCF-2303372 and CCF-1934876, ARO through W911NF-23-1-0030, Analytics at Wharton, and Wharton AI and Analytics for Business.
Appendix A Proof of Theorem 1
Before we dive into the proof, we introduce several definitions and lemmas for Theorem 1.
Definition 1 (Mixed Strategy Nash equilibrium).
In a mixed strategy Nash equilibrium, each owner $j$ adopts a mixed strategy $x^{j}$ as a distribution supported on all possible rankings. We say a profile of owners’ report $\{x^{j}\}_{j=1}^{m}$ forms a mixed strategy Bayes-Nash Equilibrium (NE) under mechanism $\mathcal{M}$ if for any owner $j\in[m]$, given others’ randomized report profile $x^{-j}=\{x^{j^{\prime}}\}_{j^{\prime}\neq j}$, the expected utility under $x^{j}$ is no worse than any other possible randomized report $\widetilde{x}^{j}$, i.e.,
$\mathop{\mathbf{E}}_{\bm{y}}\left[\mathop{\mathbf{E}}_{\pi^{j}\sim x^{j}}\mathop{\mathbf{E}}_{\pi^{-j}\sim x^{-j}}\mathrm{U}^{j}\left(\widehat{\mathrm{R}}_{\mathcal{M}}(x^{j},x^{-j};\bm{y})\right)\right]\geq\mathop{\mathbf{E}}_{\bm{y}}\left[\mathop{\mathbf{E}}_{\widetilde{\pi}^{j}\sim\widetilde{x}^{j}}\mathop{\mathbf{E}}_{\pi^{-j}\sim x^{-j}}\mathrm{U}^{j}\left(\widehat{\mathrm{R}}_{\mathcal{M}}(\widetilde{x}^{j},x^{-j};\bm{y})\right)\right].$
Definition 2 (Majorization).
We say $\bm{a}$ majorizes $\bm{b}$, denoted $\bm{a}\succeq\bm{b}$, if $\sum_{i=1}^{k}a_{(i)}\geq\sum_{i=1}^{k}b_{(i)},\forall k<n$ and $\sum_{i=1}^{n}a_{(i)}=\sum_{i=1}^{n}b_{(i)}$, where $a_{(1)}\geq\cdots\geq a_{(n)}$ and $b_{(1)}\geq\cdots\geq b_{(n)}$ are sorted in descending order from $\bm{a}$ and $\bm{b}$, respectively.
Lemma 2 (Hardy–Littlewood–Pólya inequality).
For any vector $\bm{a},\bm{b}\in\mathbb{R}^{n}$, the inequality
$\sum_{i=1}^{n}h(a_{i})\geq\sum_{i=1}^{n}h(b_{i})$
holds for all convex functions $h:\mathbb{R}\to\mathbb{R}$ if and only if $\bm{a}\succeq\bm{b}$.
Lemma 3.
For any vector $\bm{a},\bm{b}\in\mathbb{R}^{n}$, $\bm{a}^{+}+\bm{b}^{+}\succeq\bm{a}^{+}+\rho\circ\bm{b}^{+}$ for any permutation $\rho$ on $\bm{b}^{+}$, where $\bm{a}^{+}=(a_{(1)},a_{(2)},\dots,a_{(n)})$ and $\bm{b}^{+}=(b_{(1)},b_{(2)},\dots,b_{(n)})$
Proof.
Proof of Lemma 3 It is clear that $(\bm{a}^{+}+\bm{b}^{+})_{(i)}=\bm{a}^{+}_{i}+\bm{b}^{+}_{i}=a_{(i)}+b_{(i)}$. And for any permutation $\rho$, we have $(\rho\circ\bm{b}^{+})_{(i)}=\bm{b}^{+}_{(i)}=b_{(i)}$. Next, we prove the lemma by definition. The equality constraint holds since $\rho$ is a permutation. For any $k<n$, we have
$\sum_{i=1}^{k}(a_{(i)}+b_{(i)})=\sum_{i=1}^{k}a_{(i)}+\sum_{i=1}^{k}(\rho\circ\bm{b}^{+})_{(i)}\geq\sum_{i=1}^{k}(\bm{a}^{+}+\rho\circ\bm{b}^{+})_{(i)}.$
We have proved that $\bm{a}^{+}+\bm{b}^{+}\succeq\bm{a}^{+}+\rho\circ\bm{b}^{+}$ by definition. ∎
Lemma 4.
Suppose $\bm{a},\bm{a}^{\prime},\bm{b}$ are in the same descending order and $\bm{a}\succeq\bm{a}^{\prime}$, then $\bm{a}+\bm{b}\succeq\bm{a}^{\prime}+\bm{b}.$
Proof.
Proof of Lemma 4 We first verify that equality holds. Since $\sum_{i=1}^{n}a_{i}=\sum_{i=1}^{n}a_{i}^{\prime}$,
$\sum_{i=1}^{n}(a_{i}+b_{i})=\sum_{i=1}^{n}(a^{\prime}_{i}+b_{i}).$
For the inequality at any $k<n$, we have $\sum_{i=1}^{k}a_{i}\geq\sum_{i=1}^{k}a^{\prime}_{i}$. This readily implies that
$\sum_{i=1}^{k}(a_{i}+b_{i})\geq\sum_{i=1}^{k}(a^{\prime}_{i}+b_{i}).$
Therefore, by definition, we have $\bm{a}+\bm{b}\succeq\bm{a}^{\prime}+\bm{b}$. ∎
Theorem A.1 (Restatement of [36, Theorem 1]).
Under Assumption 2.1, 2.2, 2.3, it is optimal for a single owner to truthfully report the ranking of its items $\pi^{\star}$,
$\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}(\widehat{\bm{R}}(\pi^{\star}))\right]=\mathop{max}_{\pi}\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}(\widehat{\bm{R}}(\pi))\right].$
Proof.
Proof of Theorem 1 We first show that truthful reporting forms a Bayes-Nash equilibrium. Pick any owner and let its utility function be $\mathrm{U}(\widehat{\bm{R}})=\sum_{i=1}^{n}{U}(\widehat{R}_{i})$. Suppose that the other $m-1$ owners are already truth-telling and without loss of generality we can index them from $1$ to $m-1$. With slight abuse of notation, we let $\widehat{\bm{R}}^{j}(\pi)$ be the random variable for the score estimates using owner $j$’s reported ranking $\pi$ w.r.t. the randomness of review noise $\bm{z}$, i.e, $\mathop{\mathbf{Pr}}[\widehat{\bm{R}}^{j}(\pi)=\widehat{\mathrm{R}}^{j}(\pi;\bm{R}+\bm{z})]=\mathop{\mathbf{Pr}}[\bm{z}]$. Suppose that the first $m-1$ estimates use the true ranking and the true ranking $\pi^{\star}$ is the identity permutation for simplicity. It is clear that the item score estimates with all $m-1$ owners’ truthfully reported rankings are all identical,
$\widehat{\bm{R}}^{1}(\pi^{\star})=\widehat{\bm{R}}^{2}(\pi^{\star})=\cdots=\widehat{\bm{R}}^{m-1}(\pi^{\star}):=\widehat{\bm{R}}(\pi^{\star}).$
Suppose the $m$-th owner reports $\pi$. Pick any $\{\alpha^{j}\}_{j=1}^{m}$ such that $\sum_{j=1}^{m}\alpha^{j}=1$. Then, her expected utility is
$\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}(\alpha^{m}\widehat{\bm{R}}^{m}(\pi)+\sum_{j=1}^{m-1}\alpha^{j}\widehat{\bm{R}}^{j}(\pi^{\star}))\right]=\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}(\alpha^{m}\widehat{\bm{R}}^{m}(\pi)+(1-\alpha^{m})\widehat{\bm{R}}(\pi^{\star}))\right].$
Since the noise terms are exchangeable in distribution, the expectation can be replaced by averaging the term above with $n!$ permuted noise terms. Let $\widehat{\bm{R}}(\pi;\bm{z}):=\widehat{\bm{R}}(\pi;\bm{R}+\bm{z})$ and $\rho\circ\bm{z}$ be the noise vector $\bm{z}$ after permutation $\rho$. By the linearity of expectation, we have
$\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}(\alpha^{m}\widehat{\bm{R}}^{m}(\pi)+(1-\alpha^{m})\widehat{\bm{R}}(\pi^{\star}))\right]=\mathop{\mathbf{E}}_{\bm{z}}\left[\frac{1}{n!}\sum_{\rho}\mathrm{U}\left(\alpha^{m}\widehat{\mathrm{R}}^{m}(\pi;\rho\circ\bm{z})+(1-\alpha^{m})\widehat{\mathrm{R}}(\pi^{\star};\rho\circ\bm{z})\right)\right].$
Note that $\widehat{\mathrm{R}}^{m}(\pi;\rho\circ\bm{z})=\widehat{\mathrm{R}}(\pi;\rho\circ\bm{z})$ for any $\pi$. Then, it suffices to show that, for any vector $\bm{z}\in\mathbb{R}^{n}$,
$\sum_{\rho}\mathrm{U}\left(\alpha^{m}\widehat{\mathrm{R}}(\pi;\rho\circ\bm{z})+(1-\alpha^{m})\widehat{\mathrm{R}}(\pi^{\star};\rho\circ\bm{z})\right)\leq\sum_{\rho}\mathrm{U}\left(\alpha^{m}\widehat{\mathrm{R}}(\pi^{\star};\rho\circ\bm{z})+(1-\alpha^{m})\widehat{\mathrm{R}}(\pi^{\star};\rho\circ\bm{z})\right).$ (4)
Let $\bm{a}^{+}$ be the projection of $\bm{a}$ onto the standard isotonic cone $\{\bm{a}|a_{1}\geq a_{2}\geq\dots\geq a_{n}\}$. Recall that we assume without loss of generality that $R_{1}\geq R_{2}\dots\geq R_{n}$. Following from the coupling argument in the proof of Theorem 1 in [36], we have $\widehat{\mathrm{R}}(\pi^{\star};\bm{z})=(\bm{R}+\bm{z})^{+}$ and $\widehat{\mathrm{R}}(\pi;\bm{z})=\pi^{-1}\circ(\pi\circ\bm{R}+\pi\circ\bm{z})^{+},\forall\bm{z}\in\mathbb{R}^{n}$. So we can rewrite the inequality (4) as,
$\sum_{\rho}\mathrm{U}\left((1-\alpha^{m})(\bm{R}+\rho\circ\bm{z})^{+}+\alpha^{m}\pi^{-1}\circ(\pi\circ\bm{R}+\pi\circ\rho\circ\bm{z})^{+}\right)\leq\sum_{\rho}\mathrm{U}\left((\bm{R}+\rho\circ\bm{z})^{+}\right).$
We can prove this inequality with the following steps:
$\sum_{\rho}\mathrm{U}\left((1-\alpha^{m})(\bm{R}+\rho\circ\bm{z})^{+}+\alpha^{m}\pi^{-1}\circ(\pi\circ\bm{R}+\pi\circ\rho\circ\bm{z})^{+}\right)$
$\leq$ $\sum_{\rho}\mathrm{U}\left((1-\alpha^{m})(\bm{R}+\rho\circ\bm{z})^{+}+\alpha^{m}(\pi\circ\bm{R}+\pi\circ\rho\circ\bm{z})^{+}\right)$
$\leq$ $\sum_{\rho}\mathrm{U}\left((1-\alpha^{m})(\bm{R}+\rho\circ\bm{z})^{+}+\alpha^{m}(\bm{R}+\pi\circ\rho\circ\bm{z})^{+}\right)$
$=$ $\sum_{i=1}^{n}\sum_{\rho}U\left((1-\alpha^{m})(\bm{R}+\rho\circ\bm{z})^{+}_{i}+\alpha^{m}(\bm{R}+\pi\circ\rho\circ\bm{z})^{+}_{i}\right)$
$\leq$ $\sum_{i=1}^{n}\sum_{\rho}U\left((1-\alpha^{m})(\bm{R}+\rho\circ\bm{z})^{+}_{i}+\alpha^{m}(\bm{R}+\rho\circ\bm{z})^{+}_{i}\right)$
$=$ $\sum_{\rho}\mathrm{U}\left((\bm{R}+\rho\circ\bm{z})^{+}\right)$
The first inequality is due to Lemmas 2 and 3, as $\pi^{-1}$ is some permutation on $\bm{b}^{+}=(\pi\circ\bm{R}+\pi\circ\rho\circ\bm{z})^{+}$. The second inequality follows from Lemmas 2, 4 as well as Lemma 2.4 of [36], which shows that $(\bm{R}+\pi\circ\rho\circ\bm{z})^{+}\succeq(\pi\circ\bm{R}+\pi\circ\rho\circ\bm{z})^{+},\forall\rho$. The third inequality follows from Lemma 2 and 3 where for each distinct permutation $\rho_{j}$, $(1-\alpha^{m})(\bm{R}+\rho_{j}\circ\bm{z})^{+}_{i},\alpha^{m}(\bm{R}+\rho_{j}\circ\bm{z})^{+}_{i}$ are the $j$-th entry of $\bm{a}^{+},\bm{b}^{+}\in\mathbb{R}^{n!}$ and $\pi$ is some permutation on $\bm{b}^{+}$. The equalities are by the definition of the utility function and its linearity.
It now remains to prove that truthful NE gives every owner the highest equilibrium utility among all possible equilibria of the game. Again, we pick arbitrary owner $i$. Under truthful NE, the owner’s utility is $\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}^{i}(\widehat{\bm{R}}(\pi^{\star}))\right]$, where $\widehat{\bm{R}}(\pi^{\star})$ is the item score estimate under the truthful ranking. We compare it to any (possibly mixed) Nash equilibrium (see Definition 1). Consider any pure strategy profile in this mixed Nash equilibrium, $\bm{\pi}=(\pi^{1},\pi^{2},\dots,\pi^{m})\sim\bm{x}=(x^{1},x^{2},\dots,x^{m})$. This sampled profile of reported ranking corresponds to the weighted average of score estimates, $\sum_{j=1}^{m}\alpha^{j}\widehat{\bm{R}}(\pi^{j})$. We can see that the owner’s expected utility under this pure strategy profile is no larger than that under the truthful strategy profile,
$\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}^{i}\left(\sum_{j=1}^{m}\alpha^{j}\widehat{\bm{R}}(\pi^{j})\right)\right]$ $\leq\sum_{j=1}^{m}\alpha^{j}\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}^{i}(\widehat{\bm{R}}(\pi^{j}))\right]$ (5)
$\leq\mathop{max}_{j\in[m]}\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}^{i}(\widehat{\bm{R}}(\pi^{j}))\right]$
$\leq\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}^{i}(\widehat{\bm{R}}(\pi^{\star}))\right],$
where the first inequality is due to Jensen’s inequality and linearity of expectation, the second inequality is by the property of arithmetic mean, last inequality is due to Theorem A.1. Hence, the owner’s expected utility under this Nash equilibrium, taking an expectation over the distribution of pure strategy profile, $\mathop{\mathbf{E}}_{\bm{\pi}\sim\bm{x}}\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}^{i}(\sum_{j=1}^{m}\alpha^{j}\widehat{\bm{R}}^{j})\right]\leq\mathop{\mathbf{E}}_{\bm{z}}\left[\mathrm{U}^{i}(\widehat{\bm{R}}(\pi^{\star}))\right]$, is no larger than the expected utility under the truthful NE. ∎
Remark 7.
In fact, one may have noticed that the proof of payoff dominance above does not rely on whether the non-truthful strategy profile forms an equilibrium or not. Hence, a slightly stronger result holds, that is, the truthful NE gives every owner the highest utility among all possible strategy profiles of the game. However, we would like to also point out that this additional observation does not imply that the truthful report is necessarily a dominant strategy for any owner, i.e., the best response to any other strategy profile of the other owners. Nor is the truthful NE necessarily unique in this game. That said, given its payoff-dominant property, there is other strong evidence, especially from empirical human preference and behavioral theory, to expect the truthful equilibrium outcome, as pointed out in Remark 1.
Appendix B Proof of Proposition 1
We prove that for any ownership instance $\mathcal{O}=\{\mathcal{I}^{j}\}_{j\in[m]}$, we can construct a different ownership instance $\mathcal{O}^{\prime}=\{\overline{\mathcal{I}}^{j}\}_{j\in[m]}$ such that any $L$-strong partition ${\mathcal{S}}$ in $\mathcal{O}$ is a 1-strong partition in $\mathcal{O}^{\prime}$.
Pick any bipartite graph $\mathcal{G}=(\mathcal{A},\mathcal{P},\mathcal{E})$ from the ownership instance $\mathcal{O}=\{\mathcal{I}^{j}\}_{j\in[m]}$. Suppose we were to find $L$-strong partition ${\mathcal{S}}$ in $\mathcal{G}$. We can construct a bipartite graph $\mathcal{G}^{\prime}=(\mathcal{A}^{\prime},\mathcal{P}^{\prime},\mathcal{E}^{\prime})$, where the item set remains the same $\mathcal{P}^{\prime}=\mathcal{P}$, the owner set $\mathcal{A}^{\prime}$ corresponds to all the $L$ element subsets of $\mathcal{A}$, $\mathcal{A}^{\prime}=\{\ell|u^{\ell}\subseteq\mathcal{A},|u^{\ell}|=L\}$ and the edge set captures the common items of every $L$ owners, =$\mathcal{E}^{\prime}=\{(i,\ell)|i\in\bigcap_{j\in u^{\ell}}\mathcal{I}^{j}\}$. We show that for any $L$-strong partition ${\mathcal{S}}=\{{\mathcal{S}}_{1},{\mathcal{S}}_{2},\dots,{\mathcal{S}}_{K}\}$ of $\mathcal{G}$, there is an equivalent $1$-strong greedy partition ${\mathcal{S}}^{\prime}=\{{\mathcal{S}}^{\prime}_{1},{\mathcal{S}}^{\prime}_{2},\dots,{\mathcal{S}}^{\prime}_{K}\}$ of $\mathcal{G}^{\prime}$ such that ${\mathcal{S}}_{k}={\mathcal{S}}^{\prime}_{k},\forall k\in K$.
The proof is a straightforward verification of this bipartite graph construction. Pick any ${\mathcal{S}}_{k}\in{\mathcal{S}}$ with $|{\mathcal{T}}_{k}|\geq L$ in $\mathcal{G}$, there exists an ${\mathcal{S}}^{\prime}_{k}$ with $|{\mathcal{T}}^{\prime}_{k}|=1$ in $\mathcal{G}^{\prime}$. That is, we have ${\mathcal{T}}^{\prime}_{k}=\{\ell\}$ for some $u^{\ell}\subseteq{\mathcal{T}}_{k}$, then ${\mathcal{S}}^{\prime}_{k}={\mathcal{S}}_{k}\subseteq\bigcap_{j\in u^{\ell}}\mathcal{I}^{j}$ by construction. Therefore, ${\mathcal{S}}_{k}={\mathcal{S}}^{\prime}_{k},\forall k\in[K]$.
Appendix C Proof of Proposition 2
We consider the weight function $w(x)=\mathop{max}\{x,1\}-1$. For any partition with $K$ blocks, the objective function is $\operatorname{obj}({\mathcal{S}})=n-K$. We show that finding the optimal partition under such an objective function is at least as hard as finding the minimum set. We prove via a standard hardness reduction argument: for any set cover instance, we can construct a bipartite graph instance in which the partition that maximizes the objective function $\operatorname{obj}({\mathcal{S}})=n-K$ can be turned into a minimum set cover.
Pick arbitrary set cover instance with $m$ subsets of $n$ item, $\mathcal{V}_{1},\mathcal{V}_{2},\dots,\mathcal{V}_{m}$. We construct a bipartite graph $\mathcal{G}=(\mathcal{A},\mathcal{P},\mathcal{E})$ where $\mathcal{A}=[m],\mathcal{P}=[n]$ and =$\mathcal{E}=\{(p_{i},a_{j})|p_{i}\in\mathcal{V}_{j},j\in[m]\}$. There exists an optimal partition ${\mathcal{S}}^{*}$ with $\operatorname{obj}({\mathcal{S}}^{*})=n-K^{*}$, if and only if the minimum set cover has size $K^{*}$. For the “if” direction, we construct a partition ${\mathcal{S}}^{*}$ with $\operatorname{obj}({\mathcal{S}}^{*})=n-K^{*}$ from the set cover of size $K^{*}$. For each subset selected by the set cover, we construct a partition block with items in the subset minus the items already in the partition. This partition must be valid and have $K^{*}$ blocks. For the “only if” direction, we construct the set cover of size $K^{*}$ from the optimal partition ${\mathcal{S}}^{*}$. For every block ${\mathcal{S}}_{k}^{*}$ of the optimal partition ${\mathcal{S}}^{*}$, we pick any vertex $a_{j}\in{\mathcal{T}}_{k}^{*}$ from the corresponding owner set. Based on the $K^{*}$ distinct vertices, we select the subsets with the same indices, and it forms a set cover of all $n$ items by construction.
Appendix D Proof of Theorem 2
Proof.
Proof of Theorem 2 The proof is structured as follows. First, we characterize the necessary conditions under which Mechanism 3 is truthful, according to Lemma 1. Second, we show that for any Mechanism 3 under the necessary condition, a Mechanism 2 can be constructed to elicit as least as much ranking information.
Fix any ownership relation $\{\mathcal{I}^{j}\}_{j=1}^{m}$. For notational convenience, we let $\beta_{i}^{j}=0,\forall i\not\in\mathcal{I}^{j}$, as $\beta_{i}^{j}$ does not affect the mechanism output anyway. Our characterization of truthful Mechanism 3 is based on the following lemma (restatement of Lemma), which shows that truthful condition holds for Mechanism 3 only if its parameters admit certain partition structure as detailed in Equation (1).
Lemma 5 (Restatement of Lemma 1).
Under Assumptions 2.1, 2.2, and 2.3, if a Mechanism 3 with parameters $\big{\{}\mathfrak{S}^{j},\bm{\beta}^{j}\big{\}}_{j=1}^{m}$ is truthful, then the following two conditions must both hold:
  1. (I)
    Each owner has balanced influence on the items within each of its partition blocks for any input, i.e.,
    for $\text{for any }j\in[m],{\mathcal{S}}\in\mathfrak{S}^{j},i,i^{\prime}\in{\mathcal{S}},\qquad\text{we have }\,\omega^{j}_{i}=\omega^{j}_{i^{\prime}},$
    where $\omega_{i}^{j}=\alpha^{j}\beta^{j}_{i}/\sum_{j^{\prime}\in\mathcal{J}^{i}}\alpha^{j^{\prime}}\beta^{j^{\prime}}_{i}$ denotes owner $j$’s relative influence on the output score of item $i$.
  2. (II)
    The parameters $\big{\{}\mathfrak{S}^{j},\bm{\beta}^{j}\big{\}}_{j=1}^{m}$ has a valid partition structure in the following sense,
    for $\text{for any }j,j^{\prime}\in[m],{\mathcal{S}}\in\mathfrak{S}^{j},{\mathcal{S}}^{\prime}\in\mathfrak{S}^{j^{\prime}},\text{ if }{\mathcal{S}}\cap{\mathcal{S}}^{\prime}\neq\varnothing,\text{ then }\beta^{j}_{i}=\beta^{j}_{i^{\prime}},\text{ for all }i,i^{\prime}\in{\mathcal{S}}\cup{\mathcal{S}}^{\prime}.$ (6)
Moreover, these two conditions are equivalent to each other.
We start by observing a few corollaries from Lemma 1, which will be useful for our proof of Theorem 2. First, Condition (II) of the lemma is applicable to the situation with $j=j^{\prime}$, in which case ${\mathcal{S}}={\mathcal{S}}^{\prime}$ since other sets in $\mathfrak{S}^{j}$ does not overlap with ${\mathcal{S}}$. Hence we have $\beta^{j}_{i}=\beta^{j}_{i^{\prime}}$ for each $i,i^{\prime}\in{\mathcal{S}}$. Second, let us now consider the case $j\not=j^{\prime}$ and let ${\mathcal{S}}\in\mathfrak{S}^{j},{\mathcal{S}}^{\prime}\in\mathfrak{S}^{j^{\prime}}$ satisfying ${\mathcal{S}}\cap{\mathcal{S}}^{\prime}\neq\varnothing$. Then the lemma implies $\beta^{j}_{i}=\beta^{j}_{i^{\prime}},\text{ for all }i,i^{\prime}\in{\mathcal{S}}\cup{\mathcal{S}}^{\prime}$. A consequence of this is that either $\beta^{j}_{i}=0$ (i.e., $j$’s report does not affect item $i$), or $j$ must own every item in ${\mathcal{S}}\cup{\mathcal{S}}^{\prime}$ since $j$ has a non-zero value for every item in ${\mathcal{S}}\cup{\mathcal{S}}^{\prime}$. In the later case, if ${\mathcal{S}}^{\prime}\not\subseteq{\mathcal{S}}$, then some element $\overline{i}$ in ${\mathcal{S}}^{\prime}\setminus{\mathcal{S}}$ owned by $j$ must belong to another set $\overline{{\mathcal{S}}}\in\mathfrak{S}^{j}$. Condition (II) of Lemma 1 then implies that the weights in $\overline{{\mathcal{S}}},{\mathcal{S}}$ must all be the same, i.e., $\beta^{j}_{i}=\beta^{j}_{i^{\prime}}$ for any $i\in{\mathcal{S}}$ and $i^{\prime}\in\overline{{\mathcal{S}}}$.
Observations above hint at a constructive proof of the theorem by iteratively merging any two “unnecessarily isolated” partition blocks such as the ${\mathcal{S}},\overline{{\mathcal{S}}}$ as above in the parameters $\big{\{}\mathfrak{S}^{j},\bm{\beta}^{j}\big{\}}_{j=1}^{m}$, until it becomes a Mechanism 2. Formally, we will show that for any Mechanism 3 with a valid partition structure specified in Equation (1), we can construct a Mechanism 2 with some item set partition that elicits no less ranking information and thus has no worse statistical efficiency.
Our construction hinges on a MERGE operator that merges all those “unnecessarily isolated” partition blocks like the ${\mathcal{S}},\overline{{\mathcal{S}}}$ above. Specifically, if some owner $j\in[m]$ has two distinct blocks ${\mathcal{S}},\overline{{\mathcal{S}}}\in\mathfrak{S}^{j}$ that satisfy the following conditions — for any $i\in{\mathcal{S}},\overline{i}\in\overline{{\mathcal{S}}}$: (1) $\beta^{j}_{i}=\beta^{j}_{\overline{i}}$; and (2) $\beta^{j^{\prime}}_{i}=\beta^{j^{\prime}}_{\overline{i}}\neq 0$ for any other owner $j^{\prime}$ who fully owns all items in ${\mathcal{S}}\cup\overline{{\mathcal{S}}}$ — then MERGE combines ${\mathcal{S}},\overline{{\mathcal{S}}}$ as one set in $j$’s $\mathfrak{S}^{j}$ with the same parameter $\beta^{j}_{i}$ for any $i\in{\mathcal{S}}\cup\overline{{\mathcal{S}}}$. Notice that MERGE strictly increases the amount of ranking information elicited from the owners, as the orderings of items among merged blocks ${\mathcal{S}},\overline{{\mathcal{S}}}$ are now used in the calibration.
We exhaustively apply MERGE on the parameters for every owner’s every pair of item sets until there is no more valid merge, and then remove any block ${\mathcal{S}}$ with zero weights, i.e., $\beta^{j}_{i}=0,\forall i\in{\mathcal{S}}$. The last procedure does not affect the mechanism but makes our following analysis cleaner. It is also obvious that the merge process must terminate, since the merged block can only be as large as each owner’s item set. In addition, after each merge, the value of any ${\bm{\beta}}^{j}$ remains unchanged; it can be verified that the condition in Equation (1) holds for the updated partition blocks. Let the resulting parameters be $\big{\{}\widetilde{\mathfrak{S}}^{j},{\bm{\beta}}^{j}\big{\}}_{j=1}^{m}$. We claim that $\big{\{}\widetilde{\mathfrak{S}}^{j},{\bm{\beta}}^{j}\big{\}}_{j=1}^{m}$ satisfies that, for any $j,j^{\prime}\in[m]$, ${\mathcal{S}}\in\widetilde{\mathfrak{S}}^{j},{\mathcal{S}}^{\prime}\in\widetilde{\mathfrak{S}}^{j^{\prime}}$, ${\mathcal{S}},{\mathcal{S}}^{\prime}$ are either identical or disjoint.
We prove by contradiction: if this is not the case, then there must exist at least one valid merge. Suppose ${\mathcal{S}}\cap{\mathcal{S}}^{\prime}\neq\varnothing$. Since $\big{\{}\widetilde{\mathfrak{S}}^{j},{\bm{\beta}}^{j}\big{\}}_{j=1}^{m}$ satisfies the condition in Equation (1), we have shown above that for owner $j$, she must (1) either have $\beta^{j}_{i}=0,\forall i\in{\mathcal{S}}\cup{\mathcal{S}}^{\prime}$, (2) or $j$ must own every item in ${\mathcal{S}}\cup{\mathcal{S}}^{\prime}$ since $j$ has a non-zero value for every item in ${\mathcal{S}}\cup{\mathcal{S}}^{\prime}$. The first case is eliminated as blocks with zero weights are removed. In the second case, some element $\overline{i}$ in ${\mathcal{S}}^{\prime}\setminus{\mathcal{S}}$ owned by $j$ must belong to another set $\overline{{\mathcal{S}}}\in\mathfrak{S}^{j}$, and ${\mathcal{S}},\overline{{\mathcal{S}}}$ can be merged, which would reach the contradiction. To check that the two condition for a valid merge is satisfied, we can observe the following, according to the condition in Equation (1):
  1. 1.
    For this owner $j$, we have $\beta^{j}_{i}=\beta^{j}_{i^{\prime}}=\beta^{j}_{\overline{i}}$ for any $\overline{i}\in\overline{{\mathcal{S}}}\setminus{\mathcal{S}}^{\prime}$. Hence, $\beta^{j}_{i}=\beta^{j}_{i^{\prime}},\forall i,i^{\prime}\in{\mathcal{S}}\cup\overline{{\mathcal{S}}}$.
  2. 2.
    For any other owner $j^{\prime}$ who owns all items in ${\mathcal{S}}\cup\overline{{\mathcal{S}}}$, we pick two items $i_{1}\in{\mathcal{S}}\cap{\mathcal{S}}^{\prime}$ and $i_{2}\in({\mathcal{S}}^{\prime}\setminus{\mathcal{S}})\cap\overline{{\mathcal{S}}}$. Let $i_{1}\in{{\mathcal{S}}}_{1}$ and $i_{2}\in{{\mathcal{S}}}_{2}$ from $\mathfrak{S}^{j^{\prime}}$. Since ${\mathcal{S}}_{1}\cap{\mathcal{S}}\neq\varnothing$, $\beta^{j^{\prime}}_{i_{1}}=\beta^{j^{\prime}}_{i},\forall i\in{\mathcal{S}}$. Since ${\mathcal{S}}_{2}\cap\overline{{\mathcal{S}}}\neq\varnothing$, $\beta^{j^{\prime}}_{i_{2}}=\beta^{j^{\prime}}_{i},\forall i\in\overline{{\mathcal{S}}}$. Since ${\mathcal{S}}_{1}\cap{\mathcal{S}}\neq\varnothing$ and $i_{1},i_{2}\in{\mathcal{S}}$, $\beta^{j^{\prime}}_{i_{1}}=\beta^{j^{\prime}}_{i_{2}}$. Hence, $\beta^{j^{\prime}}_{i}=\beta^{j^{\prime}}_{i^{\prime}}=\beta^{j^{\prime}}_{i},\forall i,i^{\prime}\in\overline{{\mathcal{S}}}\cup{\mathcal{S}}$.
Finally, consider $\mathfrak{S}:=\bigcup_{j=1}^{m}\widetilde{\mathfrak{S}}^{j}$ (after removing repeated item sets). One can verify that any two blocks ${\mathcal{S}},{\mathcal{S}}^{\prime}\in\mathfrak{S}$ are disjoint, though the union of all its blocks is not necessarily $[n]$ (recall that zero weight blocks have been removed). We claim that Mechanism 2 with parameter $\mathfrak{S}$ is at least as statistically efficient as Mechanism 3 with parameter $\big{\{}\widetilde{\mathfrak{S}}^{j},{\bm{\beta}}^{j}\big{\}}_{j=1}^{m}$, and thus, Mechanism 3 with parameter $\big{\{}{\mathfrak{S}}^{j},{\bm{\beta}}^{j}\big{\}}_{j=1}^{m}$ before MERGE. To see this, let us first observe that Mechanism 3 under $\big{\{}\widetilde{\mathfrak{S}}^{j},{\bm{\beta}}^{j}\big{\}}_{j=1}^{m}$ proceeds in the following steps: (1) pick each distinct block ${\mathcal{S}}\in\mathfrak{S}=\bigcup_{j=1}^{m}\widetilde{\mathfrak{S}}^{j}$, (2) elicit rankings of items in ${\mathcal{S}}$ from any owner $j$ such that ${\mathcal{S}}\in\widetilde{\mathfrak{S}}^{j}$, (3) determine a weighted average of adjusted review scores based on the rank-calibrated score from those owners with weight $\alpha^{j}\beta_{i}^{j}$. These procedures are identical to Mechanism 2 with parameter $\mathfrak{S}$, except that Mechanism 2 elicits from all owners with complete ownership of a block ${\mathcal{S}}\in\mathfrak{S}$. This completes the proof of our theorem.
Proof.
Proof of Lemma 1
We will prove that the truthfulness of Mechanism 3 implies condition (I), and condition (I) implies condition (II). Finally, we will prove condition (II) also implies condition (I), showing their equivalence.
We start by proving the first step, i.e., truthfulness of Mechanism 3 implies condition (I). We prove by contradiction. Suppose that for some owner $j\in[m]$ and for one of her partition blocks ${\mathcal{S}}\in\mathfrak{S}^{j}$, there exists $i,i^{\prime}\in{\mathcal{S}}$ such that $\omega^{j}_{i}\neq\omega^{j}_{i^{\prime}}$. Without loss of generality, let $\omega^{j}_{i}<\omega^{j}_{i^{\prime}}$. We construct an instance with owner utility $\{\mathrm{U}^{j}\}_{j=1}^{m}$, review scores and ground-truth scores $\{y_{i},R_{i}\}_{i=1}^{n}$ such that the unbalanced influence would give the owner $j$ incentive to untruthfully report the pairwise order between $i,i^{\prime}$. Let $y_{i}=R_{i}>R_{i^{\prime}}=y_{i^{\prime}}$. For any other item in the block ${\mathcal{S}}$, let its review score and ground-truth score be equal, and be smaller than the ground-truth scores of $i,i^{\prime}$ — it is now possible to only misreport the ranking between $i,i^{\prime}$, since the pairwise order of $i,i^{\prime}$ is independent of other items in this block. In addition, since the owner’s report strategy in block ${\mathcal{S}}$ does not affect the items outside block ${\mathcal{S}}$, it suffices to compare the owner’s report strategy in block ${\mathcal{S}}$. Let the utility of owner $j$ be ${U}^{j}(\widehat{R})=\mathop{max}\{\widehat{R}-c,0\}$ for some constant $c$ to be specified below. The utility function is convex w.r.t. the item’s adjusted review score $\widehat{R}$, conforming to Assumption 2.3.
We now compare two reporting strategy of this owner: one is to truthfully report the ground-truth ranking, the other is to report that item $i^{\prime}$ is better than $i$, and the ground-truthful ranking for the rest of the items in this block. It suffices to show that the owner has strictly higher utility in the latter reporting strategy. Recall that in this problem instance, the review scores and ground-truth scores of the other items are equal, and are smaller than the ground-truth scores of $i,i^{\prime}$. Isotonic regression implies that the adjusted score of all other items remain the same as the ground-truth score in both reporting strategy, so it suffices to analyze adjusted score of $i,i^{\prime}$ and compare the owner $j$’s utility on these two items.
If the owner $j$ is truthful, the adjusted score is $\widehat{R}_{i}=R_{i},\widehat{R}_{i^{\prime}}=R_{i^{\prime}}$. If the owner $j$ chooses the misreporting strategy, the calibrated scores based on the owner $j$’s ranking are $\widetilde{R}_{i}^{j}=\widetilde{R}_{i^{\prime}}^{j}=\frac{1}{2}(R_{i^{\prime}}+R_{i})$, resulting in the final adjusted scores, $\widetilde{R}_{i}=\frac{1}{2}\omega_{i}^{j}(R_{i^{\prime}}+R_{i})+(1-\omega_{i}^{j})R_{i}$ and $\widetilde{R}_{i^{\prime}}=\frac{1}{2}\omega_{i^{\prime}}^{j}(R_{i^{\prime}}+R_{i})+(1-\omega_{i^{\prime}}^{j})R_{i^{\prime}}.$ Let $\varepsilon=\frac{R_{i}-R_{i^{\prime}}}{2}$; it can be verified that $\widetilde{R}_{i}={R}_{i}-\omega_{i}^{j}\varepsilon$ and $\widetilde{R}_{i^{\prime}}={R}_{i^{\prime}}+\omega_{i^{\prime}}^{j}\varepsilon$. Notice that since $0\leq\omega_{i}^{j}<\omega_{i^{\prime}}^{j}\leq 1$, we have ${R}_{i}\geq\widetilde{R}_{i}>\widetilde{R}_{i^{\prime}}>R_{i^{\prime}}$ and $\omega_{i^{\prime}}^{j}\varepsilon-\omega_{i}^{j}\varepsilon>0$. Pick any $c\in[R_{i^{\prime}},R_{i^{\prime}}+\omega_{i^{\prime}}^{j}\varepsilon-\omega_{i}^{j}\varepsilon)$, we can derive the owner $j$’s utility for the two items under the misreport and under the truthful report, respectively:
$U^{j}(\widetilde{R}_{i})+U^{j}(\widetilde{R}_{i^{\prime}})=$ $\mathop{max}\{{R}_{i}-\omega_{i}^{j}\varepsilon-c,0\}+\mathop{max}\{{R}_{i^{\prime}}+\omega_{i^{\prime}}^{j}\varepsilon-c,0\}$
$=$ ${R}_{i}-\omega_{i}^{j}\varepsilon-c+{R}_{i^{\prime}}+\omega_{i^{\prime}}^{j}\varepsilon-c$
$U^{j}(\widehat{R}_{i})+U^{j}(\widehat{R}_{i^{\prime}})=$ $\mathop{max}\{{R}_{i}-c,0\}+\mathop{max}\{{R}_{i^{\prime}}-c,0\}$
$=$ ${R}_{i}-c$
We can see that $U^{j}(\widetilde{R}_{i})+U^{j}(\widetilde{R}_{i^{\prime}})-U^{j}(\widehat{R}_{i})-U^{j}(\widehat{R}_{i^{\prime}})={R}_{i^{\prime}}+\omega_{i^{\prime}}^{j}\varepsilon-\omega_{i}^{j}\varepsilon-c>0$ (by our choice of $c$), so misreporting the ordering between $j$ and $j^{\prime}$ is strictly better than truthful report.
We now prove that the condition (I) and (II) are equivalent and, therefore, are both necessary condition for Mechanism 3 to be truthful.
(I)$\Rightarrow$(II): We prove by contradiction. Suppose that there exists $j,j^{\prime}\in[m],{\mathcal{S}}\in\mathfrak{S}^{j},{\mathcal{S}}^{\prime}\in\mathfrak{S}^{j^{\prime}}$ such that ${\mathcal{S}}\cap{\mathcal{S}}^{\prime}\neq\varnothing$ and there exists $i_{1},i_{2}\in{\mathcal{S}}\cup{\mathcal{S}}^{\prime},\beta^{j}_{i_{1}}\neq\beta^{j}_{i_{2}}$. We consider two possible cases:
  1. 1.
    If $i_{1},i_{2}\in{\mathcal{S}}$, then we can construct the problem instance with $\alpha^{j}=1$ being the only non-zero weight such that $\omega^{j}_{i_{1}}=\beta^{j}_{i_{1}}\neq\beta^{j}_{i_{2}}=\omega^{j}_{i_{2}}$, which is a contradiction to the given condition. Similar contradiction arises when $i_{1},i_{2}\in{\mathcal{S}}^{\prime}$.
  2. 2.
    Otherwise, let $i_{1}\in{\mathcal{S}}$ and $i_{2}\in{\mathcal{S}}^{\prime}$ without loss of generality. Since ${\mathcal{S}}\cap{\mathcal{S}}^{\prime}\neq\varnothing$, let $i_{0}\in{\mathcal{S}}\cap{\mathcal{S}}^{\prime}$. Note that the argument above already shows items within each block must have the same $\beta_{i}^{j}$ values, i.e., $\beta_{i_{0}}^{j}=\beta_{i_{1}}^{j}$ and $\beta_{i_{0}}^{j^{\prime}}=\beta_{i_{2}}^{j^{\prime}}$. Consequently, we have $\beta_{i_{0}}^{j}=\beta_{i_{1}}^{j}\neq\beta_{i_{2}}^{j}$ and $\beta_{i_{0}}^{j^{\prime}}=\beta_{i_{2}}^{j^{\prime}}$. We consider a problem instance with $\alpha^{j}=\alpha^{j^{\prime}}=1$ being only non-zero weight, thus $w_{i_{0}}^{j^{\prime}}=\frac{\beta_{i_{0}}^{j^{\prime}}}{\beta_{i_{0}}^{j}+\beta_{i_{0}}^{j^{\prime}}}$ and $w_{i_{2}}^{j^{\prime}}=\frac{\beta_{i_{2}}^{j^{\prime}}}{\beta_{i_{2}}^{j}+\beta_{i_{2}}^{j^{\prime}}}$. Since $i_{0},i_{2}\in{\mathcal{S}}^{\prime}$, we have $\omega^{j^{\prime}}_{i_{0}}\neq\omega^{j^{\prime}}_{i_{2}}$ because $\beta_{i_{0}}^{j^{\prime}}=\beta_{i_{2}}^{j^{\prime}}$ but $\beta_{i_{0}}^{j}\neq\beta_{i_{2}}^{j}$. Contradiction is reached.
(II)$\Rightarrow$(I): We consider arbitrary owner $j\in[m]$ and any partition block ${\mathcal{S}}\in\mathfrak{S}^{j}$. Without loss of generality, suppose $|{\mathcal{S}}|\geq 2$ since singleton set does not have item rankings. Pick any two distinct items $i,i^{\prime}\in{\mathcal{S}}$ in the block. We make the following observations from the condition in Equation (1):
  1. 1.
    $\beta_{i}^{j}=\beta_{i^{\prime}}^{j}$ for the owner $j$ herself. This is implied by instantiating Equation (1) to $j=j^{\prime}$, in which case ${\mathcal{S}}={\mathcal{S}}^{\prime}$ since other sets in $\mathfrak{S}^{j}$ does not overlap with ${\mathcal{S}}$. The condition applies to any two items in ${\mathcal{S}}$.
  2. 2.
    $\beta_{i}^{j^{\prime}}=\beta_{i^{\prime}}^{j^{\prime}}$, for any owner $j^{\prime}(\neq j)$ who own any of $i,i^{\prime}$, i.e., $j^{\prime}\in\mathcal{J}^{i}\cup\mathcal{J}^{i^{\prime}}$. This is implied by instantiating Equation (1) to the $j^{\prime},j$ and some ${\mathcal{S}}^{\prime}\in\mathfrak{S}^{j^{\prime}}$ such that ${\mathcal{S}}^{\prime}\cap{\mathcal{S}}\neq\varnothing$. Any items in ${\mathcal{S}}^{\prime}\cup{\mathcal{S}}$ such as $i,i^{\prime}$ must have the same $\beta$ values, i.e., $\beta_{i}^{j^{\prime}}=\beta_{i^{\prime}}^{j^{\prime}}$.
  3. 3.
    $\beta_{i}^{j^{\prime}}=\beta_{i^{\prime}}^{j^{\prime}}=0$, for any owner $j^{\prime}$ who own neither of $i,i^{\prime}$, i.e., $j^{\prime}\not\in\mathcal{J}^{i}\cup\mathcal{J}^{i^{\prime}}$. This follows from the mechanism construction where we set $\beta_{i}^{j^{\prime}}=0$ for any $i\not\in\mathcal{I}^{j^{\prime}}$.
The three cases above implies that, for any input credentials $\{\alpha^{j}\}_{j=1}^{m}$, we must always have $\omega^{j}_{i}=\omega^{j}_{i^{\prime}}$ by definition for any two items $i,i^{\prime}$ in any of $j$’s partition block ${\mathcal{S}}$, which proves the condition (I).
Appendix E Proof of Theorem 3
A Convenient Bipartite Graph View of the Ownership Relation. To ease our discussion in the formal proofs regarding the optimal partition, we start by representing the overlapping ownership model as a bipartite graph $\mathcal{G}=(\mathcal{P},\mathcal{A},\mathcal{E})$. The vertex sets $\mathcal{P},\mathcal{A},$ with $|\mathcal{P}|=n,|\mathcal{A}|=m$ correspond to the disjoint set of $n$ items (papers) and $m$ owners (authors), respectively. The edge set $\mathcal{E}\subseteq\mathcal{P}\times\mathcal{A}$ corresponds to the ownership relation: the owner $j$ owns the item $i$, if and only if there is an edge $(p_{i},a_{j})\in\mathcal{E}$ between the vertices $p_{i}\in\mathcal{P}$ and $a_{j}\in\mathcal{A}$. The number of edges $|E|=\sum_{j\in[m]}|\mathcal{I}^{j}|$ is precisely the $N$ in the theorem statement, which captures the order of the problem’s input size. Let $E=(e_{i}^{j})_{m\times n}$ be the $0-1$ bi-adjacency matrix of the ownership in which $e_{i}^{j}=1$ for each edge $(p_{i},a_{j})\in\mathcal{E}$. Hence, the set of items owned by $j$ is exactly the neighbor set of vertex $a_{j}$, i.e., $\mathcal{I}^{j}=\{p_{i}\in\mathcal{P}|e_{i}^{j}=1\}$.
We now formalize the partition-based scheme in Mechanism 2 using the bipartite graph characterization. Specifically, we consider the partition $\mathfrak{S}=\{{\mathcal{S}}_{1},\ldots,{\mathcal{S}}_{K}\}$ on the vertex set $\mathcal{P}$ of the bipartite graph. By definition of a partition, we have $\bigcup_{k=1}^{K}{\mathcal{S}}_{k}=\mathcal{P}$ and ${\mathcal{S}}_{k^{\prime}}\cap{\mathcal{S}}_{k}=\varnothing,\forall k\neq k^{\prime}\in[K]$. For each partition block ${\mathcal{S}}_{k}$, we denote ${\mathcal{T}}_{k}=\{j\in\mathcal{A}|{\mathcal{S}}_{k}\subseteq\mathcal{I}^{j}\}$ as the set of owners who owns all items in ${\mathcal{S}}_{k}$. For the blocks with $|{\mathcal{S}}_{k}|\leq 1$ or $|{\mathcal{T}}_{k}|\leq 0$, the mechanism simply uses the raw review score. Otherwise, by construction, the sub-bipartite-graph of $\mathcal{G}$ induced by each $({\mathcal{S}}_{k},{\mathcal{T}}_{k})$ and their incidence edges $\{(p_{i},a_{j})\in\mathcal{E}|p_{i}\in{\mathcal{S}}_{k},a_{j}\in{\mathcal{T}}_{k}\}$ form a complete bipartite graph, i.e., a complete overlapping ownership. Therefore, we can also visualize the partition of the partially overlapping ownership using the bi-adjacency matrix.
We now utilize the notations above to formally prove Theorem 3.
Proof.
Proof of Theorem 3 Pick any ownership instance represented as a bipartite graph $\mathcal{G}$ and any wellness function $w\in\mathcal{W}$. Let $\operatorname{OPT}(\mathcal{G}):=\operatorname{obj}(\mathfrak{S}^{*})$, where $\mathfrak{S}^{*}=\{{\mathcal{S}}_{k}^{*}\}_{k=1}^{K}$ is the optimal $1$-strong partition of $\mathcal{G}$. Consider the greedy algorithm that keeps picking the largest residual item set owned by some owner, until all items are assigned to some block, as described in Algorithm 4. Let $\mathfrak{S}$ denote the partition obtained by Algorithm 4. By construction, $\mathfrak{S}$ is 1-strong — i.e., there is at least one owner who owns all items of each block in $\mathfrak{S}$. We let $\operatorname{ALG}(\mathcal{G}):=\operatorname{obj}(\mathfrak{S})$ be the objective value of the partition obtained by Algorithm 4. The remainder of the proof will argue that the output $\mathfrak{S}$ simultaneously satisfies,
$\frac{\operatorname{ALG}(\mathcal{G})}{\operatorname{OPT}(\mathcal{G})}=\frac{\operatorname{obj}(\mathfrak{S})}{\operatorname{obj}(\mathfrak{S}^{*})}\geq\inf\big{\{}\frac{w(x)}{w^{\prime}_{-}(x)x}\big{|}w^{\prime}_{-}(x)>0,x\geq 2\big{\}},$
for every $w\in\mathcal{W}$.
Let $\mathcal{G}\setminus j$ denote the induced subgraph of $\mathcal{G}$ by removing the vertex $j$ as well as all its adjacent item nodes and corresponding edges from $\mathcal{G}$. Formally, in this induced subgraph $\mathcal{G}\setminus j=(\mathcal{P}^{\prime},\mathcal{A}^{\prime},\mathcal{E}^{\prime})$, we have $\mathcal{P}^{\prime}=\mathcal{P}\setminus\mathcal{I}^{j},\mathcal{A}^{\prime}=\mathcal{A}\setminus\{j\}$ and =$\mathcal{E}^{\prime}=\{(i,j)|(i,j)\in\mathcal{E}\text{ and }i\in\mathcal{P}^{\prime},j\in\mathcal{A}^{\prime}\}$.
We now prove the approximation ratio of the greedy algorithm by induction on the owner set’s size $m=|\mathcal{A}|$. For the base case, for any bipartite graph $\mathcal{G}=(\mathcal{P},\mathcal{A},\mathcal{E})$ with $|\mathcal{A}|=1$, it is clear that the greedy approach finds the exact optimal partition. For the inductive case, we show that if the approximation guarantee holds for any bipartite graph $\mathcal{G}=(\mathcal{P},\mathcal{A},\mathcal{E})$ with $|\mathcal{A}|\leq k$, then it must hold for any bipartite graph $\mathcal{G}$ with $|\mathcal{A}|=k+1$ as well.
Let $j$ be the very first owner picked by Algorithm 4 on the bipartite graph instance $\mathcal{G}$ with $|\mathcal{A}|=k+1$. Thus, the first block (i.e., item set) picked by the algorithm must be $\mathcal{I}^{j}$. Let $\mathfrak{S}^{*}$ be the optimal partition of $\mathcal{G}$ under weight $w$. Note that the algorithm output $\mathfrak{S}$ is independent of $w$, but $\mathfrak{S}^{*}$ typically depends on $w$. For each block ${\mathcal{S}}_{k}^{*}\in\mathfrak{S}^{*}$, let $p_{k}^{*}:=\left|{\mathcal{S}}_{k}^{*}\right|$ be the number of items in this block and $a_{k}:=\left|{\mathcal{S}}^{*}_{k}\cap\mathcal{I}^{j}\right|$ be the number of items covered by the greedy algorithm’s first choice $\mathcal{I}^{j}$. Hence, $\left|\mathcal{I}^{j}\right|=\sum_{k=1}^{K}a_{k}$. We can then bound the approximation ratio of the greedy algorithm by decomposing the objective value as follows,
$\frac{\operatorname{ALG}(\mathcal{G})}{\operatorname{OPT}(\mathcal{G})}=$ $\frac{w(\mathcal{I}^{j})+\operatorname{ALG}(\mathcal{G}\setminus j)}{\sum_{k=1}^{K}w(p^{*}_{k})-\sum_{k=1}^{K}w(p^{*}_{k}-a_{k})+\sum_{k=1}^{K}w(p^{*}_{k}-a_{k})}$
$\geq$ $\frac{w(\mathcal{I}^{j})+\operatorname{ALG}(\mathcal{G}\setminus j)}{\sum_{k=1}^{K}w(p^{*}_{k})-\sum_{k=1}^{K}w(p^{*}_{k}-a_{k})+\operatorname{OPT}(\mathcal{G}\setminus j)},$
where the inequality uses the suboptimality of the partition $\{{\mathcal{S}}^{*}_{k}\setminus\mathcal{I}^{j}\}_{k=1}^{K}$ for the graph $\mathcal{G}\setminus j$.
We will bound the approximation ratio by considering the two parts $\frac{w\left(\sum_{k=1}^{K}a_{k}\right)}{\sum_{k=1}^{K}w(p^{*}_{k})-\sum_{k=1}^{K}w(p^{*}_{k}-a_{k})}$ and $\frac{\operatorname{ALG}(\mathcal{G}\setminus j)}{\operatorname{OPT}(\mathcal{G}\setminus j)}$ separately, and then apply the mediant inequality to derive the lower bound. The second part can be easily bounded according to the induction hypothesis. However, the first part could be unbounded in general (this may happen when $p^{*}_{k}$’s are all at most 1). Fortunately, the analysis for such special case can be carried out separately. Specifically, if $\mathop{max}_{k\in[K]}p^{*}_{k}\leq 1$ in $\mathcal{G}$, by convexity of $w$ function, this implies that any 1-strong partition of the items are at least as good as $\mathfrak{S}^{*}$, thus the greedy algorithm is optimal.
Next, we consider the more typical situation with $p:=\mathop{max}_{k\in[K]}p^{*}_{k}\geq 2$. Here, invoking the mediant inequality and a technical Lemma 6 below with $p\geq 2$, we can lower bound the approximation ratio as follows,
$\frac{w(\mathcal{I}^{j})+\operatorname{ALG}(\mathcal{G}\setminus j)}{\sum_{k=1}^{K}w(p^{*}_{k})-\sum_{k=1}^{K}w(p^{*}_{k}-a_{k})+\operatorname{OPT}(\mathcal{G}\setminus j)}$
$\geq$ $\mathop{min}\bigg{\{}\frac{w\left(\sum_{k=1}^{K}a_{k}\right)}{\sum_{k=1}^{K}w(p^{*}_{k})-\sum_{k=1}^{K}w(p^{*}_{k}-a_{k})},\ \frac{\operatorname{ALG}(\mathcal{G}\setminus j)}{\operatorname{OPT}(\mathcal{G}\setminus j)}\bigg{\}}$
$\geq$ $\inf\left\{\frac{w(x)}{w^{\prime}_{-}(x)x}\big{|}w^{\prime}_{-}(x)>0,x\geq 2\right\}.$
Finally, for running time analysis, Algorithm 4 can be efficiently implemented in almost linear time using appropriately chosen data structures. Specifically, we employ the red-black tree data structure which can retrieve the maximum element $j^{*}=\arg\mathop{max}_{j\in[m]}|\mathcal{I}^{j}\setminus\overline{\mathcal{I}}|$, i.e., the maximal residual item sets among all owners, using $O(\log m)$ time. We will then update the item sets of all the owners who have at least one item in $\mathcal{I}^{j^{*}}\setminus\overline{\mathcal{I}}$. The number of operations is at most the number of edges connected to removed items from $j^{*}$ in the bipartite graph. We then update the red-black tree structure using the new sizes of the residual item sets of the involved items. Note that the update of each owner’s item set size takes $O(\log m)$ time by first deleting it in the red-black tree and then re-inserting it. The above recursion takes at most $m$ rounds since each round at least one owner’s residual item set will become empty and there are only $m$ owners. In total, the total running time is thus $O(|\mathcal{E}|\log m)$.
Lemma 6.
Consider any $\{p_{k}^{*}\}_{k\in[K]}$ and $\{a_{k}\}_{k\in[K]}$ such that $p^{*}_{k}\geq a_{k}\geq 0$ for any $k\in[K]$ and $\sum_{k=1}^{K}a_{k}\geq p:=\mathop{max}_{k\in[K]}p^{*}_{k}$. For any $w\in\mathcal{W}$ with $w(p)>0$, we have
$\frac{w\left(\sum_{k=1}^{K}a_{k}\right)}{\sum_{k=1}^{K}w(p^{*}_{k})-\sum_{k=1}^{K}w(p^{*}_{k}-a_{k})}\geq\inf_{x\geq p}\frac{w(x)}{w^{\prime}_{-}(x)x}.$
Proof.
Proof of Lemma 6 Any $w\in\mathcal{W}$ is convex, so its left derivative $w^{\prime}_{-}$ is non-deceasing such that $w(p^{*}_{k})-w(p^{*}_{k}-a_{k})\leq w^{\prime}_{-}(p_{k}^{*})a_{k}$. In addition, $w^{\prime}_{-}(p^{*}_{k})\geq 0$ since $w(0)=0,w(p^{*}_{k})\geq 0$. Let $\sum_{k=1}^{K}a_{k}=a$, we have
$\frac{w\left(\sum_{k=1}^{K}a_{k}\right)}{\sum_{k=1}^{K}w(p^{*}_{k})-\sum_{k=1}^{K}w(p^{*}_{k}-a_{k})}\geq\frac{w(a)}{\sum_{k=1}^{K}w^{\prime}_{-}(p^{*}_{k})a_{k}}.$
Since $a\geq\mathop{max}_{k\in[K]}p^{*}_{k}=p$, thus $w^{\prime}_{-}(p_{k}^{*})\leq w^{\prime}_{-}(a)$ for any $k$ by convexity of $w$. Since $w(p)>0$ and $w(0)=0$, we must have $0<w^{\prime}_{-}(p)\leq w^{\prime}_{-}(x)$ for any $x\geq p$. Therefore,
$\frac{w(a)}{\sum_{k=1}^{K}w^{\prime}_{-}(p^{*}_{k})a_{k}}\geq\frac{w(a)}{\sum_{k=1}^{K}w^{\prime}_{-}(a)a_{k}}=\frac{w(a)}{w^{\prime}_{-}(a)a}\geq\inf_{x\geq p}\frac{w(x)}{w^{\prime}_{-}(x)x}.$
E.1 Tightness of Greedy’s Approximation Ratio in the Monomial Class
One implication of Theorem 3 is that for $\alpha$-th degree polynomial function $w=|{\mathcal{S}}_{k}|^{\alpha}$, the greedy algorithm has $\frac{1}{\alpha}$-approximation. Next, we show that this approximation ratio is provably tight for every weight in the monomial class.
Consider an instance with $n=MN$ items and $m=M+L$ owners for some positive integer $L,M,N$ such that $M^{L-1}\mid N$ and $N(1-1/M)^{L}\geq L$. The ownership relations are as follows (see also Figure 8 for the visualization of the instance):
  • There are $M$ owners $1,\cdots,M$, each of whom owns a disjoint set of $N$ items. Thus, collectively, they own all the $MN$ items. Let the items owned by these owners as $\mathcal{I}^{1},\dots,\mathcal{I}^{M}$ respectively.
  • We denote the item owned by the remaining $L$ owners as $\mathcal{I}^{\ell+M}$ for $\ell\in[1,\dots,L]$. Let $\left|\mathcal{I}^{\ell+M}\cap\mathcal{I}^{\ell}\right|=(1-1/M)^{\ell-1}N/M+1$ and $\left|\mathcal{I}^{\ell+M}\cap\mathcal{I}^{\ell^{\prime}}\right|=(1-1/M)^{\ell-1}N/M,\forall\ell^{\prime}\neq\ell\in[M]$. Hence, $|\mathcal{I}^{\ell+M}|=(1-1/M)^{\ell-1}N+1$.
Observe that $\mathcal{I}^{M+1}$ is the largest item set and $\mathcal{I}^{M+2}$ becomes the largest item set after $\mathcal{I}^{M+1}$ is removed. Therefore, we can see that the greedy algorithm proceeds by picking the $\ell$-th owner of the last $L$ owners from $\ell\in[1,\dots,L]$ and lastly the first $M$ owners, i.e., when each of whom has remaining item less than $N-\sum_{\ell=1}^{L}(1-1/M)^{\ell-1}N/M=(1-1/M)^{L}N$. As such, the greedy partition achieves the objective of at least $\operatorname{ALG}(\mathcal{G})\leq\sum_{\ell=1}^{L}[(1-1/M)^{\ell-1}N+1]^{\alpha}+M(1-1/M)^{\alpha L}N^{\alpha}.$
On the other hand, consider a partition strategy that only picks the first $M$ owners, which achieves the objective $\operatorname{OPT}(\mathcal{G})=MN^{\alpha}$. Simply using it as a lower bound for the optimal objective, we can bound the approximation ratio as,
$\frac{\operatorname{ALG}(\mathcal{G})}{\operatorname{OPT}(\mathcal{G})}$ $\leq\frac{\sum_{\ell=1}^{L}[(1-1/M)^{\ell-1}N+1]^{\alpha}+M(1-1/M)^{\alpha L}N^{\alpha}}{MN^{\alpha}}$
$=\frac{\sum_{\ell=1}^{L}[(1-1/M)^{\ell-1}N+1]^{\alpha}}{MN^{\alpha}}+(1-1/M)^{\alpha L}.$
For any constant $M>1$, we have as $N,L\to\infty$,
$\lim_{N,L\to\infty}\frac{\sum_{\ell=1}^{L}[(1-1/M)^{\ell-1}N+1]^{\alpha}}{MN^{\alpha}}+(1-1/M)^{\alpha L}$ $=\lim_{N,L\to\infty}\sum_{\ell=1}^{L}(1-1/M)^{(\ell-1)\alpha}/M$
$=\frac{1/M}{1-(1-1/M)^{\alpha}}.$
Finally, as $M\to\infty$, the ratio matches with the worst case approximation guarantee $1/\alpha$,
$\lim_{M\to\infty}\frac{1/M}{1-(1-1/M)^{\alpha}}=\lim_{M\to\infty}\frac{1}{\alpha-1/M}=1/\alpha.$
[…]
Figure 8: Illustration of the worst case instance for the greedy algorithm. $l=1,\cdots,L$.
Meanwhile, despite the hardness result for the size-focused objective in Proposition 2, we would like to point out that the efficient algorithm for optimal partition under the $\alpha$-th degree polynomial objective may exist. In particular, consider the following instance with $m=n+1$, where the optimal partition cannot realize the set-cover problem for the hardness reduction. The owner $i$ owns the items $\{2i-1,2i\}$,$\forall i\in[n]$ and owner $n+1$ owns the items $\{1,3,\cdots,2n-1\}$. In this case, the optimal partition is not the minimum set cover, since
$\operatorname{obj}\left(\left\{\{1,3,\cdots,2n-1\},\{2\},\cdots\{2n\}\right\}\right)=n^{\alpha}+n>\operatorname{obj}\left(\left\{\{1,2\},\cdots,\{2n-1,2n\}\right\}\right)=n\times 2^{\alpha}.$
Hence, it remains an open question on the hardness of finding the optimal partition under the $\alpha$-th degree polynomial objective.
Appendix F Additional Experiments
In Figure 9, we show how MSE improvement influences the accuracy of more stringent paper selections, e.g., spotlight (top 5%) and oral (top 1.5%) presentations. We can observe that the calibration performance has similar trend of improvement as the poster acceptance.
(a) ICLR 2021 Oral
(b) ICLR 2022 Oral
(c) ICLR 2023 Oral
(d) ICLR 2021 Spotlight
(e) ICLR 2022 Spotlight
(f) ICLR 2023 Spotlight
Figure 9: The percentile precision of Isotonic Mechanisms for Oral and Spotlight under varying level review noise $\sigma$ in ICLR.